### v1

# Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study over Open-ended Question Answering

Abstract
Recent works integrating Knowledge Graphs (KGs) have led to promising improvements in enhancing reasoning accuracy of Large Language Models (LLMs). However, current benchmarks mainly focus on closed tasks, leaving a gap in the assessment of more complex, real-world scenarios. This gap has also obscured the evaluation of KGs’ potential to mitigate the problem of hallucination in LLMs. To fill the gap, we introduce OKGQA, a new benchmark specifically designed to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios. OKGQA is designed to closely reflect the complexities of practical applications using questions from different types, and incorporates specific metrics to measure both the reduction in hallucinations and the enhancement in reasoning capabilities. To consider the scenario in which KGs may have varying levels of mistakes, we further propose another experiment setting OKGQA-P to assess model performance when the semantics and structure of KGs are deliberately perturbed and contaminated. OKGQA aims to (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on methods and future directions for leveraging KGs to reduce LLMs’ hallucination. We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs. The code of this paper is released at https://anonymous.4open.science/r/OKGQA-CBB0.
1 Introduction
Contemporary LLMs are prone to producing hallucinations due to gaps in their knowledge (Gekhman et al., 2024; Lee et al., 2023), primarily when the training data contain misinformation, biases, or inaccuracies. These flaws can arise when we simply maximize the log-likelihood, leading to responses that may seem plausible, but are often irrelevant or incorrect (Weng, 2024). This issue is especially concerning in scenarios where accuracy and reliability are essential, such as healthcare (He et al., 2023) and science (Taylor et al., 2022).
To address this critical limitation, researchers employ diverse strategies to augment the LLMs by integrating external knowledge graphs (KGs) (Pan et al., 2024; Luo et al., 2023a; Hu et al., 2023; Sun et al., 2023; van Breugel et al., 2021). KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023). Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance. This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent. Due to their reliability and interpretability, KGs are considered a promising method to improve the reliability of LLM reasoning.
However, current benchmarks for testing the capabilities of these LLM+KG models are predominantly closed-ended, restricting responses to a limited set of entities/relations (Talmor et al., 2019; Mihaylov et al., 2018; Jin et al., 2020) or a set of logical forms (Yih et al., 2016; Talmor & Berant, 2018; Puerto et al., 2023) derived from specific facts of KG. Hence, they can only test a very limited subset of the LLM’s tendency to hallucinate, leaving a gap in the assessment of complex, real-world scenarios. Particularly, standard metrics such as FActScore (Min et al., 2023) and SAFE (Wei et al., 2024) for evaluating the hallucination rate of LLMs require open-ended settings, i.e., questions are phrased as a statement which requires a longer answer.
Motivation: Recognizing the research gap, our study aims to contribute to the existing knowledge from two perspectives: 1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and 2) conduct a comparative analysis to shed light on methods and directions for leveraging KGs to reduce LLMs’ hallucination.
To achieve this, we develop a new benchmark, Open-ended Knowledge-Graphs Question Answering (OKGQA), specifically designed to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios. OKGQA is designed to closely reflect the complexities of practical applications using questions of different types, and incorporates specific metrics to measure both the reduction in hallucinations and enhancements in reasoning capabilities. Specifically, OKGQA covers 850 queries from ten categories, representing different levels of real-world QA types. All queries are synthesized by LLMs with pre-defined templates mentioned in Table 1, and we use Wikipedia as references for all the entities mentioned in the query. All the questions are designed to be open-ended and cannot be simply answered by generating logical forms or retrieving specific facts from the knowledge graph. To consider the scenarios in which KGs may have varying levels of mistakes (i.e., attributes may be mislabeled, relations may not exist, etc.), we propose another experiment setup OKGQA-P (discussed in Section 3.2) to assess model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated. We assess the hallucination ratio and the quality of the response (detailed metrics can be found in Section 3.3) of the tested models.
Based on our experiments, we find that (1) retrieving KG information can indeed mitigate factual errors in LLMs’ responses, especially for queries requiring additional reasoning; (2) directly performing reasoning in the LLM itself (e.g., internal reasoning strategies like Chain-of-thoughts (Kim et al., 2023) and self-consistency (Wang et al., 2022)) may cause bias and hallucination; (3) Subgraphs generally achieve the best performance across different query types, especially for simpler types of queries; (4) Integrating KGs can effectively reduce hallucinations in LLMs, even when KGs are contaminated to a certain degree. Overall, our main contributions include:
  • We present OKGQA, a benchmark tailored for evaluating LLMs enhanced with KGs under open-ended, real-world question-answering scenarios. The benchmark extends the assessment of closed-ended question answering to an open-ended setting, which can further support the assessment of hallucination of LLMs.
  • To consider the scenarios in which KGs may have varying levels of mistakes, we propose another experiment setup OKGQA-P to assess model performance under conditions in which the semantics and structure of KGs are deliberately perturbed and contaminated.
  • We conduct a series of experiments on OKGQA and OKGQA-P, analyzing the effectiveness of various retrieval methods and LLMs of different magnitudes, providing insights for future research and development. Our results underscore the importance of integrating LLMs with KGs to help reduce hallucinations, even in circumstances where KGs are contaminated.
2 Related Work
Due to the stochastic decoding process of Large language models (LLMs), i.e., sampling the next token in the sequence, LLMs exhibit probabilistic behavior, potentially yielding varied outputs of the same input across different instances (Agrawal et al., 2023). In addition, they also face challenges in accurately interpreting phrases or terms when the context is vague and resides in a knowledge gap region of the model, leading to outputs that may sound plausible but are often irrelevant or incorrect. This “hallucinations" undermines the reliability of LLMs (Huang et al., 2023). One emerging research trend is enhancing LLMs through integrating external knowledge graphs (Agrawal et al., 2023). KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023). Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance. This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent. Due to their reliability and interpretability, KGs are considered a promising method to improve the reliability of LLM reasoning.
Researchers employ diverse strategies to augment the LLMs by integrating external KGs. For example, KAPING (Baek et al., 2023b) matches entities in questions to retrieve related triples from knowledge graphs for zero-shot question answering. Wu et al. (2023) finds that converting these triples into textualized statements can further enhance LLM performance. StructGPT (Jiang et al., 2023b) propose to convert user query into structured formats (e.g., SPARQL) for information extraction from KGs. Following the succuess of internal reasoning-enhancement methods like Chain-of-thoughts (CoT) (Wei et al., 2022), Reflexion (Shinn et al., 2024), and Tree-of-thoughts (ToT), He et al. (2022) propose “rethinking with retrieval" to use decomposed reasoning steps from CoT prompting to retrieve external knowledge, leading to more accurate and faithful explanations. IR-CoT (Trivedi et al., 2022) interleaves the generation of CoT with knowledge retrieval from corresponding KGs, iteratively guiding both retrieval and reasoning for multi-step questions. MindMap (Wen et al., 2023) introduce a plug-and-play approach to evoke graph-of-thoughts reasoning in LLMs. Similarly, Reasoning on Graphs (RoG) (Luo et al., 2023b) use KGs to create faithful reasoning paths based on various relations, enabling interpretable and accurate reasoning in LLMs.
However, current benchmarks for testing the capabilities of these LLM+KG models are predominantly closed-ended, restricting responses to a limited set of entities/relations or a set of logical forms derived from specific facts of KG. Hence, they can only test a very limited subset of the LLM’s tendency to hallucinate, leaving a gap in the assessment of complex, real-world scenarios. Particularly, standard metrics such as FActScore (Min et al., 2023) and SAFE (Wei et al., 2024) for evaluating the hallucination rate of LLMs require open-ended settings, i.e., questions are phrased as a statement which requires a longer answer. Compared with previous works, our proposed OKGQA is tailored for evaluating LLMs enhanced with KGs under open-ended, real-world question-answering scenarios. The benchmark extends the assessment of closed-ended question answering to an open-ended setting, which can further support the assessment of hallucination of LLMs.
3 OKGQA: A Open-ended Knowledge Graph Question-Answering Benchmark
OKGQA is a benchmark designed to assess LLMs enhanced with KGs under open-ended, real-world questions answering scenarios. OKGQA is designed to closely reflect the complexities of practical applications using questions of different types and incorporates specific metrics to measure both the reduction in hallucinations and enhancement in reasoning capabilities. In this section, we discuss the construction of the dataset, and the construction of OKGQA-P which is a variant of OKGQA where the KGs’ semantics and structure are deliberately perturbed and contaminated.
Motivation of Open-ended QA instead of Close-ended QA: Current benchmarks for testing the capabilities of LLM+KGs models are predominantly close-ended [^1] https://en.wikipedia.org/wiki/Open-ended_question#Examples, which demand a short answer such as ‘yes’ and ‘no’. This restricts the responses to a limited set of entities/relations (Talmor et al., 2019; Mihaylov et al., 2018; Jin et al., 2020) or a set of logical forms (Yih et al., 2016; Talmor & Berant, 2018; Puerto et al., 2023) derived from specific KG facts. Hence, they can only test a very limited subset of the LLM’s tendency to hallucinate, leaving a gap in the assessment of more complex, real-world scenarios. Particularly, standard metrics (like FActScore (Min et al., 2023) and SAFE (Wei et al., 2024)) for evaluating the hallucination rate of LLMs require open-ended settings, i.e., questions are phrased as a statement which requires a longer answer.
Open-ended QA Task Definition: The tasks we design require to first understand the scope of the question, then optionally retrieve relevant information from multiple parts of the knowledge graph, then finally synthesize a coherent and informative response. The ideal output should be a paragraph that fully addresses the question with accurate and factual responses. We verify the response based on the metrics specified in Section 5 to reflect the models’ capabilities and faithfulness.
Statistics (on average)
Tokens in query 23.97
Total number of queries 850
Number of unique DBPedia entities 816
Before Pruning $\rightarrow$ After PPR Pruning
Tokens in subgraph 348,715 $\rightarrow$ 2,452
Number of nodes 7,171 $\rightarrow$ 48
Number of Edges 8,213 $\rightarrow$ 152
Avg. Degree 1.15 $\rightarrow$ 3.17
Clustering Coefficient 0.00 $\rightarrow$ 0.69
Graph Density 0.00 $\rightarrow$ 0.07
Query Type Simple Complex
Descriptive 78 11
Explanatory 195 56
Predictive 110 55
Comparative 72 74
Critical 182 17
Total 637 213
(a) Dataset statistics and query types
(b) Distribution of sub-query types
Figure 1: (left) Dataset statistics and query types, (right) Sub-query type distribution
3.1 Dataset Construction
Queries. We utilize a template-based methodology to systematically generate a diverse range of queries using LLMs, including categories such as descriptive, explanatory, predictive, comparative, and critical queries. Details regarding the specific templates and example queries can be found in Table 1, while the corresponding prompts are provided in the Appendix A.1. To ensure that the generated queries accurately represent real-world scenarios and complexities, we integrate the corresponding Wikipedia pages of each entity as contextual information for the LLMs’ generation. Furthermore, we follow a human-in-the-loop process to enhance the instructions for generating (the details are given in Appendix A.4). Intuitively, starting with an initial instruction, we generate a large number of query candidates, then use an LLM as automatic evaluator to evaluate the quality of the query, denoted as a set of scores from different metrics $s_{\text{auto}}$ ranging from 1 to 10, with higher scores indicating better performance. Then, these queries are manually evaluated and assign a set of human-label score $s_{\text{human}}$ correspond to $s_{\text{auto}}$, normalized as the same range as $s_{\text{auto}}$. We then optimize the input instructions by iteratively generate instructions to minimize the gap between $s_{\text{human}}$ and $s_{\text{auto}}$. These queries are categorized by difficulty and the naturalness of their phrasing. The statistics of the queries can be found in Figure 1.
Table 1: Query types and examples in OKGQA. Brown is used to highlight the placeholders (e.g., [person], [event]) in description, while Teal refers to the specific entities to replace the placeholders. The distribution of the sub-query types can be referred to in Figure 1(b).
Type Sub-Type Description / Template Example
Descriptive Character Description
Describe a [person]’s significant contributions
during their career.
Please describe Albert Einstein’s contributions to the
field of physics.
Event Description
Provide a detailed description of the background
and course of an [event].
Please provide a detailed description of the background
and course of the French Revolution.
Explanatory Cause Explanation
Why did [person] take [action] at [time]?
Why did Nixon choose to resign from the presidency
in 1974?
Relationship Explanation
Explain the relationship between [entity A] and
[entity B] and its significance.
Explain the relationship between Alexander the Great
and Aristotle and its significance.
Predictive Trend Prediction
Based on the historical behavior of [entity], what
do you think it might do in the future?
Based on Tesla’s historical behavior, in which fields do
you think it might innovate in the future?
Outcome Prediction
Based on the current situation, how do you predict
[event] will develop?
Based on the current international situation, how do you
predict climate change policies will develop?
Comparative Contrast Analysis
Compare and contrast the similarities and differences
between [entity A] and [entity B] in [aspect].
Compare and contrast the leadership styles of Steve Jobs
and Bill Gates.
Historical Comparison
Compare the impact of [historical event A] and
[historical event B].
Compare the impact of World War I and World War II
on the global order.
Critical Evaluation and Reflection
How do you evaluate the impact of [person/event]
on [field]? Please explain your viewpoint.
How do you evaluate Martin Luther King’s impact on
the civil rights movement? Please explain your viewpoint.
Application and Practice
How do you think [theory/method] can be applied
to [practical issue]?
How do you think machine learning technology can be
applied to medical diagnostics?
KG Sub-graphs. To reduce the size of KGs, following previous works (Luo et al., 2023a; Yih et al., 2016; Talmor & Berant, 2018), we construct subgraphs of DBpedia (all the queries in OKGQA can be reasoned based on DBPedia [^2] https://www.dbpedia.org/resources/knowledge-graphs/) by extracting all triples contained within the k-hop neighbors from the question entities in the query. For our experiments, we set k=2. To further reduce the size of the 2-hop subgraphs, we leverage personalized page-rank (PPR) (Bahmani et al., 2010) to prune the nodes/edges that are not relevant to the query (the details of the PPR algorithm are discussed in Appendix A.5). We compare the statistics of subgraphs before and after PPR pruning in Figure 1(a).
3.2 OKGQA-P: Benchmark with Noise & Perturbations in KGs
To further mimic the real situations where KGs may not be of high quality (i.e., attributes of nodes/edges may be mislabeled, relations may not exist, etc.), we propose another experiment setting OKGQA-P in this section to assess the model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated. Considering that KGs are typically annotated by humans and are generally accurate and meaningful, we introduce perturbations to edges in the KG to degrade the quality of the KGs, diminishing human comprehensibility. To quantify the degree of perturbation, we evaluate both the semantic and structual similarity between the original and the modified KG as defined below.
Notation. Let $\mathcal{F}_{\theta}$ be a KG-augmented model, and KG as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T})$, where $\mathcal{V}$ is the set of entities (nodes), $\mathcal{E}$ is the set of relation types (edges), and $\mathcal{T}=\{(v_{1},e,v_{2})|v_{1},v_{2}\in\mathcal{V},e\in\mathcal{E}\}$ is the set of triplets composed from existing entities and relations. Let $\mathcal{G}^{\prime}=(\mathcal{V},\mathcal{E^{\prime}},\mathcal{T^{\prime}})$ be the KG obtained after perturbing $\mathcal{G}$, where $\mathcal{E}^{\prime}\neq\mathcal{E}$ and $\mathcal{T}^{\prime}\neq\mathcal{T}$. Let $f(\mathcal{G},\mathcal{G}^{\prime})$ be a function that measures the similarity between $\mathcal{G}$ and $\mathcal{G}^{\prime}$. Let $g(\mathcal{G})$ be the downstream performance when evaluating $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}$.
High-level Procedure. First, we test $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}$ to get the original performance $g(\mathcal{G})$. Second, we perturb $\mathcal{G}$ to obtain $\mathcal{G}^{\prime}$. Third, we evaluate $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}^{\prime}$ to get the perturbed performance $g(\mathcal{G}^{\prime})$. Finally, we measure $g(\mathcal{G})-g(\mathcal{G}^{\prime})$ and $f(\mathcal{G},\mathcal{G}^{\prime})$ to assess how robust $\mathcal{F}_{\theta}$ is, i.e., to assess the model performance under conditions where KGs’ semantics and structure are deliberately perturbed.
To quantify how much the perturbed KG has deviated from the original KG, i.e., $f(\mathcal{G},\mathcal{G}^{\prime})$, we leverage metrics from Raman et al. (2020) for capturing semantics (ATS) and structural (SC2D, SD2) similarity between KGs. Intuitively, ATS leverages an pre-trained LM for link prediction to measure the probability of each edge from $\mathcal{G}^{\prime}$ existing in $\mathcal{G}$, while SC2D and SD2 measures the structural similarity between two KGs based on local clustering coefficient and degree distribution. For each of the three metrics, higher value indicates higher similarity. The detailed description can be found in Appendix A.7. The results after perturbation can be found in Figure 4.
For the perturbation methods, we consider four perturbation heuristics based on (Raman et al., 2020) as follows: Relation Swapping (RS) randomly chooses two edges from $\mathcal{T}$ and swaps their relations. Relation Replacement (RR) randomly chooses an edge $v_{1},e,v_{2})\in\mathcal{T}$, then replaces $e_{1}$ with another relation $e_{2}=\text{argmin}_{r\in\mathcal{R}}S_{\mathcal{G}}(v_{1},e,v_{2})$, where $S_{\mathcal{G}}(v_{1},e,v_{2})$ uses ATS to measure the semantics similarity between two edges. Edge Rewiring (ER) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, then replaces $v_{2}$ with another entity $v_{3}\in\mathcal{E}\backslash\mathcal{N}_{1}(v_{1})$, where $\mathcal{N}_{1}(v_{1})$ represents the 1-hop neighborhood of $v_{1}$. Edge Deletion (ED) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$ and deletes it. We control the perturbation level based on the percentage of KG edges being perturbed.
3.3 Evaluation Metrics
To quantify the hallucinations of LLMs, we leverage two public metrics, FActScore (Min et al., 2023) (factual precision in atomicity score) and SAFE (Wei et al., 2024) (search-augmented factuality evaluator). FActScore is designed to measure factual precision in text by decomposing a long form generation into multiple atomic facts and validates each separately against a reliable knowledge base, such as Wikipedia. We measure the proportion of facts that are supported by the knowledge source out of the total atomic facts. SAFE, on the other hand, adopts a more dynamic approach to fact-checking. It employs a language model as an investigative agent, which iteratively uses Google Search queries and reasons about whether the search results support or do not support the fact.
In addition to the hallucination metrics, we have also included four metrics, following previous works (Edge et al., 2024; Wang et al., 2023) to quantify the desirable qualities for generating sensible responses. Given the complexities inherent in our open-ended QA setting, collecting and annotating ground-truth answers pose significant challenges. Therefore, we leverage LLMs as automatic reference-free evaluators to assess these metrics based on G-Eval (Liu et al., 2023), a framework of using LLMs with chain-of-thought (CoT) and a form-filling paradigm. To enhance G-Eval’s robustness and stability, we provide the relevant Wikipedia pages as context in the prompt when calculating the metrics. The metrics are defined as follows: Context Relevance: measures how well the generated response aligns with the provided context. Comprehensiveness: evaluates the extent to which the generated answer covers all aspects and details of the question. Correctness: measures how clearly and specifically the generated answer responds to the question. Empowerment: evaluates how well the generated answer helps the reader understand the topic and make informed decisions. The detailed prompt for each metric can be found in Appendix A.1.
Models. We incorporate a range of widely used LLMs of different scales, including GPT-4o [^3] https://openai.com/index/hello-gpt-4o/, GPT-4o-mini [^4] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, Llama-3.1-8B-instruct (Dubey et al., 2024), Mistral-7B-instruct-v0.3 (Jiang et al., 2023a), and Gemma-2-9B-it (Team et al., 2024). We use the embedding model (text-embedding-3-small) from OpenAI for the G-retrieval process, considering cost and performance.
4 Exploring KG-augmented framework for reducing hallucination
Figure 2: Overview of KG-augmented framework for reducing hallucinations.
Overview. To explore whether KG-augmented approaches can mitigate LLMs’ hallucination, we investigated a range of methods and discovered that they can be encapsulated within a unified framework as shown in Figure 2. Our framework follows the paradigm of retrieval augmented generation (RAG) (Edge et al., 2024; Baek et al., 2023a), which retrieves essential information from the KGs, and then uses the retrieved knowledge to enhance the LLM’s generation. It has two components, i.e., Graph-guided retrieval (G-retrieval) and Graph-guided generation (G-generator), with a variety of algorithmic choices that combine to create diverse algorithms with different benefits for reducing hallucinations. We analyze our proposed strategies within each component in Section 5, aiming to shed light on the best practices for leveraging KGs for reducing hallucinations in LLMs.
Here we formalize the problem of context-relevant KG retrieval for open-ended question answering. Given a user query $q$, a generative pretrained language models (PLMs) first encodes the input tokens, and then models a conditional distribution $p(a|q)$ to generate an output response $a=[a_{1},\ldots,a_{T}]$. To explore whether KGs can help reduce hallucinations of PLMs, we introduce the retrieved knowledge from the KG, denoted $\mathcal{Z}\in\mathcal{G}$, which can take various forms (e.g., triplets, paths, subgraphs, as shown in Figure 2). This retrieved knowledge contains information relevant to the user query, and can either benefit the PLM at its knowledge boundaries (Baek et al., 2023b) or explicitly guide its reasoning process (Luo et al., 2023a; Sun et al., 2023). We explore which forms of knowledge should be extracted from KGs to best enhance the PLM. The distribution of the retrieved knowledge $\mathcal{Z}$ is modeled as $p(\mathcal{Z}|q,\mathcal{G})$, leading to a new likelihood for generating responses, expressed as $p(a|q)$.
$p(a|q)$ $=\sum_{\mathcal{Z}\subseteq\mathcal{G}}p_{\phi}(a|q,\mathcal{Z})p_{\theta}(\mathcal{Z}|q,\mathcal{G})$ (1)
$\approx p_{\phi}(a|q,\mathcal{Z}^{*})p_{\theta}(\mathcal{Z}^{*}|q,\mathcal{G}),$
where $p_{\theta}(\mathcal{Z}^{*}|q,\mathcal{G})$ represents the distribution of retrieving relevant knowledge elements from the KGs, and $p_{\phi}(a|q,\mathcal{Z}^{*})$ represents the distribution of leveraging the retrieved knowledge elements to generate the final responses, parameterized by $\phi$ and $\theta$, where $\mathcal{Z}^{*}$ is the optimal retrieved knowledge. Due to the potential exponential growth in the number of possible retrieved knowledge with graph size, efficient approximation methods are required. Consequently, the first line of Eq. 1 is approximated using the most probable retrieved knowledge $\mathcal{Z}^{*}$.
4.1 Graph-guided retrieval (G-retrieval)
The graph-guided retrieval phase focuses on extracting the most relevant elements (e.g., triplets, paths or subgraphs) from KGs to help answer the user query. Initially, the query is transformed into an embedding vector $\mathbf{q}\in\mathbb{R}^{d}$ using a language model encoder $f_{\text{LM}}(\cdot)$. For each node $v\in\mathcal{V}$ and edge $e\in\mathcal{E}$, their respective embeddings $\mathbf{e}_{v}$ and $\mathbf{e}_{e}$ are computed using the same language model to map all the information into the same vector space, and their relevance to the query is quantified using cosine similarity as: $s_{v}=\frac{\mathbf{q}\cdot\mathbf{e}_{v}}{\|\mathbf{q}\|\|\mathbf{e}_{v}\|}$ and $s_{e}=\frac{\mathbf{q}\cdot\mathbf{e}_{e}}{\|\mathbf{q}\|\|\mathbf{e}_{e}\|}$. Next, we identify the most relevant nodes and edges for the query based on the corresponding similarity, to yield a set of ‘top-$k$ nodes/edges’.
The goal of G-retrieval is to retrieve the knowledge that encompasses as many relevant nodes and edges as possible, while keeping the size manageable. To this end, we leverage a ‘prize’ and ‘cost’ trade-off strategy to guide the retrieval process, as follows: (1) Prize assignment: based on the computed similarity scores, we assign prizes to nodes and edges to quantify their relevance to the query. Specifically, we assign the top-$k$ nodes/edges with descending prizes values from $k$ to 1, while nodes and edges outside the top-$k$ receive a prize of 0. Formally: $p_{v}=\max(0,k-\text{rank}(v)+1)$ and $p_{e}=\max(0,k-\text{rank}(e)+1)$. (2) Cost allocation: to manage the retrieved knowledge size, we assign penalties as cost $C_{e}$ (by default assign to 1) during the expansion of the retrieved paths or subgraphs. These prizes and costs will be used in the retrieval process, by aiming to maximize the prizes while minimizing costs.
Triplet-retrieval: the triplet retrieval approach retrieves a fixed number $k$ of triplets with the highest total prize assigned to their respective head entity, relation, and tail entities.
Path-retrieval: based on the defined prize assignment and cost allocation, the path-retrieval approach starts from a fix number of $k$ of high-prize nodes to construct sequences $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$ by aiming to greedily maximize their scores: $S(\mathcal{P})=\sum_{i=1}^{n}p_{v_{i}}+\sum_{i=1}^{n-1}p_{e_{i}}-\sum_{i=1}^{n-1}c_{e}$. We use a priority queue to iteratively extend paths by selecting the highest-scoring extensions while avoiding cycles and adhering to maximum path length and quantity constraints. Finally, the top-scoring paths are sorted and returned. The details of path-retrieval can be found in Appendix A.6.
Sub-graph retrieval: the sub-graph retrieval approach also leverages the prize assignment and cost allocation; we follow He et al. (2024) to optimize the process using Prize-Collecting Steiner Tree (PCST) algorithm. The PCST problem seeks to build a connected subgraph $\mathcal{S}$ that maximizes the total prizes of included nodes and edges while minimizing the total cost of the edges. The score of the subgraph is defined as $S(\mathcal{S})=\sum_{n\in V_{\mathcal{S}}}p_{v_{i}}+\sum_{e\in E_{\mathcal{S}}}p_{e_{i}}-\sum_{e\in E_{\mathcal{S}}}c_{e}$. Unlike in path-retrieval, we only yield one subgraph that maximizes the total score.
4.2 Graph-guided generation (G-generator)
After retrieving $\mathcal{Z}^{*}$, the graph-guided generation phase focuses on learning to leverage this knowledge to generate the corresponding output to answer the user query. The generation is modeled as a sequential decision-making process as follows, where at each time step $t$, the next token $a_{t}$ is generated by G-Generator, conditioned on the query $q$, the retrieved knowledge $\mathcal{Z}^{*}$, and all previously generated tokens $a_{0:t-1}$.
$p(a|q,\mathcal{Z}^{*})=\prod_{t=1}^{T}p_{\theta}\left(a_{t}|q,\mathcal{Z}^{*},a_{0:t-1}\right)$ (2)
The model learns the parameters $\theta$ to maximize the likelihood of generating the correct output sequence by computing the probability distribution over possible next tokens at each time step. This ensures that the generated output not only considers the previously generated tokens but also the retrieved knowledge. The generation continues until a termination condition is met, such as generating an end-of-sequence token or reaching the maximum sequence length $T$.
5 Experiments
5.1 RQ1: Main Results - Can KGs Reduce Hallucination in LLMs?
To explore whether KGs can help reduce hallucination in LLMs, we benchmark the LLMs in different settings. We use zero-shot and few-shot prompting as baselines without injecting external knowledge. In addition, compared to adding external knowledge from KGs, we also consider leveraging LLMs’ internal knowledge to do Chain-of-thought (Kim et al., 2023), or self-consistency (Wang et al., 2022) prompting as baselines. As for LLMs augmented with KGs, we consider three KG retrieval forms: triplets, paths, and subgraphs (as shown in Figure 2) for comparison, to study the impact of G-retrieval for reducing LLMs’ hallucinations. The results can be found in Table 2 and Figure 3. We obtain some intriguing findings:
Table 2: Comparison results of various forms of information extracted from the KGs.
Models G-Eval Hallucination
Context Relevance Comprehensiveness Correctness Empowerment SAFE FActScore
Baseline: Without External Knowledge (Zero-shot prompting)
GPT-4o $68.12\%\pm 0.88\%$ $65.41\%\pm 0.79\%$ $60.41\%\pm 0.38\%$ $62.41\%\pm 0.84\%$ $82.47\%\pm 0.62\%$ $55.34\%\pm 0.93\%$
GPT-4o-mini $63.21\%\pm 0.49\%$ $60.11\%\pm 0.47\%$ $55.43\%\pm 0.63\%$ $58.72\%\pm 0.62\%$ $80.14\%\pm 0.89\%$ $50.23\%\pm 1.01\%$
llama-3.1-8b-instruct $57.12\%\pm 0.91\%$ $54.74\%\pm 1.20\%$ $49.01\%\pm 0.61\%$ $52.21\%\pm 0.71\%$ $79.33\%\pm 0.91\%$ $45.14\%\pm 0.32\%$
mistral-7B-Instruct-v0.3 $55.71\%\pm 1.21\%$ $52.00\%\pm 1.31\%$ $47.03\%\pm 0.94\%$ $50.13\%\pm 1.04\%$ $78.27\%\pm 0.83\%$ $44.37\%\pm 1.23\%$
gemma-2-9b-it $53.63\%\pm 1.33\%$ $50.00\%\pm 1.33\%$ $45.72\%\pm 0.71\%$ $48.15\%\pm 0.93\%$ $77.11\%\pm 0.78\%$ $40.94\%\pm 0.83\%$
Baseline: Without External Knowledge (4-shot prompting)
GPT-4o $70.61\%\pm 0.62\%$ $67.43\%\pm 0.81\%$ $62.33\%\pm 0.37\%$ $64.51\%\pm 0.12\%$ $83.39\%\pm 0.53\%$ $57.45\%\pm 0.78\%$
GPT-4o-mini $65.53\%\pm 0.94\%$ $62.33\%\pm 1.03\%$ $57.23\%\pm 0.68\%$ $60.47\%\pm 0.83\%$ $81.62\%\pm 0.69\%$ $52.34\%\pm 0.76\%$
llama-3.1-8b-instruct $59.43\%\pm 0.32\%$ $56.31\%\pm 0.78\%$ $51.27\%\pm 0.32\%$ $54.33\%\pm 0.41\%$ $80.27\%\pm 0.78\%$ $47.24\%\pm 1.03\%$
mistral-7B-Instruct-v0.3 $57.34\%\pm 1.04\%$ $54.13\%\pm 1.31\%$ $49.27\%\pm 0.84\%$ $52.46\%\pm 0.94\%$ $79.12\%\pm 0.87\%$ $45.13\%\pm 1.42\%$
gemma-2-9b-it $55.24\%\pm 1.49\%$ $52.27\%\pm 1.21\%$ $47.14\%\pm 0.36\%$ $50.36\%\pm 0.51\%$ $78.00\%\pm 0.77\%$ $44.32\%\pm 1.58\%$
Var-1: With CoT Prompting
GPT-4o - CoT $72.76\%\pm 0.92\%$ $69.56\%\pm 0.74\%$ $64.48\%\pm 0.63\%$ $66.69\%\pm 0.69\%$ $80.07\%\pm 0.83\%$ $54.30\%\pm 0.87\%$
GPT-4o - CoT+SC $75.81\%\pm 0.65\%$ $71.62\%\pm 0.74\%$ $66.55\%\pm 0.75\%$ $68.74\%\pm 0.15\%$ $79.03\%\pm 0.48\%$ $53.23\%\pm 0.78\%$
llama-3.1-8b-instruct - CoT $61.54\%\pm 0.95\%$ $58.35\%\pm 1.05\%$ $53.31\%\pm 0.71\%$ $56.42\%\pm 0.83\%$ $77.07\%\pm 0.85\%$ $46.15\%\pm 0.54\%$
llama-3.1-8b-instruct - CoT+SC $63.69\%\pm 0.32\%$ $60.44\%\pm 0.59\%$ $55.46\%\pm 0.52\%$ $58.53\%\pm 1.11\%$ $76.00\%\pm 0.63\%$ $45.05\%\pm 0.97\%$
mistral-7B-Instruct-v0.3 - CoT $59.58\%\pm 0.43\%$ $56.23\%\pm 2.31\%$ $51.28\%\pm 1.31\%$ $54.33\%\pm 0.72\%$ $75.04\%\pm 0.95\%$ $43.03\%\pm 1.03\%$
mistral-7B-Instruct-v0.3 - CoT+SC $61.35\%\pm 0.93\%$ $58.33\%\pm 1.02\%$ $53.42\%\pm 0.79\%$ $56.47\%\pm 0.85\%$ $74.30\%\pm 0.21\%$ $42.00\%\pm 0.29\%$
gemma-2-9b-it - CoT $57.34\%\pm 1.05\%$ $54.12\%\pm 0.32\%$ $49.27\%\pm 0.85\%$ $52.12\%\pm 0.95\%$ $72.07\%\pm 1.05\%$ $40.13\%\pm 0.49\%$
gemma-2-9b-it - CoT+SC $59.42\%\pm 0.27\%$ $56.27\%\pm 0.84\%$ $51.34\%\pm 1.42\%$ $54.34\%\pm 1.31\%$ $71.09\%\pm 0.43\%$ $39.85\%\pm 1.03\%$
Var-2: With Triplets Extracted from KGs Provided
GPT-4o $74.62\%\pm 0.65\%$ $70.44\%\pm 0.79\%$ $65.37\%\pm 0.72\%$ $67.12\%\pm 0.71\%$ $89.20\%\pm 1.42\%$ $72.53\%\pm 0.83\%$
GPT-4o-mini $69.50\%\pm 0.81\%$ $65.03\%\pm 0.92\%$ $60.21\%\pm 0.65\%$ $63.43\%\pm 1.01\%$ $87.52\%\pm 0.34\%$ $67.73\%\pm 0.95\%$
llama-3.1-8b-instruct $63.45\%\pm 1.13\%$ $59.33\%\pm 1.05\%$ $54.23\%\pm 0.75\%$ $57.33\%\pm 0.12\%$ $85.37\%\pm 0.72\%$ $62.37\%\pm 0.82\%$
mistral-7B-Instruct-v0.3 $61.34\%\pm 0.31\%$ $57.21\%\pm 0.89\%$ $52.29\%\pm 0.32\%$ $55.12\%\pm 0.43\%$ $84.21\%\pm 0.84\%$ $60.28\%\pm 1.05\%$
gemma-2-9b-it $59.25\%\pm 1.06\%$ $55.29\%\pm 0.44\%$ $50.15\%\pm 0.85\%$ $53.73\%\pm 0.95\%$ $83.18\%\pm 0.43\%$ $58.13\%\pm 0.91\%$
GPT-4o - CoT+SC $76.71\%\pm 0.53\%$ $72.34\%\pm 0.21\%$ $67.33\%\pm 1.31\%$ $69.64\%\pm 0.33\%$ $88.11\%\pm 0.57\%$ $71.45\%\pm 0.53\%$
Var-3: With Paths Extracted from KGs Provided
GPT-4o $78.71\%\pm 0.53\%$ $74.53\%\pm 0.31\%$ $69.42\%\pm 0.23\%$ $71.63\%\pm 0.61\%$ $90.20\%\pm 0.59\%$ $75.61\%\pm 0.51\%$
GPT-4o-mini $73.64\%\pm 0.93\%$ $69.41\%\pm 0.22\%$ $64.35\%\pm 0.72\%$ $67.52\%\pm 0.82\%$ $88.22\%\pm 0.34\%$ $70.53\%\pm 0.24\%$
llama-3.1-8b-instruct $67.51\%\pm 0.46\%$ $63.62\%\pm 1.39\%$ $58.41\%\pm 0.93\%$ $61.57\%\pm 0.94\%$ $86.33\%\pm 0.94\%$ $65.42\%\pm 0.95\%$
mistral-7B-Instruct-v0.3 $65.48\%\pm 0.94\%$ $61.37\%\pm 1.01\%$ $56.34\%\pm 0.23\%$ $59.45\%\pm 0.43\%$ $85.26\%\pm 0.85\%$ $63.31\%\pm 1.33\%$
gemma-2-9b-it $63.35\%\pm 1.37\%$ $59.23\%\pm 0.91\%$ $54.31\%\pm 0.91\%$ $57.41\%\pm 0.27\%$ $84.13\%\pm 0.21\%$ $61.23\%\pm 1.04\%$
GPT-4o - CoT+SC $80.87\%\pm 0.42\%$ $76.60\%\pm 0.65\%$ $71.54\%\pm 0.53\%$ $73.79\%\pm 1.21\%$ $89.11\%\pm 0.63\%$ $74.53\%\pm 0.24\%$
Var-4: With Subgraphs Extracted from KGs Provided
GPT-4o $80.81\%\pm 0.43\%$ $76.63\%\pm 0.65\%$ $71.57\%\pm 0.51\%$ $73.70\%\pm 0.62\%$ $90.83\%\pm 0.63\%$ $75.33\%\pm 0.29\%$
GPT-4o-mini $75.70\%\pm 0.44\%$ $71.51\%\pm 0.83\%$ $66.43\%\pm 0.76\%$ $69.60\%\pm 0.65\%$ $88.71\%\pm 0.72\%$ $70.12\%\pm 0.87\%$
llama-3.1-8b-instruct $69.61\%\pm 0.84\%$ $65.45\%\pm 0.93\%$ $60.41\%\pm 0.65\%$ $63.42\%\pm 0.45\%$ $86.12\%\pm 0.35\%$ $65.44\%\pm 0.87\%$
mistral-7B-Instruct-v0.3 $67.55\%\pm 0.87\%$ $63.35\%\pm 0.43\%$ $58.37\%\pm 0.71\%$ $61.45\%\pm 0.32\%$ $85.21\%\pm 0.81\%$ $63.12\%\pm 0.94\%$
gemma-2-9b-it $65.45\%\pm 0.95\%$ $61.23\%\pm 1.0\%$ $56.31\%\pm 0.35\%$ $59.40\%\pm 0.85\%$ $84.51\%\pm 0.99\%$ $63.74\%\pm 0.49\%$
GPT-4o - CoT+SC $82.90\%\pm 0.57\%$ $78.72\%\pm 0.61\%$ $73.64\%\pm 0.43\%$ $75.80\%\pm 0.75\%$ $89.12\%\pm 0.94\%$ $75.42\%\pm 1.31\%$
Retrieving KG information can indeed mitigate factual errors in the responses. Methods integrating knowledge extracted from KGs show clear improvements in factual accuracy and comprehension scores compared to the baselines. For example, under Var-2 (triplet retrieval), GPT-4o achieves a FActScore of 72.55% $\pm$ 0.85%, which is a significant increase over the baseline score of 55.35% $\pm$0.95%. Moreover, these methods can be combined with strategies like CoT+SC, enhancing response quality with minimal increase in hallucination ratio. The radar chart in Figure 3 further emphasizes that in most query types, integrating knowledge retrieved from KGs mitigates the hallucination issue compared to baselines, particularly in query types such as “Evaluation and Reflection," “Outcome Prediction," and “Cause Explanation," which require more reasoning and analysis rather than merely listing information. The findings also apply to open-source models like mistral-7B-Instruct-v0.3 and Llama-3.1-8B-instruct, illustrating the consistency of reliability of integrating KGs in LLM reasoning.
Figure 3: Comparison results of different forms of information over different queries.
Directly performing reasoning in the LLM itself does not mitigate hallucinations. We benchmark the hallucination ratio of LLMs using internal reasoning strategies like Chain-of-thought and Self-consistency, as shown in Var-1 in Table 2. It shows that these methods can indeed improve response quality (i.e., correctness, comprehensiveness and empowerment) compared to baselines, but do not consistently improve factuality, and sometimes even diminish factuality. This showcases that relying solely on internal reasoning does not mitigate hallucination, emphasizing the importance of external knowledge to counter the hallucination issue.
Subgraph retrieval generally achieves best performance across different query types, especially for simpler queries. We demonstrate the performance of different retrieval methods across different query types in Figure 3, showing that subgraphs generally achieve best performance. Especially for simpler queries (“Character Description" and “Event Description" which generally don’t require intensive reasoning), subgraph retrieval methods achieves better performance compared to paths and triplets. Even for queries like “Relationship Explanation" and “Cause Explanation" which require step-wise reasoning, subgraph methods still demonstrate promising performance. This suggests that while different forms of retrieved knowledge each offer unique benefits suitable for specific types of queries, subgraphs provide consistently good performance broadly.
5.2 RQ2: How Are KG-Aware Methods Affected by Noise / Perturbations in KGs?
We benchmark different KG-augmented LLMs on our OKGQA-P setting, where we deliberately perturb and contaminate the semantics and structure of KGs to simulate the real-world situation where KGs may not have high quality. Specifically, we consider different perturbation methods discussed in Section 3.2 and control the perturbation level based on the percentage of KG edges being perturbed. We first illustrate how much the perturbed KG has been deviated from the original KG with the increase of perturbation level, shown in Figure 4.
(a) Edge Deletion
(b) Edge Rewiring
(c) Relation Replacement
(d) Relation Swapping
Figure 4: Performance Metrics (ATS, SC2D, SD2) vs. Perturbation Level for Different Perturbation Methods.
It shows that the perturbation methods like edge deletion, rewiring and swapping have relatively weak influence on ATS (which intuitively measures semantic similarity), even as the perturbation level increases. For the edge deletion methods, only if the perturbation level reaches 1.0, the ATS goes to 0, otherwise, the ATS is very high. Relation replacement has a relatively strong impact on ATS, i.e., on the semantics of the perturbed KGs. For SC2D and SD2 (which intuitively measure structural similarity), the four methods show similar levels of decrease in most cases.
(a) Edge Deletion
(b) Edge Rewiring
(c) Relation Replacement
(d) Relation Swapping
Figure 5: Performance Metric (FActScore) vs. Perturbation Level for Different Perturbation Methods and Different Retrieval Methods. FS-T refers to FActScore metric using triplets, FS-P refers to using paths, and FS-SG refers to using sub-graphs.
Figure 5 illustrates the hallucination ratio using different methods on the perturbed KGs. We observe that (1) FS-SG consistently outperforms FS-T and FS-P even at higher perturbation levels, demonstrating its robustness by maintaining higher scores as perturbations increase. (2) FS-T and FS-P exhibit similar trends, each showing a significant performance drop as perturbation levels increase. Particularly, performance of FS-T and FS-P deteriorate when the perturbation level reaches 50%, i.e., becoming worse than the baseline using CoT. (3) On the setting using Relation Replacement which severely harms the semantics of the KGs, FS-T and FS-P decline more sharply than FS-SG. However, it still outperforms the baseline when the perturbation level is smaller than 40%.
In conclusion, our study demonstrates that leveraging external knowledge from KGs can effectively reduce hallucinations in LLMs, even when KGs are contaminated to a certain degree. In real-world applications, KGs such as Wikidata and unlikely to experience severe perturbations due to their continuous updates and community-driven quality control, making our findings applicable in most real-world scenarios.
6 Conclusion
In this paper, we present OKGQA, a benchmark tailored for evaluating LLMs enhanced with KGs under open-ended, real-world question answering scenarios. The benchmark extends the assessment from close-ended question answering to the open-ended setting, to support the assessment of hallucination in LLMs. To further mimic real-world scenarios where KGs may not have high quality, we propose another experiment setup OKGQA-P to assess model performance under conditions where KG’s semantics and structure are deliberately perturbed and contaminated. We conduct a series of experiments on OKGQA and OKGQA-P, analyzing the effectiveness of various retrieval methods and LLMs of different magnitudes, providing insights for further research and development. Our results underscore the significance of integrating KGs with LLMs to help reduce hallucination of LLMs, even in circumstances where the KGs are contaminated.
Appendix A Appendix
A.1 Prompt List
In this section, we present all the prompts required for the main experiments. To enhance clarity, we provide only one example in the prompt labeled as Example 1; the other few-shot examples utilized are labeled as Other In-Context Few-shots within the prompt.
A.1.1 Knowledge-Augmented Language Model Prompting
System Instruction: “You are a helpful assistant designed to answer the users’ open-ended questions. Your task is to provide accurate, concise, and useful information to foster understanding and solve problems. Whether the questions relate to complex scientific concepts, historical events, practical advice, or everyday life, your goal is to assist by offering thoughtful and informative responses."
Example 1:
Question: How do you think blockchain technology can be applied to secure online voting systems? Please provide specific suggestions?
Below are the facts that might be relevant to answer the question: (’Bitcoin’, ’wikiPageWikiLink’, ’Bill_Gates’), (’Bitcoin’, ’wikiPageWikiLink’, ’Non-fungible_token’), (’Decentralized_application’, ’wikiPageWikiLink’, ’Cardano_(blockchain_platform)’), (’Non-fungible_token’, ’wikiPageWikiLink’, ’Decentralized_application’), (’Hyperledger’, ’wikiPageWikiLink’, ’Blockchain’), (’Blockchain’, ’implementations’, ’Bitcoin’), (’Blockchain’, ’currencyName’, ’Litecoin’), (’Thing’, ’type’, ’Privacy_and_blockchain’), (’Thing’, ’type’, ’Saraju_Mohanty’), (’Thing’, ’type’, ’Blockchain’), (’Thing’, ’type’, ’Satoshi_Nakamoto’), (’Thing’, ’type’, ’Pseudonym’), (’Thing’, ’type’, ’Metaverse’), (’Thing’, ’type’, ’PricewaterhouseCoopers’)
Answer: Blockchain technology can significantly enhance the security of online voting systems through its decentralized, transparent, and immutable nature. By distributing data across multiple nodes, blockchain eliminates single points of failure, reducing the risk of hacking or vote tampering. The system allows for verifiable, yet anonymous, vote recording through cryptographic methods, ensuring voter privacy while preventing fraud such as double voting. Additionally, once recorded, votes cannot be altered, safeguarding election integrity. Smart contracts can further automate the voting process, enabling instant and accurate tallying, while the system’s resilience to cyberattacks like DDoS ensures uninterrupted functionality.
Other In-Context Few-shots
Question: {Question}
Below are the facts that might be relevant to answer the question: {Retrieved-knowledge}
Answer:
A.1.2 OKGQA Query Generation Prompting
System Instruction: Generate open-ended questions about different types: character description, event description, cause explanation, relationship explanation, trend prediction, outcome prediction, contrast analysis, historical comparison, evaluation and reflection, and application and practice.
Some templates are provided below: character description: describe a [person]’s significant contributions during their career. Example: Please describe Albert Einstein’s contributions to the field of physics. event description: Provide a detailed description of the background and course of an [event]. Example: Please provide a detailed description of the background and course of the French Revolution. cause explanation: Why did [person] take [action] at [time]? Example: Why did Nixon choose to resign from the presidency in 1974? relationship explanation: Explain the relationship between [entity A] and [entity B] and its significance. Example: Explain the relationship between Alexander the Great and Aristotle and its significance. trend prediction: Based on the historical behavior of [entity], what do you think it might do in the future? Example: Based on Tesla’s historical behavior, in which fields do you think it might innovate in the future? outcome prediction: Based on the current situation, how do you predict [event] will develop? Example: Based on the current international situation, how do you predict climate change policies will develop? contrast analysis: Compare and contrast the similarities and differences between [entity A] and [entity B] in [aspect]. Example: Compare and contrast the leadership styles of Steve Jobs and Bill Gates. historical comparison: Compare the impact of [historical event A] and [historical event B]. Example: Compare the impact of World War I and World War II on the global order evaluation and reflection: How do you evaluate the impact of [person/event] on [field]? Please explain your viewpoint. Example: How do you evaluate Martin Luther King’s impact on the civil rights movement? Please explain your viewpoint. application and practice: How do you think [theory/method] can be applied to [practical issue]? Please provide specific suggestions. Example: How do you think machine learning technology can be applied to medical diagnostics? Please provide specific suggestions. Generate the questions, the type of the questions, the placeholders, the naturalness of your generated questions (choose from high, medium, and unnatural), the difficulty of the generated questions (choose from hard, medium and easy) and dbpedia_entities (link the placeholders to dbpedia entities) in JSON format.
Other In-Context Few-shots
Generation:
A.2 Prompts for Instruction Tuner
Act as an "Instruction Tuner" for the LLM, you will be given the inputs: (1) the {Current Instruction} used to guide the LLMs’s evaluation, including specific examples with ground truth labels; (2) {Current Errors} that emerged with this instruction was applied to the dataset.
The current errors are presented in the following format: (1) INPUT: {input text} (2) PREDICTED OUTPUT: {predicted label}, (3) EXPECTED OUTPUT: {ground truth label}. Carefully analyze these errors and craft a revised concise instruction for the LLM to fit the expected outputs. Include 2-3 examples at the end of your response to demonstrate how the new instruction would be applied.
A.3 Metrics Prompt for G-eval
System Instruction: “You are a helpful assistant designed to evaluate the quality of the response to a query. Your task is to rate the response on one metric defined as below:"
Empowerment Criteria: Evaluate whether the ‘Actual Output’ can help the reader understand the topic and make informed decisions regarding the ‘Input’. A response with high empowerment provides accurate information and explanations that enhance the reader’s understanding. When evaluating empowerment, consider the relevance of the information provided in the ‘Actual Output’ to the ’Input’ and the ’Retrieval Context’.
Comprehensiveness Criteria: Evalute the extent to which the ’Actual Output’ covers all aspects and details of the question ’Input’. A comprehensive answer should thoroughly address every part of the question, leaving no important points unaddressed. When evaluating comprehensiveness, consider the relevance of the information provided in the ’Actual Output’ to the ’Input’ and the ’Retrieval Context’.
Correctness Criteria: Measure how clearly and specifically the ’Actual output’ responds to the question ’input’. A highly direct response stays focused on the question, providing clear and unambiguous information. When evaluating correctness, consider the relevance of the information provided in the ’Actual Output’ to the ’Input’ and the ’Retrieval Context’.
Context Relevance Criteria: Evaluate the extent to which the ’Actual output’ incorporates relevant information from the ’Retrieval Context’. This includes assessing whether the output adheres to the thematic, factual, and situational specifics presented in the ’Retrieval Context’. Relevant responses not only address the direct query but also align closely with the contextual elements provided, ensuring a seamless and coherent transition between the ’Retrieval Context’ and the ’Actual Output’. The most contextually relevant responses demonstrate an understanding and appropriate reflection of the given circumstances, historical facts, or conceptual background, thereby contributing to the overall accuracy and utility of the information provided.
A.4 Query Construction
In this section, we discuss the details of the query construction of OKGQA. It follows a template-based method with LLMs to generate a diverse range of queries. To ensure that the generated queries accurately represent real-world scenarios and complexities, we integrate the corresponding Wikipedia pages of each entity as contextual information for the LLMs’ generation. Furthermore, we follow a human-in-the-loop process to optimize the instruction used for generation, as shown in Figure 6. Intuitively, starting with an initial instruction, we generate a large number of query candidates, then use an LLM as automatic evaluator to evaluate the quality of the query, denoted as a set of scores from different metrics $s_{\text{auto}}$ ranging from 1 to 10, with higher scores indicating better performance. Then, these queries are manually evaluated and assign a set of human-label score $s_{\text{human}}$ correspond to $s_{\text{auto}}$, normalized as the same range as $s_{\text{auto}}$.We then optimize the input instructions by iteratively generate instructions to minimize the gap between $s_{\text{human}}$ and $s_{\text{auto}}$. This process is quite mimic the way of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and inherits the benefit that labeling the reward or penalty of the LLMs’ output is much easier than labeling the output directly.
Figure 6: Process of human-in-the-loop of query construction.
Specifically, we consider five angles to measure the quality of the generated queries: (1) Naturalness: assessing how fluid and human-like the query sounds; (2) Relevance: measuring whether the query pertains directly to the entity and the context provided; (3) Specificity: determining the level of detail and granularity included in the query, ensuring it is not too broad or vague; (4) Novelty: evaluating the uniqueness of the query, ensuring it is not just a repetitive or common question; (5) Actionability: gauging whether the query prompts clear, definite answers or actions that are feasible within the given context. Each of these angles contributes to a holistic evaluation of the query’s effectiveness and relevance in real-world applications.
A.5 Personalized PageRank (PPR)
In this section, we discuss the details of the PPR algorithm used in Section 1 to prune the graph from DBPedia and concentrate on nodes most pertinent to the central nodes of interest. The PPR is calculated using the iterative formula:
$\mathbf{p}=\alpha\mathbf{A}^{\top}\mathbf{p}+(1-\alpha)\mathbf{s},$ (3)
[…]
A.6 Prize-Cost-based Path Retrieval
The path-retrieval method is designed to construct and evaluate paths in a graph based on predefined prize assignments and cost allocations. The objective is to form sequences of nodes and edges, represented as $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$, that maximize the overall score and minimize the costs. To efficiently manage the exploration of potential paths, we utilizes a priority queue, a data structure that allows paths to be organized based on their scores, ensuring that the highest-scoring paths are processed first. The method starts by picking a number of starting nodes with high prizes. We then expand these starting points by exploring neighboring nodes. For each neighbor, the method calculates a new score. This score is the sum of the neighbor’s prize and the edge’s prize minus the edge’s cost. If this neighbor hasn’t been visited before, which helps avoid looping, the algorithm adds this neighbor to the path. This new path is then added to the priority queue. This expansion keeps going until the path reaches a maximum length or can’t be extended further. The algorithm keeps track of paths already explored to avoid repetition and ensure paths don’t loop back on themselves. When no more paths can be added or the priority queue is empty, the algorithm sorts the paths by their scores from highest to lowest.
A.7 KG Similarity Metrics
To measure how much the perturbed KGs are deviated from the original KG used in Section 3. We re-present the according metrics from (Raman et al., 2020) below:
Aggregated Triple Score (ATS)  ATS measures semantic similarity between two KGs. Let $s_{\mathcal{G}}$ be an edge (triple) scoring function, such that $s_{\mathcal{G}}(e_{1},r,e_{2})$ measures how likely edge $(e_{1},r,e_{2})$ is to exist in $\mathcal{G}$. Also, assume $s_{\mathcal{G}}$ has been pre-trained on $\mathcal{G}$ for link prediction. Then, ATS is defined as $f_{\text{ATS}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=\frac{1}{|\mathcal{T}^{{}^{\prime}}|}\sum_{(e_{1},r,e_{2})\in\mathcal{T}^{{}^{\prime}}}s_{\mathcal{G}}(e_{1},r,e_{2})\in[0,1]$, which denotes the mean $s_{\mathcal{G}}$ score across all edges in $\mathcal{G^{\prime}}$. Intuitively, if a high percentage of edges in $\mathcal{G^{\prime}}$ are also likely to exist in $\mathcal{G}$ (i.e., high ATS), then we say that $\mathcal{G^{\prime}}$ and $\mathcal{G}$ have high semantic similarity. $s_{\mathcal{G}}$ is task-specific, as KGs from different tasks may differ greatly in semantics. We use the $s_{\mathcal{G}}$ from (Li et al., 2016); While ATS captures semantic KG differences, it is not sensitive to KG connectivity structure. Note that $f_{\text{ATS}}(\mathcal{G},\mathcal{G})$ may not equal 1, since $s_{\mathcal{G}}$ may not perfectly generalize to KGs beyond those it was trained on.
Similarity in Clustering Coefficient Distribution (SC2D)  SC2D measures structural similarity between two KGs and is derived from the local clustering coefficient (Saramäki et al., 2007; Onnela et al., 2005; Fagiolo, 2007). For a given entity in $\mathcal{G}$ (treated here as undirected), the local clustering coefficient is the fraction of possible triangles through the entity that exist (i.e., how tightly the entity’s neighbors cluster around it). For entity $e_{i}\in\mathcal{E}$, the local clustering coefficient is defined as $c_{i}=2\text{Tri}(e_{i})/(\text{deg}(e_{i})(\text{deg}(e_{i})-1))$, where $\text{Tri}(e_{i})$ is the number of triangles through $e_{i}$, and $\text{deg}(e_{i})$ is the degree of $e_{i}$. For each relation $r\in\mathcal{R}$, let $\mathcal{G}^{r}$ be the subgraph of $\mathcal{G}$ consisting of all edges in $\mathcal{T}$ with $r$ That is, $\mathcal{G}^{r}=(\mathcal{E},r,\mathcal{T}^{{}^{\prime}})$, where $\mathcal{T}^{{}^{\prime}}=\{(e,r,e^{\prime})\hskip 2.84526pt|\hskip 2.84526pte,e^{\prime}\in\mathcal{E}\}$. Let $\mathbf{c}^{r}$ denote the $|\mathcal{E}|$-dimensional clustering coefficient vector for $\mathcal{G}^{r}$, where the $i$th element of $\mathbf{c}^{r}$ is $c_{i}$. Then, the mean clustering coefficient vectors for $\mathcal{G}$ and $\mathcal{G}^{\prime}$ are $\mathbf{c}_{o}=\frac{1}{|\mathcal{R}|}\sum_{r\in\mathcal{R}}\mathbf{c}^{r}$ and $\mathbf{c}_{p}=\frac{1}{|\mathcal{R}^{\prime}|}\sum_{r\in\mathcal{R}^{\prime}}\mathbf{c}^{r}$, respectively. SC2D is defined as $f_{\text{SC2D}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=1-\frac{\lVert\mathbf{c}_{o}-\mathbf{c}_{p}\rVert_{2}}{\lVert\mathbf{c}_{o}-\mathbf{c}_{p}\rVert_{2}+1}\in[0,1]$, with higher value indicating higher similarity.
Similarity in Degree Distribution (SD2)  SD2 also measures structural similarity between two KGs, while addressing SC2D’s ineffectiveness when the KGs’ entities have tiny local clustering coefficients (e.g., the item KG used by recommender systems is roughly bipartite). In such cases, SC2D is always close to one regardless of perturbation method, thus rendering SC2D useless. Let $\mathbf{d}^{r}$ denote the $|\mathcal{E}|$-dimensional degree vector for $\mathcal{G}^{r}$, where the $i$th element of $\mathbf{d}^{r}$ is $\text{deg}(e_{i})$. Then, the mean degree vectors for $\mathcal{G}$ and $\mathcal{G}^{\prime}$ are $\mathbf{d}_{o}=\frac{1}{|\mathcal{R}|}\sum_{r\in\mathcal{R}}\mathbf{d}^{r}$ and $\mathbf{d}_{p}=\frac{1}{|\mathcal{R}^{\prime}|}\sum_{r\in\mathcal{R}^{\prime}}\mathbf{d}^{r}$, respectively. SD2 is defined as $f_{\text{SD2}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=1-\frac{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}}{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}+1}\in[0,1]$, with higher value indicating higher similarity.

### v2

@@
 Abstract
-Recent works integrating Knowledge Graphs (KGs) have led to promising improvements in enhancing reasoning accuracy of Large Language Models (LLMs).
-However, current benchmarks mainly focus on closed tasks, leaving a gap in the assessment of more complex, real-world scenarios.
+Recent works integrating Knowledge Graphs (KGs) have led to promising improvements in enhancing the reasoning accuracy of Large Language Models (LLMs).
+However, current benchmarks focus mainly on closed-ended tasks, leaving a gap in the assessment of more complex real-world scenarios.
 This gap has also obscured the evaluation of KGs’ potential to mitigate the problem of hallucination in LLMs.
 To fill the gap, we introduce OKGQA, a new benchmark specifically designed to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios.
-OKGQA is designed to closely reflect the complexities of practical applications using questions from different types, and incorporates specific metrics to measure both the reduction in hallucinations and the enhancement in reasoning capabilities.
-To consider the scenario in which KGs may have varying levels of mistakes, we further propose another experiment setting OKGQA-P to assess model performance when the semantics and structure of KGs are deliberately perturbed and contaminated.
-OKGQA aims to (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on methods and future directions for leveraging KGs to reduce LLMs’ hallucination.
-We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs.
-The code of this paper is released at https://anonymous.4open.science/r/OKGQA-CBB0.
+OKGQA is designed to closely reflect the complexities of practical applications using questions from different types, and incorporates specific metrics to measure both hallucination ratio and the enhancement in reasoning capabilities.
+To consider the scenario in which KGs may have varying levels of mistakes, we propose another benchmark variant OKGQA-P to assess model performance when the semantics and structure of KGs are deliberately perturbed and contaminated.
+OKGQA aims to (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on method design.
+We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs to reduce hallucination.
+The code and data are released at link.
+Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study over Open-ended Question Answering
+Yuan Sui1, Yufei He1, Zifeng Ding2, Bryan Hooi1 1 National University of Singapore 2 University of Cambridge {yuansui, yufei.he, bhooi}@comp.nus.edu.sg, zd320@cam.ac.uk
 1 Introduction
@@
 This issue is especially concerning in scenarios where accuracy and reliability are essential, such as healthcare (He et al., 2023) and science (Taylor et al., 2022).
-To address this critical limitation, researchers employ diverse strategies to augment the LLMs by integrating external knowledge graphs (KGs) (Pan et al., 2024; Luo et al., 2023a; Hu et al., 2023; Sun et al., 2023; van Breugel et al., 2021).
-KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023).
-Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance.
-This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent.
-Due to their reliability and interpretability, KGs are considered a promising method to improve the reliability of LLM reasoning.
-However, current benchmarks for testing the capabilities of these LLM+KG models are predominantly closed-ended, restricting responses to a limited set of entities/relations (Talmor et al., 2019; Mihaylov et al., 2018; Jin et al., 2020) or a set of logical forms (Yih et al., 2016; Talmor & Berant, 2018; Puerto et al., 2023) derived from specific facts of KG.
-Hence, they can only test a very limited subset of the LLM’s tendency to hallucinate, leaving a gap in the assessment of complex, real-world scenarios.
-Particularly, standard metrics such as FActScore (Min et al., 2023) and SAFE (Wei et al., 2024) for evaluating the hallucination rate of LLMs require open-ended settings, i.e., questions are phrased as a statement which requires a longer answer.
-Motivation: Recognizing the research gap, our study aims to contribute to the existing knowledge from two perspectives: 1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and 2) conduct a comparative analysis to shed light on methods and directions for leveraging KGs to reduce LLMs’ hallucination.
-To achieve this, we develop a new benchmark, Open-ended Knowledge-Graphs Question Answering (OKGQA), specifically designed to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios.
-OKGQA is designed to closely reflect the complexities of practical applications using questions of different types, and incorporates specific metrics to measure both the reduction in hallucinations and enhancements in reasoning capabilities.
-Specifically, OKGQA covers 850 queries from ten categories, representing different levels of real-world QA types.
-All queries are synthesized by LLMs with pre-defined templates mentioned in Table 1, and we use Wikipedia as references for all the entities mentioned in the query.
-All the questions are designed to be open-ended and cannot be simply answered by generating logical forms or retrieving specific facts from the knowledge graph.
-To consider the scenarios in which KGs may have varying levels of mistakes (i.e., attributes may be mislabeled, relations may not exist, etc.), we propose another experiment setup OKGQA-P (discussed in Section 3.2) to assess model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated.
-We assess the hallucination ratio and the quality of the response (detailed metrics can be found in Section 3.3) of the tested models.
-Based on our experiments, we find that (1) retrieving KG information can indeed mitigate factual errors in LLMs’ responses, especially for queries requiring additional reasoning; (2) directly performing reasoning in the LLM itself (e.g., internal reasoning strategies like Chain-of-thoughts (Kim et al., 2023) and self-consistency (Wang et al., 2022)) may cause bias and hallucination; (3) Subgraphs generally achieve the best performance across different query types, especially for simpler types of queries; (4) Integrating KGs can effectively reduce hallucinations in LLMs, even when KGs are contaminated to a certain degree.
-Overall, our main contributions include:
-  • We present OKGQA, a benchmark tailored for evaluating LLMs enhanced with KGs under open-ended, real-world question-answering scenarios.
-The benchmark extends the assessment of closed-ended question answering to an open-ended setting, which can further support the assessment of hallucination of LLMs.
-  • To consider the scenarios in which KGs may have varying levels of mistakes, we propose another experiment setup OKGQA-P to assess model performance under conditions in which the semantics and structure of KGs are deliberately perturbed and contaminated.
-  • We conduct a series of experiments on OKGQA and OKGQA-P, analyzing the effectiveness of various retrieval methods and LLMs of different magnitudes, providing insights for future research and development.
-Our results underscore the importance of integrating LLMs with KGs to help reduce hallucinations, even in circumstances where KGs are contaminated.
-2 Related Work
-Due to the stochastic decoding process of Large language models (LLMs), i.e., sampling the next token in the sequence, LLMs exhibit probabilistic behavior, potentially yielding varied outputs of the same input across different instances (Agrawal et al., 2023).
-In addition, they also face challenges in accurately interpreting phrases or terms when the context is vague and resides in a knowledge gap region of the model, leading to outputs that may sound plausible but are often irrelevant or incorrect.
-This “hallucinations" undermines the reliability of LLMs (Huang et al., 2023).
-One emerging research trend is enhancing LLMs through integrating external knowledge graphs (Agrawal et al., 2023).
-KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023).
-Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance.
-This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent.
-Due to their reliability and interpretability, KGs are considered a promising method to improve the reliability of LLM reasoning.
-Researchers employ diverse strategies to augment the LLMs by integrating external KGs.
-For example, KAPING (Baek et al., 2023b) matches entities in questions to retrieve related triples from knowledge graphs for zero-shot question answering.
-Wu et al.
-(2023) finds that converting these triples into textualized statements can further enhance LLM performance.
-StructGPT (Jiang et al., 2023b) propose to convert user query into structured formats (e.g., SPARQL) for information extraction from KGs.
-Following the succuess of internal reasoning-enhancement methods like Chain-of-thoughts (CoT) (Wei et al., 2022), Reflexion (Shinn et al., 2024), and Tree-of-thoughts (ToT), He et al.
-(2022) propose “rethinking with retrieval" to use decomposed reasoning steps from CoT prompting to retrieve external knowledge, leading to more accurate and faithful explanations.
-IR-CoT (Trivedi et al., 2022) interleaves the generation of CoT with knowledge retrieval from corresponding KGs, iteratively guiding both retrieval and reasoning for multi-step questions.
-MindMap (Wen et al., 2023) introduce a plug-and-play approach to evoke graph-of-thoughts reasoning in LLMs.
-Similarly, Reasoning on Graphs (RoG) (Luo et al., 2023b) use KGs to create faithful reasoning paths based on various relations, enabling interpretable and accurate reasoning in LLMs.
-However, current benchmarks for testing the capabilities of these LLM+KG models are predominantly closed-ended, restricting responses to a limited set of entities/relations or a set of logical forms derived from specific facts of KG.
-Hence, they can only test a very limited subset of the LLM’s tendency to hallucinate, leaving a gap in the assessment of complex, real-world scenarios.
-Particularly, standard metrics such as FActScore (Min et al., 2023) and SAFE (Wei et al., 2024) for evaluating the hallucination rate of LLMs require open-ended settings, i.e., questions are phrased as a statement which requires a longer answer.
-Compared with previous works, our proposed OKGQA is tailored for evaluating LLMs enhanced with KGs under open-ended, real-world question-answering scenarios.
-The benchmark extends the assessment of closed-ended question answering to an open-ended setting, which can further support the assessment of hallucination of LLMs.
+Figure 1: Comparison between Open-ended Question and Close-ended Question over Knowledge Graph.
+To address this critical limitation, researchers have turned to leveraging external knowledge graphs (KGs) as a complementary (Pan et al., 2024; Luo et al., 2023a; Hu et al., 2023; Sun et al., 2023; van Breugel et al., 2021).
+KGs provide several key advantages: they contain structured and explicit factual knowledge, including domain-specific information, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023).
+Furthermore, KGs offer traceability, allowing each piece of information to be traced back to its source.
+This traceability enables both verification of generated context and transparency in reasoning paths, making KGs a promising source for improving the reliability of LLM reasoning.
+Find more details of related works in §B.
+However, current benchmarks for testing the capabilities of these LLM+KG models are predominantly closed-ended, restricting responses to a limited set of entities or relations (Talmor et al., 2019; Mihaylov et al., 2018; Jin et al., 2020) or a set of logical forms (Yih et al., 2016; Talmor and Berant, 2018; Puerto et al., 2023) derived from specific facts of KG (as shown in Figure 1).
+These benchmarks, while useful for evaluating basic retrieval and reasoning, are insufficient for testing the LLMs’ ability for faithful generation (i.e., minimize hallucination in LLM responses), leaving a gap in the assessment of complex, real-world scenarios.
+Why Open-ended Question?
+A key limitation of closed-ended benchmarks is: they cannot assess the phenomenon of “more talk, more mistakes”, where longer, more complex responses tend to increase hallucination Qiu et al.
+(2024).
+Unlike closed-ended benchmarks that require short, specific answers, open-ended questions demand more detailed and nuanced explanations.
+This format also enables direct evaluation metrics like FActScore (Min et al., 2023) and SAFE (Wei et al., 2024), which measure factual consistency of LLM responses by decomposing them into atomic statements for comparison with external information.
+However, for short, rigid responses, this decomposition is not feasible, rendering these metrics inapplicable.
+Moreover, open-ended evaluations align more closely with human cognition Lu et al.
+(2024); Amirizaniani et al.
+(2024), where context-rich and interpretative answers are expected compared to simplistic and mechanical responses.
+Find more details of intuition in Appendix §A.1.
+Motivation: Recognizing the research gap, our study aims to contribute to the existing knowledge from two perspectives: 1) explore whether KGs can make LLMs more trustworthy in the open-ended setting, and 2) conduct a comparative analysis to shed light on methods and directions for leveraging KGs to reduce LLMs’ hallucination.
+To achieve this, we develop a new benchmark, Open-ended Knowledge-Graphs Question Answering (OKGQA), specifically designed to assess LLMs enhanced with KGs under open-ended question answering scenarios.
+OKGQA is designed to closely reflect the complexities of practical applications using questions of different types and incorporates specific metrics to measure both the reduction in hallucinations and enhancements in reasoning capabilities.
+Specifically, all queries in OKGQA are synthesized by LLMs with pre-defined templates mentioned in Table 1 and corresponding Wikipedia context.
+To consider the scenarios in which KGs may have varying levels of mistakes (i.e., attributes may be mislabeled, relations may not exist, etc.), we propose another experiment setup OKGQA-P (§2.2) to assess model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated.
+We assess the hallucination ratio and the quality of the response (detailed metrics can be found in §4.1) of the tested models.
+Based on our experiments, we find that (1) retrieving KG information can mitigate factual errors in LLMs’ responses, especially for queries requiring additional reasoning; (2) directly performing reasoning in the LLM itself (e.g., internal reasoning strategies like Chain-of-thoughts (Kim et al., 2023) and self-consistency (Wang et al., 2022)) may cause bias and hallucination; (3) Subgraphs generally achieve the best performance across different query types, especially for simpler types of queries; (4) Integrating KGs can effectively reduce hallucinations in LLMs, even when KGs are contaminated to a certain degree.
 2 OKGQA: A Open-ended Knowledge Graph Question-Answering Benchmark
-OKGQA is a benchmark designed to assess LLMs enhanced with KGs under open-ended, real-world questions answering scenarios.
-OKGQA is designed to closely reflect the complexities of practical applications using questions of different types and incorporates specific metrics to measure both the reduction in hallucinations and enhancement in reasoning capabilities.
-In this section, we discuss the construction of the dataset, and the construction of OKGQA-P which is a variant of OKGQA where the KGs’ semantics and structure are deliberately perturbed and contaminated.
-Motivation of Open-ended QA instead of Close-ended QA: Current benchmarks for testing the capabilities of LLM+KGs models are predominantly close-ended [^1] https://en.wikipedia.org/wiki/Open-ended_question#Examples, which demand a short answer such as ‘yes’ and ‘no’.
-This restricts the responses to a limited set of entities/relations (Talmor et al., 2019; Mihaylov et al., 2018; Jin et al., 2020) or a set of logical forms (Yih et al., 2016; Talmor & Berant, 2018; Puerto et al., 2023) derived from specific KG facts.
-Hence, they can only test a very limited subset of the LLM’s tendency to hallucinate, leaving a gap in the assessment of more complex, real-world scenarios.
-Particularly, standard metrics (like FActScore (Min et al., 2023) and SAFE (Wei et al., 2024)) for evaluating the hallucination rate of LLMs require open-ended settings, i.e., questions are phrased as a statement which requires a longer answer.
-Open-ended QA Task Definition: The tasks we design require to first understand the scope of the question, then optionally retrieve relevant information from multiple parts of the knowledge graph, then finally synthesize a coherent and informative response.
-The ideal output should be a paragraph that fully addresses the question with accurate and factual responses.
-We verify the response based on the metrics specified in Section 5 to reflect the models’ capabilities and faithfulness.
+Overview.
+OKGQA is a comprehensive benchmark designed to assess how effectively LLMs enhanced with KGs perform in open-ended, real-world-like question answering scenarios.
+Unlike existing benchmarks that focus primarily on closed-ended tasks, OKGQA presents diverse open-ended question types that mirror the variable nature of practical applications.
+The benchmark evaluates how effectively models leverage KGs to generate both reliable and comprehensive answers, focusing on reducing hallucinations while maintaining reasoning capabilities of LLMs.
+To assess robustness under imperfect conditions, we also introduce OKGQA-P, a variant where KGs are deliberately perturbed and contaminated.
+2.1 Dataset Construction
 Statistics (on average)
@@
 Figure 2: (left) Dataset statistics and query types, (right) Sub-query type distribution
-3.1 Dataset Construction
 Queries.
-We utilize a template-based methodology to systematically generate a diverse range of queries using LLMs, including categories such as descriptive, explanatory, predictive, comparative, and critical queries.
-Details regarding the specific templates and example queries can be found in Table 1, while the corresponding prompts are provided in the Appendix A.1.
-To ensure that the generated queries accurately represent real-world scenarios and complexities, we integrate the corresponding Wikipedia pages of each entity as contextual information for the LLMs’ generation.
-Furthermore, we follow a human-in-the-loop process to enhance the instructions for generating (the details are given in Appendix A.4).
-Intuitively, starting with an initial instruction, we generate a large number of query candidates, then use an LLM as automatic evaluator to evaluate the quality of the query, denoted as a set of scores from different metrics $s_{\text{auto}}$ ranging from 1 to 10, with higher scores indicating better performance.
-Then, these queries are manually evaluated and assign a set of human-label score $s_{\text{human}}$ correspond to $s_{\text{auto}}$, normalized as the same range as $s_{\text{auto}}$.
-We then optimize the input instructions by iteratively generate instructions to minimize the gap between $s_{\text{human}}$ and $s_{\text{auto}}$.
-These queries are categorized by difficulty and the naturalness of their phrasing.
-The statistics of the queries can be found in Figure 1.
+We utilize a template-based methodology to generate a diverse range of queries using LLMs, including categories such as descriptive, explanatory, predictive, comparative, and critical queries.
+Details regarding specific templates and example queries can be found in Table 1, while the corresponding prompts are provided in the Appendix C.
+To ensure that the generated queries represent real-world scenarios and complexities, we employ an iterative optimization approach that utilizes both automated and human evaluation to refine the query generation process (the details are given in Appendix A.2).
+Initially, we generate a diverse set of query candidates from a seed instruction.
+These candidates undergo automated evaluation using an LLM-based evaluator, which assigns quality scores $s_{\text{auto}}$ on a scale of 1-10, with higher scores indicating better performance across multiple metrics.
+Subsequently, human evaluators assess the same queries, producing corresponding normalized scores $s_{\text{human}}$ within the same range.
+To refine the dataset, we iteratively optimize the input instructions by minimizing the discrepancy between $s_{\text{human}}$ and $s_{\text{auto}}$.
+This optimization process ensures alignment between automated and human quality assessments.
+The queries are also categorized by complexity with detailed statistics in Figure 2.
+KG Sub-graphs.
+To reduce the size of KGs, following previous works (Yih et al., 2016; Talmor and Berant, 2018), we sample subgraphs from DBpedia Knowledge Graph (Noted that all the queries in OKGQA can be reasoned based on DBPedia) by extracting all triples contained within the k-hop neighbors from the question entities in the query.
+We set k = 2 to balance graph size and computational feasibility, as increasing k leads to exponential growth in the number of edges and nodes, which can introduce excessive noise and make it more challenging to retrieve essential information from KGs [^1] This choice is also informed by common practices in other benchmarks, such as WebQSP Yih et al.
+(2016) and CWQ Talmor and Berant (2018), where 2-hop subgraphs are widely used for similar KGQA tasks..
+To further reduce the size of the 2-hop subgraphs, we leverage personalized page-rank (PPR) (Bahmani et al., 2010) to prune the nodes/edges that are not relevant to the query (the details of the PPR algorithm are discussed in Appendix A.3).
+We compare the statistics of subgraphs before and after PPR pruning in Figure 2(a).
+2.2 OKGQA-P: Benchmark with Noise & Perturbations in KGs
+To further mimic the real situations where KGs may not be of high quality (i.e., attributes of nodes/edges may be mislabeled, relations may not exist, etc.), we propose another experiment setting OKGQA-P to assess the model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated.
+We introduce various perturbation scenarios including mislabeled attributes, incorrect relations, and missing connections to test the model’s reasoning capabilities with flawed or incomplete KG data.
+To quantify the degree of perturbation, we also evaluate both the semantic and structural similarity between the original and the modified KG as defined below.
+Notation.
+Let $\mathcal{F}_{\theta}$ be a KG-augmented model, and KG as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T})$, where $\mathcal{V}$ is the set of entities (nodes), $\mathcal{E}$ is the set of relation types (edges), and $\mathcal{T}=\{(v_{1},e,v_{2})|v_{1},v_{2}\in\mathcal{V},e\in\mathcal{E}\}$ is the set of triplets composed of entities and relations.
+Let $\mathcal{G}^{\prime}=(\mathcal{V},\mathcal{E}^{\prime},\mathcal{T}^{\prime})$ be the KG after perturbing $\mathcal{G}$, where $\mathcal{E}^{\prime}\neq\mathcal{E}$ and $\mathcal{T}^{\prime}\neq\mathcal{T}$.
+Let $f(\mathcal{G},\mathcal{G}^{\prime})$ be a function that measures the similarity between $\mathcal{G}$ and $\mathcal{G}^{\prime}$.
+Let $g(\mathcal{G})$ be the downstream performance when evaluating $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}$.
+High-level Procedure.
+First, we test $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}$ to get the original performance $g(\mathcal{G})$.
+Second, we perturb $\mathcal{G}$ to obtain $\mathcal{G}^{\prime}$.
+Third, we evaluate $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}^{\prime}$ to get the perturbed performance $g(\mathcal{G}^{\prime})$.
+Finally, we measure $g(\mathcal{G})-g(\mathcal{G}^{\prime})$ and $f(\mathcal{G},\mathcal{G}^{\prime})$ to assess how robust $\mathcal{F}_{\theta}$ is, i.e., to assess the model performance under conditions where KGs’ semantics and structure are deliberately perturbed.
+To quantify how much the perturbed KG has deviated from the original KG, i.e., $f(\mathcal{G},\mathcal{G}^{\prime})$, we leverage metrics from Raman et al.
+(2020) for capturing semantics (ATS) and structural (SC2D, SD2) similarity between KGs.
+Intuitively, ATS leverages a pre-trained LM for link prediction to measure the probability of each edge from $\mathcal{G}^{\prime}$ existing in $\mathcal{G}$, while SC2D and SD2 measure the structural similarity between two KGs based on local clustering coefficient and degree distribution.
+For each of the three metrics, higher value indicates higher similarity.
+The detailed description can be found in Appendix A.5, with visualization in Figure 5.
+For the perturbation methods, we consider four perturbation heuristics based on (Raman et al., 2020) as follows: Relation Swapping (RS) randomly chooses two edges from $\mathcal{T}$ and swaps their relations.
+Relation Replacement (RR) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, and replaces the $e_{1}$ with another relation $e_{2}=\text{argmin}_{e\in\mathcal{E}}S_{\mathcal{G}}(v_{1},e,v_{2})$, where $S_{\mathcal{G}}(\cdot)$ is a KG score function adapted from ATS.
+The replacement is performed such that the resulting $(v_{1},e_{2},v_{2})$ has the closest score to the original triplet $(v_{1},e_{1},v_{2})$ to generate “harder negatives” - triplets that are semantically similar but incorrect, thereby challenging the model to different subtle distinctions.
+Noted that $S_{G}$ differs from the original ATS, as it operates on individual triplets rather than comparing two entire KGs.
+Edge Rewiring (ER) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, then replaces $v_{2}$ with another entity $v_{3}\in\mathcal{E}\backslash\mathcal{N}_{1}(v_{1})$, where $\mathcal{N}_{1}(v_{1})$ represents the 1-hop neighborhood of $v_{1}$.
+Edge Deletion (ED) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$ and deletes it.
+We control the perturbation level based on the percentage of KG edges perturbed.
+See results in Figures 5 and 6.
 Table 1: Query types and examples in OKGQA.
-Brown is used to highlight the placeholders (e.g., [person], [event]) in description, while Teal refers to the specific entities to replace the placeholders.
-The distribution of the sub-query types can be referred to in Figure 1(b).
+Brown is used to highlight the placeholders (e.g., [person], [event]) in description, while Teal highlights the specific entities to replace the placeholders.
 Type Sub-Type Description / Template Example
@@
 applied to medical diagnostics?
-KG Sub-graphs.
-To reduce the size of KGs, following previous works (Luo et al., 2023a; Yih et al., 2016; Talmor & Berant, 2018), we construct subgraphs of DBpedia (all the queries in OKGQA can be reasoned based on DBPedia [^2] https://www.dbpedia.org/resources/knowledge-graphs/) by extracting all triples contained within the k-hop neighbors from the question entities in the query.
-For our experiments, we set k=2.
-To further reduce the size of the 2-hop subgraphs, we leverage personalized page-rank (PPR) (Bahmani et al., 2010) to prune the nodes/edges that are not relevant to the query (the details of the PPR algorithm are discussed in Appendix A.5).
-We compare the statistics of subgraphs before and after PPR pruning in Figure 1(a).
-3.2 OKGQA-P: Benchmark with Noise & Perturbations in KGs
-To further mimic the real situations where KGs may not be of high quality (i.e., attributes of nodes/edges may be mislabeled, relations may not exist, etc.), we propose another experiment setting OKGQA-P in this section to assess the model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated.
-Considering that KGs are typically annotated by humans and are generally accurate and meaningful, we introduce perturbations to edges in the KG to degrade the quality of the KGs, diminishing human comprehensibility.
-To quantify the degree of perturbation, we evaluate both the semantic and structual similarity between the original and the modified KG as defined below.
-Notation.
-Let $\mathcal{F}_{\theta}$ be a KG-augmented model, and KG as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T})$, where $\mathcal{V}$ is the set of entities (nodes), $\mathcal{E}$ is the set of relation types (edges), and $\mathcal{T}=\{(v_{1},e,v_{2})|v_{1},v_{2}\in\mathcal{V},e\in\mathcal{E}\}$ is the set of triplets composed from existing entities and relations.
-Let $\mathcal{G}^{\prime}=(\mathcal{V},\mathcal{E^{\prime}},\mathcal{T^{\prime}})$ be the KG obtained after perturbing $\mathcal{G}$, where $\mathcal{E}^{\prime}\neq\mathcal{E}$ and $\mathcal{T}^{\prime}\neq\mathcal{T}$.
-Let $f(\mathcal{G},\mathcal{G}^{\prime})$ be a function that measures the similarity between $\mathcal{G}$ and $\mathcal{G}^{\prime}$.
-Let $g(\mathcal{G})$ be the downstream performance when evaluating $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}$.
-High-level Procedure.
-First, we test $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}$ to get the original performance $g(\mathcal{G})$.
-Second, we perturb $\mathcal{G}$ to obtain $\mathcal{G}^{\prime}$.
-Third, we evaluate $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}^{\prime}$ to get the perturbed performance $g(\mathcal{G}^{\prime})$.
-Finally, we measure $g(\mathcal{G})-g(\mathcal{G}^{\prime})$ and $f(\mathcal{G},\mathcal{G}^{\prime})$ to assess how robust $\mathcal{F}_{\theta}$ is, i.e., to assess the model performance under conditions where KGs’ semantics and structure are deliberately perturbed.
-To quantify how much the perturbed KG has deviated from the original KG, i.e., $f(\mathcal{G},\mathcal{G}^{\prime})$, we leverage metrics from Raman et al.
-(2020) for capturing semantics (ATS) and structural (SC2D, SD2) similarity between KGs.
-Intuitively, ATS leverages an pre-trained LM for link prediction to measure the probability of each edge from $\mathcal{G}^{\prime}$ existing in $\mathcal{G}$, while SC2D and SD2 measures the structural similarity between two KGs based on local clustering coefficient and degree distribution.
-For each of the three metrics, higher value indicates higher similarity.
-The detailed description can be found in Appendix A.7.
-The results after perturbation can be found in Figure 4.
-For the perturbation methods, we consider four perturbation heuristics based on (Raman et al., 2020) as follows: Relation Swapping (RS) randomly chooses two edges from $\mathcal{T}$ and swaps their relations.
-Relation Replacement (RR) randomly chooses an edge $v_{1},e,v_{2})\in\mathcal{T}$, then replaces $e_{1}$ with another relation $e_{2}=\text{argmin}_{r\in\mathcal{R}}S_{\mathcal{G}}(v_{1},e,v_{2})$, where $S_{\mathcal{G}}(v_{1},e,v_{2})$ uses ATS to measure the semantics similarity between two edges.
-Edge Rewiring (ER) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, then replaces $v_{2}$ with another entity $v_{3}\in\mathcal{E}\backslash\mathcal{N}_{1}(v_{1})$, where $\mathcal{N}_{1}(v_{1})$ represents the 1-hop neighborhood of $v_{1}$.
-Edge Deletion (ED) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$ and deletes it.
-We control the perturbation level based on the percentage of KG edges being perturbed.
+3 Exploring KG-augmented framework for reducing hallucination
+To explore whether KG-augmented approaches can mitigate LLMs’ hallucination, we investigate a range of methods and discover that they can be encapsulated within a unified framework as shown in Figure 3.
+Our framework follows the paradigm of retrieval augmented generation (RAG) (Edge et al., 2024; Baek et al., 2023a), which retrieves essential information from the KGs, and then uses the retrieved knowledge to enhance the LLM’s generation.
+It has two components, i.e., Graph-guided retrieval (G-retrieval) and Graph-guided generation (G-generator), with a variety of algorithmic design choices.
+We analyze our proposed strategies within each component in §4, aiming to shed light on the best practices for leveraging KGs for reducing hallucinations in LLMs.
+Figure 3: Overview of KG-augmented framework.
+Here we formalize the problem of context-relevant KG retrieval for open-ended question answering.
+Given a user query $q$, a pretrained language model (PLM) first encodes the input tokens, and models a conditional distribution $p(a|q)$ to generate an output response $a=[a_{1},\ldots,a_{T}]$.
+To explore whether KGs help reduce hallucinations of LLMs, we introduce the retrieved knowledge from the KG, denoted $\mathcal{Z}\in\mathcal{G}$, which can take various forms (e.g., triplets, paths, subgraphs, as shown in Figure 3).
+This retrieved knowledge contains information relevant to the user query, and either benefits the LLMs at its knowledge boundaries (Baek et al., 2023b) or explicitly guides its reasoning process (Luo et al., 2023a; Sun et al., 2023).
+The distribution of the retrieved knowledge $\mathcal{Z}$ is modeled as $p(\mathcal{Z}|q,\mathcal{G})$, leading to a new likelihood for generating responses, expressed as $p(a|q)$.
+$p(a|q)$ $=\sum_{\mathcal{Z}\subseteq\mathcal{G}}p_{\phi}(a|q,\mathcal{Z})p_{\theta}(\mathcal{Z}|q,\mathcal{G})$ (1)
+$\approx p_{\phi}(a|q,\mathcal{Z}^{*})p_{\theta}(\mathcal{Z}^{*}|q,\mathcal{G}),$
+where $\mathcal{Z}^{*}=\text{argmax}_{\mathcal{Z}\in\mathcal{G}}p_{\theta}(\mathcal{Z}|q,\mathcal{G})$ represents the optimal subset of retrieved knowledge from the KGs, and $p_{\phi}(a|q,\mathcal{Z}^{*})$ represents the likelihood generating the response given the query $q$ and the retrieved knowledge $\mathcal{Z}^{*}$, parameterized by $\phi$ and $\theta$.
+Due to the potential exponential growth in the number of possible retrieved knowledge with graph size, efficient approximation methods are required.
+Consequently, the first line of Eq.
+1 is approximated using the most probable retrieved knowledge $\mathcal{Z}^{*}$.
+3.1 Graph-guided retrieval (G-retrieval)
+In this section, we explore which forms of knowledge (e.g., triplets, paths, or subgraphs) should be extracted from KGs to enhance the LLMs and discuss the design of G-retrieval.
+The phase begins by encoding the user query into an embedding vector $\mathbf{q}\in\mathbb{R}^{d}$ using a language model encoder $f_{\text{LM}}(\cdot)$.
+For each node $v\in\mathcal{V}$ and edge $e\in\mathcal{E}$, their respective embeddings $\mathbf{e}_{v}$ and $\mathbf{e}_{e}$ are similarly computed using the same model, ensuring all the information is represented in a unified vector space.
+The relevance of each node or edge to the query is measured using cosine similarity as: $s_{v}=\frac{\mathbf{q}\cdot\mathbf{e}_{v}}{\|\mathbf{q}\|\|\mathbf{e}_{v}\|}$ and $s_{e}=\frac{\mathbf{q}\cdot\mathbf{e}_{e}}{\|\mathbf{q}\|\|\mathbf{e}_{e}\|}$.
+Based on these similarity scores, the most relevant nodes and edges are identified to form a set of top-$k$ nodes and edges for the query.
+To balance retrieving as many relevant nodes and edges as possible while keeping the size manageable, we adopt a prize-cost trade-off strategy: (1) Prize assignment: based on the computed similarity scores, we assign prizes to nodes and edges to quantify their relevance to the query.
+Specifically, we assign the top-$k$ nodes/edges with descending prize values from $k$ to 1, while nodes and edges outside the top-$k$ receive a prize of 0.
+Formally: $p_{v}=\max(0,k-\text{rank}(v)+1)$ and $p_{e}=\max(0,k-\text{rank}(e)+1)$.
+(2) Cost allocation: to manage the retrieved knowledge size, we assign penalties as cost $C_{e}$ during the expansion of the retrieved paths or subgraphs.
+These prizes and costs will be used in the retrieval process, by aiming to maximize the prizes while minimizing costs.
+Triplet-retrieval: the triplet retrieval approach retrieves a fixed number $k$ of triplets with the highest total prize assigned to their respective head entity, relation, and tail entities.
+Path-retrieval: based on the defined prize assignment and cost allocation, the path-retrieval approach starts from a fixed number of $k$ of high-prize nodes to construct sequences $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$ by aiming to greedily maximize their scores: $S(\mathcal{P})=\sum_{i=1}^{n}p_{v_{i}}+\sum_{i=1}^{n-1}p_{e_{i}}-\sum_{i=1}^{n-1}c_{e}$.
+We use a priority queue to iteratively extend paths by selecting the highest-scoring extensions while avoiding cycles and adhering to maximum path length and quantity constraints.
+Finally, the top-scoring paths are sorted and returned.
+The details of path-retrieval can be found in Appendix A.4.
+Sub-graph retrieval: the sub-graph retrieval approach also uses the prize assignment and cost allocation, following He et al.
+(2024a) to optimize the process using Prize-Collecting Steiner Tree (PCST) algorithm.
+The PCST problem seeks to build a connected subgraph $\mathcal{S}$ that maximizes the total prizes of included nodes and edges while minimizing the total cost of the edges.
+The score of the subgraph is defined as $S(\mathcal{S})=\sum_{n\in V_{\mathcal{S}}}p_{v_{i}}+\sum_{e\in E_{\mathcal{S}}}p_{e_{i}}-\sum_{einE_{\mathcal{S}}}c_{e}$.
+Unlike in path-retrieval, we only yield one subgraph that maximizes the total score.
+3.2 Graph-guided generation (G-generator)
+After retrieving $\mathcal{Z}^{*}$, the graph-guided generation phase learns to use this knowledge to generate the corresponding output to answer the user query.
+The generation is modeled as a sequential decision-making process, where at each time step $t$, the next token $a_{t}$ is generated by G-Generator, conditioned on the query $q$, the retrieved knowledge $\mathcal{Z}^{*}$, and all previously generated tokens $a_{0:t-1}$.
+$p(a|q,\mathcal{Z}^{*})=\prod_{t=1}^{T}p_{\theta}\left(a_{t}|q,\mathcal{Z}^{*},a_{0:t-1}\right)\vspace{-2mm}$ (2)
+The model learns the parameters $\theta$ to maximize the likelihood of generating the correct output sequence by computing the probability distribution over possible next tokens at each time step.
+This ensures that the generated output not only considers the previously generated tokens but also the retrieved knowledge.
+The generation continues until a termination condition is met, such as generating an end-of-sequence token or reaching the maximum sequence length $T$.
+4 Experiments
 4.1 Evaluation Metrics
@@
 It employs a language model as an investigative agent, which iteratively uses Google Search queries and reasons about whether the search results support or do not support the fact.
-In addition to the hallucination metrics, we have also included four metrics, following previous works (Edge et al., 2024; Wang et al., 2023) to quantify the desirable qualities for generating sensible responses.
-Given the complexities inherent in our open-ended QA setting, collecting and annotating ground-truth answers pose significant challenges.
-Therefore, we leverage LLMs as automatic reference-free evaluators to assess these metrics based on G-Eval (Liu et al., 2023), a framework of using LLMs with chain-of-thought (CoT) and a form-filling paradigm.
-To enhance G-Eval’s robustness and stability, we provide the relevant Wikipedia pages as context in the prompt when calculating the metrics.
-The metrics are defined as follows: Context Relevance: measures how well the generated response aligns with the provided context.
-Comprehensiveness: evaluates the extent to which the generated answer covers all aspects and details of the question.
-Correctness: measures how clearly and specifically the generated answer responds to the question.
-Empowerment: evaluates how well the generated answer helps the reader understand the topic and make informed decisions.
-The detailed prompt for each metric can be found in Appendix A.1.
-Models.
-We incorporate a range of widely used LLMs of different scales, including GPT-4o [^3] https://openai.com/index/hello-gpt-4o/, GPT-4o-mini [^4] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, Llama-3.1-8B-instruct (Dubey et al., 2024), Mistral-7B-instruct-v0.3 (Jiang et al., 2023a), and Gemma-2-9B-it (Team et al., 2024).
-We use the embedding model (text-embedding-3-small) from OpenAI for the G-retrieval process, considering cost and performance.
-4 Exploring KG-augmented framework for reducing hallucination
-Figure 2: Overview of KG-augmented framework for reducing hallucinations.
-Overview.
-To explore whether KG-augmented approaches can mitigate LLMs’ hallucination, we investigated a range of methods and discovered that they can be encapsulated within a unified framework as shown in Figure 2.
-Our framework follows the paradigm of retrieval augmented generation (RAG) (Edge et al., 2024; Baek et al., 2023a), which retrieves essential information from the KGs, and then uses the retrieved knowledge to enhance the LLM’s generation.
-It has two components, i.e., Graph-guided retrieval (G-retrieval) and Graph-guided generation (G-generator), with a variety of algorithmic choices that combine to create diverse algorithms with different benefits for reducing hallucinations.
-We analyze our proposed strategies within each component in Section 5, aiming to shed light on the best practices for leveraging KGs for reducing hallucinations in LLMs.
-Here we formalize the problem of context-relevant KG retrieval for open-ended question answering.
-Given a user query $q$, a generative pretrained language models (PLMs) first encodes the input tokens, and then models a conditional distribution $p(a|q)$ to generate an output response $a=[a_{1},\ldots,a_{T}]$.
-To explore whether KGs can help reduce hallucinations of PLMs, we introduce the retrieved knowledge from the KG, denoted $\mathcal{Z}\in\mathcal{G}$, which can take various forms (e.g., triplets, paths, subgraphs, as shown in Figure 2).
-This retrieved knowledge contains information relevant to the user query, and can either benefit the PLM at its knowledge boundaries (Baek et al., 2023b) or explicitly guide its reasoning process (Luo et al., 2023a; Sun et al., 2023).
-We explore which forms of knowledge should be extracted from KGs to best enhance the PLM.
-The distribution of the retrieved knowledge $\mathcal{Z}$ is modeled as $p(\mathcal{Z}|q,\mathcal{G})$, leading to a new likelihood for generating responses, expressed as $p(a|q)$.
-$p(a|q)$ $=\sum_{\mathcal{Z}\subseteq\mathcal{G}}p_{\phi}(a|q,\mathcal{Z})p_{\theta}(\mathcal{Z}|q,\mathcal{G})$ (1)
-$\approx p_{\phi}(a|q,\mathcal{Z}^{*})p_{\theta}(\mathcal{Z}^{*}|q,\mathcal{G}),$
-where $p_{\theta}(\mathcal{Z}^{*}|q,\mathcal{G})$ represents the distribution of retrieving relevant knowledge elements from the KGs, and $p_{\phi}(a|q,\mathcal{Z}^{*})$ represents the distribution of leveraging the retrieved knowledge elements to generate the final responses, parameterized by $\phi$ and $\theta$, where $\mathcal{Z}^{*}$ is the optimal retrieved knowledge.
-Due to the potential exponential growth in the number of possible retrieved knowledge with graph size, efficient approximation methods are required.
-Consequently, the first line of Eq.
-1 is approximated using the most probable retrieved knowledge $\mathcal{Z}^{*}$.
-4.1 Graph-guided retrieval (G-retrieval)
-The graph-guided retrieval phase focuses on extracting the most relevant elements (e.g., triplets, paths or subgraphs) from KGs to help answer the user query.
-Initially, the query is transformed into an embedding vector $\mathbf{q}\in\mathbb{R}^{d}$ using a language model encoder $f_{\text{LM}}(\cdot)$.
-For each node $v\in\mathcal{V}$ and edge $e\in\mathcal{E}$, their respective embeddings $\mathbf{e}_{v}$ and $\mathbf{e}_{e}$ are computed using the same language model to map all the information into the same vector space, and their relevance to the query is quantified using cosine similarity as: $s_{v}=\frac{\mathbf{q}\cdot\mathbf{e}_{v}}{\|\mathbf{q}\|\|\mathbf{e}_{v}\|}$ and $s_{e}=\frac{\mathbf{q}\cdot\mathbf{e}_{e}}{\|\mathbf{q}\|\|\mathbf{e}_{e}\|}$.
-Next, we identify the most relevant nodes and edges for the query based on the corresponding similarity, to yield a set of ‘top-$k$ nodes/edges’.
-The goal of G-retrieval is to retrieve the knowledge that encompasses as many relevant nodes and edges as possible, while keeping the size manageable.
-To this end, we leverage a ‘prize’ and ‘cost’ trade-off strategy to guide the retrieval process, as follows: (1) Prize assignment: based on the computed similarity scores, we assign prizes to nodes and edges to quantify their relevance to the query.
-Specifically, we assign the top-$k$ nodes/edges with descending prizes values from $k$ to 1, while nodes and edges outside the top-$k$ receive a prize of 0.
-Formally: $p_{v}=\max(0,k-\text{rank}(v)+1)$ and $p_{e}=\max(0,k-\text{rank}(e)+1)$.
-(2) Cost allocation: to manage the retrieved knowledge size, we assign penalties as cost $C_{e}$ (by default assign to 1) during the expansion of the retrieved paths or subgraphs.
-These prizes and costs will be used in the retrieval process, by aiming to maximize the prizes while minimizing costs.
-Triplet-retrieval: the triplet retrieval approach retrieves a fixed number $k$ of triplets with the highest total prize assigned to their respective head entity, relation, and tail entities.
-Path-retrieval: based on the defined prize assignment and cost allocation, the path-retrieval approach starts from a fix number of $k$ of high-prize nodes to construct sequences $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$ by aiming to greedily maximize their scores: $S(\mathcal{P})=\sum_{i=1}^{n}p_{v_{i}}+\sum_{i=1}^{n-1}p_{e_{i}}-\sum_{i=1}^{n-1}c_{e}$.
-We use a priority queue to iteratively extend paths by selecting the highest-scoring extensions while avoiding cycles and adhering to maximum path length and quantity constraints.
-Finally, the top-scoring paths are sorted and returned.
-The details of path-retrieval can be found in Appendix A.6.
-Sub-graph retrieval: the sub-graph retrieval approach also leverages the prize assignment and cost allocation; we follow He et al.
-(2024) to optimize the process using Prize-Collecting Steiner Tree (PCST) algorithm.
-The PCST problem seeks to build a connected subgraph $\mathcal{S}$ that maximizes the total prizes of included nodes and edges while minimizing the total cost of the edges.
-The score of the subgraph is defined as $S(\mathcal{S})=\sum_{n\in V_{\mathcal{S}}}p_{v_{i}}+\sum_{e\in E_{\mathcal{S}}}p_{e_{i}}-\sum_{e\in E_{\mathcal{S}}}c_{e}$.
-Unlike in path-retrieval, we only yield one subgraph that maximizes the total score.
-4.2 Graph-guided generation (G-generator)
-After retrieving $\mathcal{Z}^{*}$, the graph-guided generation phase focuses on learning to leverage this knowledge to generate the corresponding output to answer the user query.
-The generation is modeled as a sequential decision-making process as follows, where at each time step $t$, the next token $a_{t}$ is generated by G-Generator, conditioned on the query $q$, the retrieved knowledge $\mathcal{Z}^{*}$, and all previously generated tokens $a_{0:t-1}$.
-$p(a|q,\mathcal{Z}^{*})=\prod_{t=1}^{T}p_{\theta}\left(a_{t}|q,\mathcal{Z}^{*},a_{0:t-1}\right)$ (2)
-The model learns the parameters $\theta$ to maximize the likelihood of generating the correct output sequence by computing the probability distribution over possible next tokens at each time step.
-This ensures that the generated output not only considers the previously generated tokens but also the retrieved knowledge.
-The generation continues until a termination condition is met, such as generating an end-of-sequence token or reaching the maximum sequence length $T$.
-5 Experiments
-5.1 RQ1: Main Results - Can KGs Reduce Hallucination in LLMs?
-To explore whether KGs can help reduce hallucination in LLMs, we benchmark the LLMs in different settings.
-We use zero-shot and few-shot prompting as baselines without injecting external knowledge.
-In addition, compared to adding external knowledge from KGs, we also consider leveraging LLMs’ internal knowledge to do Chain-of-thought (Kim et al., 2023), or self-consistency (Wang et al., 2022) prompting as baselines.
-As for LLMs augmented with KGs, we consider three KG retrieval forms: triplets, paths, and subgraphs (as shown in Figure 2) for comparison, to study the impact of G-retrieval for reducing LLMs’ hallucinations.
-The results can be found in Table 2 and Figure 3.
-We obtain some intriguing findings:
+Figure 4: Comparison results of different forms of information over different queries.
 Table 2: Comparison results of various forms of information extracted from the KGs.
@@
 Var-3: With Paths Extracted from KGs Provided
-GPT-4o $78.71\%\pm 0.53\%$ $74.53\%\pm 0.31\%$ $69.42\%\pm 0.23\%$ $71.63\%\pm 0.61\%$ $90.20\%\pm 0.59\%$ $75.61\%\pm 0.51\%$
+[…]
 GPT-4o-mini $73.64\%\pm 0.93\%$ $69.41\%\pm 0.22\%$ $64.35\%\pm 0.72\%$ $67.52\%\pm 0.82\%$ $88.22\%\pm 0.34\%$ $70.53\%\pm 0.24\%$
@@
 Var-4: With Subgraphs Extracted from KGs Provided
-GPT-4o $80.81\%\pm 0.43\%$ $76.63\%\pm 0.65\%$ $71.57\%\pm 0.51\%$ $73.70\%\pm 0.62\%$ $90.83\%\pm 0.63\%$ $75.33\%\pm 0.29\%$
+[…]
 GPT-4o-mini $75.70\%\pm 0.44\%$ $71.51\%\pm 0.83\%$ $66.43\%\pm 0.76\%$ $69.60\%\pm 0.65\%$ $88.71\%\pm 0.72\%$ $70.12\%\pm 0.87\%$
@@
 gemma-2-9b-it $65.45\%\pm 0.95\%$ $61.23\%\pm 1.0\%$ $56.31\%\pm 0.35\%$ $59.40\%\pm 0.85\%$ $84.51\%\pm 0.99\%$ $63.74\%\pm 0.49\%$
-GPT-4o - CoT+SC $82.90\%\pm 0.57\%$ $78.72\%\pm 0.61\%$ $73.64\%\pm 0.43\%$ $75.80\%\pm 0.75\%$ $89.12\%\pm 0.94\%$ $75.42\%\pm 1.31\%$
+[…]
+In addition to the hallucination metrics, we also propose four metrics to quantify the desirable qualities for generating sensible responses (Edge et al., 2024; Wang et al., 2023).
+Given the complexities inherent in an open-ended setting, collecting and annotating ground-truth answers poses significant challenges.
+Therefore, we leverage LLMs as automatic reference-free evaluators to assess these metrics based on G-Eval (Liu et al., 2023), a framework of using LLMs with chain-of-thought (CoT) and a form-filling paradigm.
+To enhance G-Eval’s robustness and stability, we provide the relevant Wikipedia pages as context in the prompt when calculating the metrics.
+The metrics are defined as follows: Context Relevance: measures how well the generated response aligns with the provided context.
+Comprehensiveness: assesses how thoroughly the answer addresses all aspects and details of the question.
+Correctness: measures the clarity and specificity of the generated answer to the question.
+Empowerment: evaluates how well the generated answer helps the reader understand the topic and make informed decisions.
+The detailed prompt can be found in Appendix C.
+Models.
+We incorporate a range of widely used LLMs of different scales, including GPT-4o, GPT-4o-mini, Llama-3.1-8B-instruct (Dubey et al., 2024), Mistral-7B-instruct-v0.3 (Jiang et al., 2023a), and Gemma-2-9B-it (Team et al., 2024).
+We use the embedding model (text-embedding-3-small) from OpenAI for the G-retrieval process, considering cost and performance.
+4.2 RQ1: Main Results - Can KGs Reduce Hallucination in LLMs?
+To explore whether KGs can help reduce hallucination in LLMs, we benchmark the LLMs in different settings.
+We use zero-shot and few-shot prompting as baselines without injecting external knowledge.
+In addition, compared to adding external knowledge from KGs, we also consider leveraging LLMs’ internal knowledge to do Chain-of-thought (Kim et al., 2023), or self-consistency (Wang et al., 2022) as baselines.
+For LLMs augmented with KGs, we consider three KG retrieval forms: triplets, paths, and subgraphs (as shown in Figure 3) to study the impact of G-retrieval for reducing LLMs’ hallucinations.
+The results are shown in Table 2 and Figure 4.
+We obtain some intriguing findings:
 Retrieving KG information can indeed mitigate factual errors in the responses.
@@
 Moreover, these methods can be combined with strategies like CoT+SC, enhancing response quality with minimal increase in hallucination ratio.
-The radar chart in Figure 3 further emphasizes that in most query types, integrating knowledge retrieved from KGs mitigates the hallucination issue compared to baselines, particularly in query types such as “Evaluation and Reflection," “Outcome Prediction," and “Cause Explanation," which require more reasoning and analysis rather than merely listing information.
-The findings also apply to open-source models like mistral-7B-Instruct-v0.3 and Llama-3.1-8B-instruct, illustrating the consistency of reliability of integrating KGs in LLM reasoning.
-Figure 3: Comparison results of different forms of information over different queries.
+The radar chart in Figure 4 further emphasizes that in most query types, integrating knowledge retrieved from KGs mitigates the hallucination issue compared to baselines, particularly in query types such as “Evaluation and Reflection,” “Outcome Prediction,” and “Cause Explanation,” which require more reasoning and analysis rather than merely listing information.
+The findings also apply to open-source models like mistral-7B-Instruct-v0.3 and Llama-3.1-8B-instruct, illustrating the consistency of the finding.
 Directly performing reasoning in the LLM itself does not mitigate hallucinations.
 We benchmark the hallucination ratio of LLMs using internal reasoning strategies like Chain-of-thought and Self-consistency, as shown in Var-1 in Table 2.
-It shows that these methods can indeed improve response quality (i.e., correctness, comprehensiveness and empowerment) compared to baselines, but do not consistently improve factuality, and sometimes even diminish factuality.
-This showcases that relying solely on internal reasoning does not mitigate hallucination, emphasizing the importance of external knowledge to counter the hallucination issue.
-Subgraph retrieval generally achieves best performance across different query types, especially for simpler queries.
-We demonstrate the performance of different retrieval methods across different query types in Figure 3, showing that subgraphs generally achieve best performance.
-Especially for simpler queries (“Character Description" and “Event Description" which generally don’t require intensive reasoning), subgraph retrieval methods achieves better performance compared to paths and triplets.
-Even for queries like “Relationship Explanation" and “Cause Explanation" which require step-wise reasoning, subgraph methods still demonstrate promising performance.
-This suggests that while different forms of retrieved knowledge each offer unique benefits suitable for specific types of queries, subgraphs provide consistently good performance broadly.
-5.2 RQ2: How Are KG-Aware Methods Affected by Noise / Perturbations in KGs?
-We benchmark different KG-augmented LLMs on our OKGQA-P setting, where we deliberately perturb and contaminate the semantics and structure of KGs to simulate the real-world situation where KGs may not have high quality.
-Specifically, we consider different perturbation methods discussed in Section 3.2 and control the perturbation level based on the percentage of KG edges being perturbed.
-We first illustrate how much the perturbed KG has been deviated from the original KG with the increase of perturbation level, shown in Figure 4.
+It shows that these methods can improve response quality (i.e., correctness, comprehensiveness, and empowerment) compared to baselines, but do not consistently improve factuality, and sometimes even diminish factuality.
+This shows that relying solely on internal reasoning does not mitigate hallucination, emphasizing the importance of external knowledge to counter the hallucination issue.
 (a) Edge Deletion
@@
 Figure 5: Performance Metrics (ATS, SC2D, SD2) vs. Perturbation Level for Different Perturbation Methods.
-It shows that the perturbation methods like edge deletion, rewiring and swapping have relatively weak influence on ATS (which intuitively measures semantic similarity), even as the perturbation level increases.
-For the edge deletion methods, only if the perturbation level reaches 1.0, the ATS goes to 0, otherwise, the ATS is very high.
-Relation replacement has a relatively strong impact on ATS, i.e., on the semantics of the perturbed KGs.
-For SC2D and SD2 (which intuitively measure structural similarity), the four methods show similar levels of decrease in most cases.
 (a) Edge Deletion
@@
 FS-T refers to FActScore metric using triplets, FS-P refers to using paths, and FS-SG refers to using sub-graphs.
+Subgraph retrieval generally achieves best performance across different query types, especially for simpler queries.
+We demonstrate the performance of different retrieval methods across different query types in Figure 4, showing that subgraphs achieve the best performance.
+Especially for simpler queries (“Character Description” and “Event Description” which do not require intensive reasoning), subgraph retrieval methods achieve better performance compared to paths and triplets.
+Even for queries like “Relationship Explanation” and “Cause Explanation” which require stepwise reasoning, subgraph methods still demonstrate promising performance.
+This suggests that while different forms of retrieved knowledge offer unique benefits for specific types of queries, subgraphs provide consistently good performance.
+4.3 RQ2: How Are KG-Aware Methods Affected by Noise / Perturbations in KGs?
+We benchmark different KG-augmented LLMs on our OKGQA-P setting, where we deliberately perturb and contaminate the semantics and structure of KGs to simulate the real-world situation where KGs may not have high quality.
+Specifically, we consider different perturbation methods discussed in §2.2 and control the perturbation level based on the percentage of KG edges being perturbed.
+We first illustrate how much the perturbed KG has been deviated from the original KG with the increase of perturbation level, shown in Figure 5.
+It shows that the perturbation methods like edge deletion, rewiring and swapping have relatively weak influence on ATS (which intuitively measures semantic similarity), even as the perturbation level increases.
+For the edge deletion methods, only if the perturbation level reaches 1.0, the ATS goes to 0, otherwise, the ATS remains higher compared to other settings.
+Relation replacement has a relatively stronger impact on ATS, i.e., on the semantics of the perturbed KGs.
+For SC2D and SD2 (which intuitively measure structural similarity), the four methods show similar levels of decrease in most cases.
 Figure 6 illustrates the hallucination ratio using different methods on the perturbed KGs.
@@
 However, it still outperforms the baseline when the perturbation level is smaller than 40%.
-In conclusion, our study demonstrates that leveraging external knowledge from KGs can effectively reduce hallucinations in LLMs, even when KGs are contaminated to a certain degree.
-In real-world applications, KGs such as Wikidata and unlikely to experience severe perturbations due to their continuous updates and community-driven quality control, making our findings applicable in most real-world scenarios.
+Our findings on OKGQA-P illustrate that the effectiveness of KG-derived information diminishes with a perturbation level at 50%, surpassing this level leads to a further decrease in performance.
+We think that before this perturbation level at 50%, incorporating external knowledge from KGs can mitigate hallucinations in LLMs compared to baseline using CoT.
+In practical scenarios, platforms like Wikidata are less likely to be severely perturbed than 50% due to ongoing updates and community-based quality control, ensuring the relevance of our findings in real-world applications.
 5 Conclusion
-In this paper, we present OKGQA, a benchmark tailored for evaluating LLMs enhanced with KGs under open-ended, real-world question answering scenarios.
-The benchmark extends the assessment from close-ended question answering to the open-ended setting, to support the assessment of hallucination in LLMs.
-To further mimic real-world scenarios where KGs may not have high quality, we propose another experiment setup OKGQA-P to assess model performance under conditions where KG’s semantics and structure are deliberately perturbed and contaminated.
-We conduct a series of experiments on OKGQA and OKGQA-P, analyzing the effectiveness of various retrieval methods and LLMs of different magnitudes, providing insights for further research and development.
+In this paper, we propose OKGQA and variant OKGQA-P, to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios.
+Unlike existing benchmarks that focus primarily on closed ended tasks, OKGQA presents diverse open-ended question types that mirror the unpredictable nature of practical applications.
+We conduct a series of experiments and analyze the effectiveness of various retrieval methods and LLMs of different magnitudes, providing insights for further research.
 Our results underscore the significance of integrating KGs with LLMs to help reduce hallucination of LLMs, even in circumstances where the KGs are contaminated.
+Limitations
+Our proposed benchmark primarily use DBpedia as the knowledge source, which may not generalize well to testing scenarios requiring highly specialized or domain-specific knowledge.
+Testing domain-specific open-ended QA may require constructing sub-graphs from domain-specific KGs.
+In addition, the study assumes a static KG for reasoning and retrieval.
+In dynamic environments where knowledge is continuously updated, maintaining and integrating real-time changes remains a challenge and may requires further design.
 Appendix A Appendix
-A.1 Prompt List
-In this section, we present all the prompts required for the main experiments.
-To enhance clarity, we provide only one example in the prompt labeled as Example 1; the other few-shot examples utilized are labeled as Other In-Context Few-shots within the prompt.
-A.1.1 Knowledge-Augmented Language Model Prompting
-System Instruction: “You are a helpful assistant designed to answer the users’ open-ended questions.
-Your task is to provide accurate, concise, and useful information to foster understanding and solve problems.
-Whether the questions relate to complex scientific concepts, historical events, practical advice, or everyday life, your goal is to assist by offering thoughtful and informative responses."
-Example 1:
-Question: How do you think blockchain technology can be applied to secure online voting systems? Please provide specific suggestions?
-Below are the facts that might be relevant to answer the question: (’Bitcoin’, ’wikiPageWikiLink’, ’Bill_Gates’), (’Bitcoin’, ’wikiPageWikiLink’, ’Non-fungible_token’), (’Decentralized_application’, ’wikiPageWikiLink’, ’Cardano_(blockchain_platform)’), (’Non-fungible_token’, ’wikiPageWikiLink’, ’Decentralized_application’), (’Hyperledger’, ’wikiPageWikiLink’, ’Blockchain’), (’Blockchain’, ’implementations’, ’Bitcoin’), (’Blockchain’, ’currencyName’, ’Litecoin’), (’Thing’, ’type’, ’Privacy_and_blockchain’), (’Thing’, ’type’, ’Saraju_Mohanty’), (’Thing’, ’type’, ’Blockchain’), (’Thing’, ’type’, ’Satoshi_Nakamoto’), (’Thing’, ’type’, ’Pseudonym’), (’Thing’, ’type’, ’Metaverse’), (’Thing’, ’type’, ’PricewaterhouseCoopers’)
-Answer: Blockchain technology can significantly enhance the security of online voting systems through its decentralized, transparent, and immutable nature.
-By distributing data across multiple nodes, blockchain eliminates single points of failure, reducing the risk of hacking or vote tampering.
-The system allows for verifiable, yet anonymous, vote recording through cryptographic methods, ensuring voter privacy while preventing fraud such as double voting.
-Additionally, once recorded, votes cannot be altered, safeguarding election integrity.
-Smart contracts can further automate the voting process, enabling instant and accurate tallying, while the system’s resilience to cyberattacks like DDoS ensures uninterrupted functionality.
-Other In-Context Few-shots
-Question: {Question}
-Below are the facts that might be relevant to answer the question: {Retrieved-knowledge}
-Answer:
-A.1.2 OKGQA Query Generation Prompting
-System Instruction: Generate open-ended questions about different types: character description, event description, cause explanation, relationship explanation, trend prediction, outcome prediction, contrast analysis, historical comparison, evaluation and reflection, and application and practice.
-Some templates are provided below: character description: describe a [person]’s significant contributions during their career.
-Example: Please describe Albert Einstein’s contributions to the field of physics.
-event description: Provide a detailed description of the background and course of an [event].
-Example: Please provide a detailed description of the background and course of the French Revolution.
-cause explanation: Why did [person] take [action] at [time]?
-Example: Why did Nixon choose to resign from the presidency in 1974?
-relationship explanation: Explain the relationship between [entity A] and [entity B] and its significance.
-Example: Explain the relationship between Alexander the Great and Aristotle and its significance.
-trend prediction: Based on the historical behavior of [entity], what do you think it might do in the future?
-Example: Based on Tesla’s historical behavior, in which fields do you think it might innovate in the future?
-outcome prediction: Based on the current situation, how do you predict [event] will develop?
-Example: Based on the current international situation, how do you predict climate change policies will develop?
-contrast analysis: Compare and contrast the similarities and differences between [entity A] and [entity B] in [aspect].
-Example: Compare and contrast the leadership styles of Steve Jobs and Bill Gates.
-historical comparison: Compare the impact of [historical event A] and [historical event B].
-Example: Compare the impact of World War I and World War II on the global order evaluation and reflection: How do you evaluate the impact of [person/event] on [field]?
-Please explain your viewpoint.
-Example: How do you evaluate Martin Luther King’s impact on the civil rights movement?
-Please explain your viewpoint.
-application and practice: How do you think [theory/method] can be applied to [practical issue]?
-Please provide specific suggestions.
-Example: How do you think machine learning technology can be applied to medical diagnostics?
-Please provide specific suggestions.
-Generate the questions, the type of the questions, the placeholders, the naturalness of your generated questions (choose from high, medium, and unnatural), the difficulty of the generated questions (choose from hard, medium and easy) and dbpedia_entities (link the placeholders to dbpedia entities) in JSON format.
-Other In-Context Few-shots
-Generation:
-A.2 Prompts for Instruction Tuner
-Act as an "Instruction Tuner" for the LLM, you will be given the inputs: (1) the {Current Instruction} used to guide the LLMs’s evaluation, including specific examples with ground truth labels; (2) {Current Errors} that emerged with this instruction was applied to the dataset.
-The current errors are presented in the following format: (1) INPUT: {input text} (2) PREDICTED OUTPUT: {predicted label}, (3) EXPECTED OUTPUT: {ground truth label}.
-Carefully analyze these errors and craft a revised concise instruction for the LLM to fit the expected outputs.
-Include 2-3 examples at the end of your response to demonstrate how the new instruction would be applied.
-A.3 Metrics Prompt for G-eval
-System Instruction: “You are a helpful assistant designed to evaluate the quality of the response to a query. Your task is to rate the response on one metric defined as below:"
-Empowerment Criteria: Evaluate whether the ‘Actual Output’ can help the reader understand the topic and make informed decisions regarding the ‘Input’.
-A response with high empowerment provides accurate information and explanations that enhance the reader’s understanding.
-When evaluating empowerment, consider the relevance of the information provided in the ‘Actual Output’ to the ’Input’ and the ’Retrieval Context’.
-Comprehensiveness Criteria: Evalute the extent to which the ’Actual Output’ covers all aspects and details of the question ’Input’.
-A comprehensive answer should thoroughly address every part of the question, leaving no important points unaddressed.
-When evaluating comprehensiveness, consider the relevance of the information provided in the ’Actual Output’ to the ’Input’ and the ’Retrieval Context’.
-Correctness Criteria: Measure how clearly and specifically the ’Actual output’ responds to the question ’input’.
-A highly direct response stays focused on the question, providing clear and unambiguous information.
-When evaluating correctness, consider the relevance of the information provided in the ’Actual Output’ to the ’Input’ and the ’Retrieval Context’.
-Context Relevance Criteria: Evaluate the extent to which the ’Actual output’ incorporates relevant information from the ’Retrieval Context’.
-This includes assessing whether the output adheres to the thematic, factual, and situational specifics presented in the ’Retrieval Context’.
-Relevant responses not only address the direct query but also align closely with the contextual elements provided, ensuring a seamless and coherent transition between the ’Retrieval Context’ and the ’Actual Output’.
-The most contextually relevant responses demonstrate an understanding and appropriate reflection of the given circumstances, historical facts, or conceptual background, thereby contributing to the overall accuracy and utility of the information provided.
-A.4 Query Construction
+A.1 Intuition of Open-ended Questions?
+In previous close-ended KGQA research Mihaylov et al.
+(2018); Yih et al.
+(2016); Talmor and Berant (2018), “hallucination” has not been considered at all.
+The primary reason is that close-ended KGQA models typically generate concise entity names as answers, which are directly compared against ground truth labels using metrics such as precision or accuracy.
+Any incorrect entity name is immediately reflected in the evaluation results, leaving little room for hallucination to manifest as a separate problem.
+While effective for simple tasks, we believe this closed-ended setting is too naive and fails to generalize to real-world, complex QA situations, where models must provide more detailed, nuanced responses Kantharaj et al.
+(2022).
+In contrast, our work focuses on open-ended KGQA, where LLMs are prompted to generate more elaborate answers that include reasoning paths and supporting facts derived from the background KG.
+This broader output space not only provides a larger field to assess hallucination but also increases the likelihood of factual inaccuracies, as LLMs tend to produce more errors when generating longer and more complex responses.
+While hallucination is minimal in close-ended KGQA due to the short and factual nature of entity names, open-ended KGQA highlights the susceptibility of LLMs to hallucinate when generating rich, explanatory answers.
+By introducing this open-ended setting, OKGQA, we (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on method design.
+We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs to reduce hallucination.
+A.2 Query Construction
 In this section, we discuss the details of the query construction of OKGQA.
@@
 Furthermore, we follow a human-in-the-loop process to optimize the instruction used for generation, as shown in Figure 7.
-Intuitively, starting with an initial instruction, we generate a large number of query candidates, then use an LLM as automatic evaluator to evaluate the quality of the query, denoted as a set of scores from different metrics $s_{\text{auto}}$ ranging from 1 to 10, with higher scores indicating better performance.
-Then, these queries are manually evaluated and assign a set of human-label score $s_{\text{human}}$ correspond to $s_{\text{auto}}$, normalized as the same range as $s_{\text{auto}}$.We then optimize the input instructions by iteratively generate instructions to minimize the gap between $s_{\text{human}}$ and $s_{\text{auto}}$.
-This process is quite mimic the way of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and inherits the benefit that labeling the reward or penalty of the LLMs’ output is much easier than labeling the output directly.
-Figure 6: Process of human-in-the-loop of query construction.
+Intuitively, starting with an initial instruction, we generate a large number of query candidates, then use an LLM as an automatic evaluator to evaluate the quality of the query, denoted as a set of scores from different metrics $s_{\text{auto}}$ ranging from 1 to 10, with higher scores indicating better performance.
+Then, these queries are manually evaluated and assigned a set of human-label scores $s_{\text{human}}$ corresponding to $s_{\text{auto}}$, normalized to the same range as $s_{\text{auto}}$.
+We then optimize the input instructions by iteratively generating instructions to minimize the gap between $s_{\text{human}}$ and $s_{\text{auto}}$.
+This process quite mimics the way of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and inherits the benefit that labeling the reward of the LLMs’ output is much easier than labeling the output directly.
+Figure 7: Human-in-the-loop of query construction.
 Specifically, we consider five angles to measure the quality of the generated queries: (1) Naturalness: assessing how fluid and human-like the query sounds; (2) Relevance: measuring whether the query pertains directly to the entity and the context provided; (3) Specificity: determining the level of detail and granularity included in the query, ensuring it is not too broad or vague; (4) Novelty: evaluating the uniqueness of the query, ensuring it is not just a repetitive or common question; (5) Actionability: gauging whether the query prompts clear, definite answers or actions that are feasible within the given context.
 Each of these angles contributes to a holistic evaluation of the query’s effectiveness and relevance in real-world applications.
-A.5 Personalized PageRank (PPR)
-In this section, we discuss the details of the PPR algorithm used in Section 1 to prune the graph from DBPedia and concentrate on nodes most pertinent to the central nodes of interest.
+We have three evaluators participating in the manual assessment of query quality.
+All of the evaluators are computer science majors with fluent English skills.
+As the evaluation centers on various linguistic metrics such as naturalness, relevance, specificity, novelty, and actionability, we only require the evaluators to possess a fundamental understanding of English without restricting their majors.
+We calculate the Pearson correlation coefficients between human and LLM scores as shown in Table 3.
+It shows that as the rounds progress, agreement between humans and LLMs increases, suggesting that iterative feedback improves alignment between human annotation and LLM responses.
+Table 3: Pearson correlation coefficients between human and LLM scores across rounds.
+Metric Round 1 Round 2 Round 3 Round 4
+Naturalness 0.60 0.65 0.69 0.74
+Relevance 0.55 0.59 0.64 0.70
+Specificity 0.46 0.54 0.60 0.65
+Novelty 0.49 0.57 0.63 0.67
+Actionability 0.33 0.41 0.48 0.53
+In addition, we also consider verifying the inter-rater reliability across three evaluators as shown in Table 4.
+We report the Cohen’s Kappa coefficient for each pair of evaluators as follows.
+Using the Landis and Koch (1977) interpretation guidelines, the Cohen’s Kappa coefficients for Naturalness and Relevance (ranging from 0.79 to 0.85) fall within the “Substantial” to “Almost Perfect” categories, indicating strong inter-rater reliability for these metrics.
+This reflects a shared understanding of the evaluation criteria, resulting in consistent ratings among evaluators.
+For Specificity, Novelty, and Actionability, the coefficients range from 0.58 to 0.68, placing them primarily in the “Moderate” to “Substantial” categories.
+These results suggest moderate reliability for these metrics, likely due to subjective interpretation and less clearly defined evaluation guidelines.
+Novelty, with lower coefficients around 0.61 to 0.63, highlights variability in ratings, suggesting that evaluators may have differing perspectives on what qualifies as novel (but the inter-rater reliability is still considered “Substantial”).
+Meanwhile, Actionability performs slightly better, nearing the “Substantial” range, indicating moderately consistent criteria.
+Table 4: Cohen’s Kappa coefficient for various metrics.
+Metric Evaluator 1 & 2 Evaluator 1 & 3 Evaluator 2 & 3
+Naturalness 0.85 0.83 0.84
+Relevance 0.81 0.79 0.80
+Specificity 0.65 0.63 0.66
+Novelty 0.60 0.58 0.61
+Actionability 0.67 0.65 0.68
+A.3 Personalized PageRank (PPR)
+In this section, we discuss the details of the PPR algorithm used in §2.1 to prune the graph from DBPedia and concentrate on nodes most pertinent to the central nodes of interest.
 The PPR is calculated using the iterative formula:
@@
 […]
-A.6 Prize-Cost-based Path Retrieval
+A.4 Prize-Cost-based Path Retrieval
 The path-retrieval method is designed to construct and evaluate paths in a graph based on predefined prize assignments and cost allocations.
 The objective is to form sequences of nodes and edges, represented as $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$, that maximize the overall score and minimize the costs.
-To efficiently manage the exploration of potential paths, we utilizes a priority queue, a data structure that allows paths to be organized based on their scores, ensuring that the highest-scoring paths are processed first.
+To efficiently manage the exploration of potential paths, we utilize a priority queue, a data structure that allows paths to be organized based on their scores, ensuring that the highest-scoring paths are processed first.
 The method starts by picking a number of starting nodes with high prizes.
@@
 When no more paths can be added or the priority queue is empty, the algorithm sorts the paths by their scores from highest to lowest.
-A.7 KG Similarity Metrics
-To measure how much the perturbed KGs are deviated from the original KG used in Section 3. We re-present the according metrics from (Raman et al., 2020) below:
+A.5 KG Similarity Metrics
+To measure how much the perturbed KGs are deviated from the original KG used in §2. We re-present the according metrics from (Raman et al., 2020) below:
 Aggregated Triple Score (ATS)  ATS measures semantic similarity between two KGs.
@@
 Similarity in Degree Distribution (SD2)  SD2 also measures structural similarity between two KGs, while addressing SC2D’s ineffectiveness when the KGs’ entities have tiny local clustering coefficients (e.g., the item KG used by recommender systems is roughly bipartite).
-In such cases, SC2D is always close to one regardless of perturbation method, thus rendering SC2D useless.
+In such cases, SC2D is always close to one regardless of the perturbation method, thus rendering SC2D useless.
 Let $\mathbf{d}^{r}$ denote the $|\mathcal{E}|$-dimensional degree vector for $\mathcal{G}^{r}$, where the $i$th element of $\mathbf{d}^{r}$ is $\text{deg}(e_{i})$.
@@
 SD2 is defined as $f_{\text{SD2}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=1-\frac{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}}{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}+1}\in[0,1]$, with higher value indicating higher similarity.
+Appendix B Related Work
+Due to the stochastic decoding process of Large Language Models (LLMs), i.e., sampling the next token in the sequence, LLMs exhibit probabilistic behaviors: (1) potentially yielding varied outputs of the same input across different instances (Agrawal et al., 2023); (2) cannot accurately interpret phrases or terms when the context is vague and resides in a knowledge gap of the model.
+This will lead to outputs that may sound plausible but are often irrelevant or incorrect.
+This will lead to outputs that may sound plausible but are often irrelevant or incorrect.
+This “hallucinations” undermines the reliability of LLMs (Huang et al., 2023).
+One emerging research trend is enhancing LLMs through integrating external knowledge graphs (Agrawal et al., 2023).
+KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023; Sui et al., 2022).
+Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance.
+This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent.
+Researchers employ diverse strategies to augment the LLMs by integrating external KGs (Sui et al., 2024; He et al., 2024b).
+For example, KAPING (Baek et al., 2023b) matches entities in questions to retrieve related triples from knowledge graphs for zero-shot question answering.
+Wu et al.
+(2023) finds that converting these triples into textualized statements can further enhance LLM performance.
+StructGPT (Jiang et al., 2023b) propose to convert user query into structured formats (e.g., SPARQL) for information extraction from KGs.
+Following the succuess of internal reasoning-enhancement methods like Chain-of-thoughts (CoT) (Wei et al., 2022), Reflexion (Shinn et al., 2024), and Tree-of-thoughts (ToT), He et al.
+(2022) propose “rethinking with retrieval” to use decomposed reasoning steps from CoT prompting to retrieve external knowledge, leading to more accurate and faithful explanations.
+IR-CoT (Trivedi et al., 2022) interleaves the generation of CoT with knowledge retrieval from corresponding KGs, iteratively guiding both retrieval and reasoning for multi-step questions.
+MindMap (Wen et al., 2023) introduce a plug-and-play approach to evoke graph-of-thoughts reasoning in LLMs.
+Similarly, RoG (Luo et al., 2023b) use KGs to create faithful reasoning paths based on various relations, enabling interpretable reasoning in LLMs.
+However, current benchmarks for testing the capabilities of these LLM+KG models are predominantly closed-ended, restricting responses to a limited set of entities/relations or a set of logical forms derived from specific facts of KG.
+Hence, they can only test a very limited subset of the LLM’s tendency to hallucinate, leaving a gap in the assessment of complex, real-world scenarios.
+Particularly, standard metrics such as FActScore (Min et al., 2023) and SAFE (Wei et al., 2024) for evaluating the hallucination rate of LLMs require open-ended settings, i.e., questions are phrased as a statement which requires a longer answer.
+Compared with previous works, our proposed OKGQA is tailored for evaluating LLMs enhanced with KGs under open-ended, real-world question-answering scenarios.
+The benchmark extends the assessment of closed-ended question answering to an open-ended setting, which can further support the assessment of hallucination of LLMs.
+Appendix C Prompt List
+In this section, we present all the prompts required for the main experiments.
+To enhance clarity, we provide only one example in the prompt labeled as Example 1; the other few-shot examples utilized are labeled as Other In-Context Few-shots within the prompt.
+C.1 Knowledge-Augmented Generation
+System Instruction: “You are a helpful assistant designed to answer the users’ open-ended questions.
+Your task is to provide accurate, concise, and useful information to foster understanding and solve problems.
+Whether the questions relate to complex scientific concepts, historical events, practical advice, or everyday life, your goal is to assist by offering thoughtful and informative responses.”
+In-Context Few-shots
+Question: {Question}
+Below are the facts that might be relevant to answer the question: {Retrieved-knowledge}
+Answer:
+C.2 OKGQA Query Generation Prompting
+System Instruction: “Generate open-ended questions about different types: character description, event description, cause explanation, relationship explanation, trend prediction, outcome prediction, contrast analysis, historical comparison, evaluation and reflection, and application and practice.
+Please provide specific suggestions.
+Generate the questions, the type of the questions, the placeholders, the naturalness of your generated questions (choose from high, medium, and unnatural), the difficulty of the generated questions (choose from hard, medium and easy) and DBPedia_entities (link the placeholders to DBPedia entities) in JSON format.”
+Example 1: as shown in Figure 8.
+Figure 8: Example 1 Demonstration.
+Other In-Context Few-shots
+Generation:
+C.3 Prompts for Instruction Tuner
+Act as an “Instruction Tuner” for the LLM, you will be given the inputs: (1) the {Current Instruction} used to guide the LLMs’s evaluation, including specific examples with ground truth labels; (2) {Current Errors} that emerged with this instruction are applied to the dataset.
+The current errors are presented in the following format: (1) INPUT: {input text} (2) PREDICTED OUTPUT: {predicted label}, (3) EXPECTED OUTPUT: {ground truth label}.
+Carefully analyze these errors and craft a revised concise instruction for the LLM to fit the expected outputs.
+Include 2-3 examples at the end of your response to demonstrate how the new instruction would be applied.
+C.4 Metrics Prompt for G-eval
+System Instruction: “You are a helpful assistant designed to evaluate the quality of the response to a query. Your task is to rate the response on one metric defined as below:”
+Empowerment Criteria: Evaluate whether the “Actual Output” can help the reader understand the topic and make informed decisions regarding the “Input”.
+A response with high empowerment provides accurate information and explanations that enhance the reader’s understanding.
+When evaluating empowerment, consider the relevance of the information provided in the “Actual Output” to the “Input” and the “Retrieval Context”.
+Comprehensiveness Criteria: Evaluate the extent to which the “Actual Output” covers all aspects and details of the question “Input”.
+A comprehensive answer should thoroughly address every part of the question, leaving no important points unaddressed.
+When evaluating comprehensiveness, consider the relevance of the information provided in the “Actual Output” to the “Input” and the “Retrieval Context”.
+Correctness Criteria: Measure how clearly and specifically the “Actual output” responds to the question “input”.
+A highly direct response stays focused on the question, providing clear and unambiguous information.
+When evaluating correctness, consider the relevance of the information provided in the “Actual Output” to the “Input” and the “Retrieval Context”.
+Context Relevance Criteria: Evaluate the extent to which the “Actual output” incorporates relevant information from the “Retrieval Context”.
+This includes assessing whether the output adheres to the thematic, factual, and situational specifics presented in the “Retrieval Context”.
+Relevant responses not only address the direct query but also align closely with the contextual elements provided, ensuring a seamless and coherent transition between the “Retrieval Context” and the “Actual Output”.
+The most contextually relevant responses demonstrate an understanding and appropriate reflection of the given circumstances, historical facts, or conceptual background, thereby contributing to the overall accuracy and utility of the information provided.
+Response: [Respond with metric and the corresponding score.]

### v3

@@
 We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs to reduce hallucination.
-The code and data are released at link.
 Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study Over Open-ended Question Answering
@@
 1 Introduction
-Contemporary LLMs are prone to producing hallucinations due to gaps in their knowledge (Gekhman et al., 2024; Lee et al., 2023), primarily when the training data contain misinformation, biases, or inaccuracies.
-These flaws can arise when we simply maximize the log-likelihood, leading to responses that may seem plausible, but are often irrelevant or incorrect (Weng, 2024).
-This issue is especially concerning in scenarios where accuracy and reliability are essential, such as healthcare (He et al., 2023) and science (Taylor et al., 2022).
+Contemporary LLMs are prone to producing hallucinations due to gaps in their knowledge Gekhman et al.
+(2024); Lee et al.
+(2023).
+These inaccuracies commonly stem from misinformation, biases, or errors in the training data, and lead to responses that seem plausible but may be irrelevant or incorrect Weng (2024).
+This issue is particularly concerning in high-stakes contexts such as healthcare He et al.
+(2023) and science Taylor et al.
+(2022) [^1] Code and data are released at https://anonymous.4open.science/r/OKGQA-CBB0.
 Figure 1: Comparison between Open-ended Question and Close-ended Question over Knowledge Graph.
-To address this critical limitation, researchers have turned to leveraging external knowledge graphs (KGs) as a complementary (Pan et al., 2024; Luo et al., 2023a; Hu et al., 2023; Sun et al., 2023; van Breugel et al., 2021).
-KGs provide several key advantages: they contain structured and explicit factual knowledge, including domain-specific information, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023).
-Furthermore, KGs offer traceability, allowing each piece of information to be traced back to its source.
-This traceability enables both verification of generated context and transparency in reasoning paths, making KGs a promising source for improving the reliability of LLM reasoning.
-Find more details of related works in §B.
-However, current benchmarks for testing the capabilities of these LLM+KG models are predominantly closed-ended, restricting responses to a limited set of entities or relations (Talmor et al., 2019; Mihaylov et al., 2018; Jin et al., 2020) or a set of logical forms (Yih et al., 2016; Talmor and Berant, 2018; Puerto et al., 2023) derived from specific facts of KG (as shown in Figure 1).
-These benchmarks, while useful for evaluating basic retrieval and reasoning, are insufficient for testing the LLMs’ ability for faithful generation (i.e., minimize hallucination in LLM responses), leaving a gap in the assessment of complex, real-world scenarios.
-Why Open-ended Question?
-A key limitation of closed-ended benchmarks is: they cannot assess the phenomenon of “more talk, more mistakes”, where longer, more complex responses tend to increase hallucination Qiu et al.
+To address this limitation, researchers have turned to leveraging external knowledge graphs (KGs) as a complementary Pan et al.
+(2024); Luo et al.
+(2023a); Hu et al.
+(2023); Sui et al.
 (2024).
-Unlike closed-ended benchmarks that require short, specific answers, open-ended questions demand more detailed and nuanced explanations.
-This format also enables direct evaluation metrics like FActScore (Min et al., 2023) and SAFE (Wei et al., 2024), which measure factual consistency of LLM responses by decomposing them into atomic statements for comparison with external information.
-However, for short, rigid responses, this decomposition is not feasible, rendering these metrics inapplicable.
-Moreover, open-ended evaluations align more closely with human cognition Lu et al.
-(2024); Amirizaniani et al.
-(2024), where context-rich and interpretative answers are expected compared to simplistic and mechanical responses.
-Find more details of intuition in Appendix §A.1.
-Motivation: Recognizing the research gap, our study aims to contribute to the existing knowledge from two perspectives: 1) explore whether KGs can make LLMs more trustworthy in the open-ended setting, and 2) conduct a comparative analysis to shed light on methods and directions for leveraging KGs to reduce LLMs’ hallucination.
-To achieve this, we develop a new benchmark, Open-ended Knowledge-Graphs Question Answering (OKGQA), specifically designed to assess LLMs enhanced with KGs under open-ended question answering scenarios.
-OKGQA is designed to closely reflect the complexities of practical applications using questions of different types and incorporates specific metrics to measure both the reduction in hallucinations and enhancements in reasoning capabilities.
-Specifically, all queries in OKGQA are synthesized by LLMs with pre-defined templates mentioned in Table 1 and corresponding Wikipedia context.
-To consider the scenarios in which KGs may have varying levels of mistakes (i.e., attributes may be mislabeled, relations may not exist, etc.), we propose another experiment setup OKGQA-P (§2.2) to assess model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated.
-We assess the hallucination ratio and the quality of the response (detailed metrics can be found in §4.1) of the tested models.
-Based on our experiments, we find that (1) retrieving KG information can mitigate factual errors in LLMs’ responses, especially for queries requiring additional reasoning; (2) directly performing reasoning in the LLM itself (e.g., internal reasoning strategies like Chain-of-thoughts (Kim et al., 2023) and self-consistency (Wang et al., 2022)) may cause bias and hallucination; (3) Subgraphs generally achieve the best performance across different query types, especially for simpler types of queries; (4) Integrating KGs can effectively reduce hallucinations in LLMs, even when KGs are contaminated to a certain degree.
-2 OKGQA: A Open-ended Knowledge Graph Question-Answering Benchmark
-Overview.
+KGs offer structured and explicit factual information—often domain-specific—and allow each piece of data to be traced back to its source Zheng et al.
+(2023); Agrawal et al.
+(2023).
+This traceability not only enables verification of the model’s reasoning but also brings transparency to the decision-making process, making KGs a promising method for enhancing the reliability of LLM reasoning.
+Find more details of related works in §C.
+However, current benchmarks for testing these LLM+KG models are predominantly closed-ended Jin et al.
+(2020); Puerto et al.
+(2023), restricting the model’s output to a limited set of entities, relations Talmor et al.
+(2019); Mihaylov et al.
+(2018) or logical forms Yih et al.
+(2016); Talmor and Berant (2018).
+While these benchmarks are useful to measure retrieval and basic reasoning, they do not adequately capture whether a model is hallucinating.
+In closed-ended settings, errors can stem from incorrect retrieval or from fabricating (hallucinating) answers, yet conventional metrics (e.g., accuracy or precision) cannot distinguish between these two issues.
+This becomes problematic for more complex, real-world applications that demand nuanced answers Kantharaj et al.
+(2022).
+In contrast, our work focuses on open-ended KGQA, where LLMs are prompted to generate more elaborate answers, include reasoning paths and supporting facts derived from the KG (as shown in Figure 1).
+This broader output space offers two key advantages: First, it enables direct assessments of hallucination with metrics like FActScore Min et al.
+(2023) or SAFE Wei et al.
+(2024)), which decompose longer responses into atomic statements for factual consistency checks with external knowledge sources like Wikipedia.
+Second, it increases the likelihood of exposing factual errors which helps to assess the phenomenon indicated in Qiu et al.
+(2024): where longer, more complex responses provide more opportunities for errors to occur.
+By adopting this open-ended approach, we aim to (1) explore whether KGs can make LLMs more trustworthy in the open-ended setting, and (2) conduct a comparative analysis to shed light on methods design and direction for leveraging KGs to reduce LLMs’ hallucination.
+To achieve this, we introduce a new benchmark, Open-ended Knowledge-Graphs Question Answering (OKGQA), specifically designed to assess LLMs enhanced with KGs in an open-ended QA setting.
+OKGQA is designed to closely reflect the complexities of practical applications with diverse questions types as mentioned in Table 1, ensuring that all the queries cannot be answered simply by retrieving isolated KG facts.
+To consider the scenarios for potentially contaminated or imperfect KGs (i.e., attributes may be mislabeled, relations may not exist, etc.), we also propose a variant OKGQA-P (§2.2) to assess model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated.
+In both settings, we measure the degree of hallucination and the overall quality of the responses (see §4.1 for details).
+Based on our experiments, we find that (1) integrating KG information generally mitigates factual errors, especially for queries requiring deep reasoning; (2) directly performing reasoning in the LLM itself (e.g., internal reasoning strategies like Chain-of-thought Kim et al.
+(2023) and Self-Consistency Wang et al.
+(2022)) may cause bias and hallucination; (3) subgraph-based methods often achieve the best performance for simpler query types; and (4) incorporating KGs effectively reduces hallucinations in LLMs even when the KG is partially contaminated.
+2 OKGQA: An Open-ended Knowledge Graph Question-Answering Benchmark
 OKGQA is a comprehensive benchmark designed to assess how effectively LLMs enhanced with KGs perform in open-ended, real-world-like question answering scenarios.
 Unlike existing benchmarks that focus primarily on closed-ended tasks, OKGQA presents diverse open-ended question types that mirror the variable nature of practical applications.
-The benchmark evaluates how effectively models leverage KGs to generate both reliable and comprehensive answers, focusing on reducing hallucinations while maintaining reasoning capabilities of LLMs.
-To assess robustness under imperfect conditions, we also introduce OKGQA-P, a variant where KGs are deliberately perturbed and contaminated.
+As illustrated in Figure 1, given a complex query and its corresponding subgraph in a KG, the system must be capable of understanding the relationships within the data and performing human-like reasoning over the KG content to compose a paragraph-long answer.
+In the following section, we first describe our dataset construction, including query generation via LLM templates and KG subgraph extraction with PPR pruning.
+We then introduce OKGQA-P, a benchmark variant designed to evaluate model robustness under KG perturbations, detailing our perturbation methods and the metrics used to assess semantic and structural deviations.
+Due to the page limitation, we also include some extension of our benchmark in Appendix §B, including multilingual setup, and more analysis.
 2.1 Dataset Construction
 Statistics (on average)
+[…]
 Tokens in query 23.97
-Total number of queries 850
+Total number of queries 850 $\rightarrow$ 2,050
 Number of unique DBPedia entities 816
 Before Pruning $\rightarrow$ After PPR Pruning
+[…]
 Tokens in subgraph 348,715 $\rightarrow$ 2,452
@@
 Query Type Simple Complex
+[…]
 Descriptive 78 11
@@
 Queries.
-We utilize a template-based methodology to generate a diverse range of queries using LLMs, including categories such as descriptive, explanatory, predictive, comparative, and critical queries.
-Details regarding specific templates and example queries can be found in Table 1, while the corresponding prompts are provided in the Appendix C.
+We utilize a template-based approach to generate a diverse range of queries using LLMs, including categories such as descriptive, explanatory, predictive, comparative, and critical queries.
+Details regarding specific templates and example queries can be found in Table 1, while the corresponding prompts are provided in the Appendix D.
 To ensure that the generated queries represent real-world scenarios and complexities, we employ an iterative optimization approach that utilizes both automated and human evaluation to refine the query generation process (the details are given in Appendix A.1).
@@
 KG Sub-graphs.
-To reduce the size of KGs, following previous works (Yih et al., 2016; Talmor and Berant, 2018), we sample subgraphs from DBpedia Knowledge Graph (Noted that all the queries in OKGQA can be reasoned based on DBPedia) by extracting all triples contained within the k-hop neighbors from the question entities in the query.
-We set k = 2 to balance graph size and computational feasibility, as increasing k leads to exponential growth in the number of edges and nodes, which can introduce excessive noise and make it more challenging to retrieve essential information from KGs [^1] This choice is also informed by common practices in other benchmarks, such as WebQSP Yih et al.
+To reduce the size of KGs while retaining relevant information, we follow previous work Yih et al.
+(2016); Talmor and Berant (2018) by sampling subgraphs from DBpedia (Noted that all queries in OKGQA can be answered using DBpedia).
+We extract all triples contained within the $K$-hop neighbors from the entities mentioned in the query.
+We set $K=2$ to balance graph size and computational feasibility.
+As increasing beyond $2$-hop subgraphs generally leads to exponential growth in edges and nodes Jin et al.
+(2020), which increase excessive noise and complicating information retrieval [^2] This choice is also informed by common practices in other benchmarks, such as WebQSP Yih et al.
 (2016) and CWQ Talmor and Berant (2018), where 2-hop subgraphs are widely used for similar KGQA tasks..
-To further reduce the size of the 2-hop subgraphs, we leverage personalized page-rank (PPR) (Bahmani et al., 2010) to prune the nodes/edges that are not relevant to the query (the details of the PPR algorithm are discussed in Appendix A.3).
+To further reduce the size of the 2-hop subgraphs, we leverage Personalized Page-Rank (PPR) Bahmani et al.
+(2010) to prune the nodes/edges that are not relevant to the query (the details of the PPR algorithm are discussed in Appendix A.2).
 We compare the statistics of subgraphs before and after PPR pruning in Figure 2(a).
 2.2 OKGQA-P: Benchmark with Noise & Perturbations in KGs
-To further mimic the real situations where KGs may not be of high quality (i.e., attributes of nodes/edges may be mislabeled, relations may not exist, etc.), we propose another experiment setting OKGQA-P to assess the model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated.
-We introduce various perturbation scenarios including mislabeled attributes, incorrect relations, and missing connections to test the model’s reasoning capabilities with flawed or incomplete KG data.
-To quantify the degree of perturbation, we also evaluate both the semantic and structural similarity between the original and the modified KG as defined below.
+KGs are often annotated by humans and can contain errors such as mislabeled attributes or missing relations.
+To mimic the real situations where KGs’ quality may not be fully reliable, we propose OKGQA-P to assess the model performance under deliberately perturbed and contaminated KGs.
+We introduce various perturbation scenarios including mislabeled attributes, incorrect relations, and missing connections to test how well models can handle flawed or incomplete KG data.
+To quantify the degree of perturbation, we measure the semantic and structural similarity between the original and the modified KG as defined below.
 Notation.
@@
 Let $g(\mathcal{G})$ be the downstream performance when evaluating $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}$.
-High-level Procedure.
-First, we test $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}$ to get the original performance $g(\mathcal{G})$.
-Second, we perturb $\mathcal{G}$ to obtain $\mathcal{G}^{\prime}$.
-Third, we evaluate $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}^{\prime}$ to get the perturbed performance $g(\mathcal{G}^{\prime})$.
-Finally, we measure $g(\mathcal{G})-g(\mathcal{G}^{\prime})$ and $f(\mathcal{G},\mathcal{G}^{\prime})$ to assess how robust $\mathcal{F}_{\theta}$ is, i.e., to assess the model performance under conditions where KGs’ semantics and structure are deliberately perturbed.
-To quantify how much the perturbed KG has deviated from the original KG, i.e., $f(\mathcal{G},\mathcal{G}^{\prime})$, we leverage metrics from Raman et al.
-(2020) for capturing semantics (ATS) and structural (SC2D, SD2) similarity between KGs.
-Intuitively, ATS leverages a pre-trained LM for link prediction to measure the probability of each edge from $\mathcal{G}^{\prime}$ existing in $\mathcal{G}$, while SC2D and SD2 measure the structural similarity between two KGs based on local clustering coefficient and degree distribution.
-For each of the three metrics, higher value indicates higher similarity.
-The detailed description can be found in Appendix A.5, with visualization in Figure 5.
-For the perturbation methods, we consider four perturbation heuristics based on (Raman et al., 2020) as follows: Relation Swapping (RS) randomly chooses two edges from $\mathcal{T}$ and swaps their relations.
-Relation Replacement (RR) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, and replaces the $e_{1}$ with another relation $e_{2}=\text{argmin}_{e\in\mathcal{E}}S_{\mathcal{G}}(v_{1},e,v_{2})$, where $S_{\mathcal{G}}(\cdot)$ is a KG score function adapted from ATS.
-The replacement is performed such that the resulting $(v_{1},e_{2},v_{2})$ has the closest score to the original triplet $(v_{1},e_{1},v_{2})$ to generate “harder negatives” - triplets that are semantically similar but incorrect, thereby challenging the model to different subtle distinctions.
-Noted that $S_{G}$ differs from the original ATS, as it operates on individual triplets rather than comparing two entire KGs.
-Edge Rewiring (ER) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, then replaces $v_{2}$ with another entity $v_{3}\in\mathcal{E}\backslash\mathcal{N}_{1}(v_{1})$, where $\mathcal{N}_{1}(v_{1})$ represents the 1-hop neighborhood of $v_{1}$.
-Edge Deletion (ED) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$ and deletes it.
-We control the perturbation level based on the percentage of KG edges perturbed.
-See results in Figures 5 and 6.
-Table 1: Query types and examples in OKGQA.
-Brown is used to highlight the placeholders (e.g., [person], [event]) in description, while Teal highlights the specific entities to replace the placeholders.
 Type Sub-Type Description / Template Example
@@
 and course of the French Revolution.
+[…]
 Explanatory Cause Explanation
@@
 and Aristotle and its significance.
+[…]
 Predictive Trend Prediction
@@
 predict climate change policies will develop?
+[…]
 Comparative Contrast Analysis
@@
 on the global order.
+[…]
 Critical Evaluation and Reflection
@@
 applied to medical diagnostics?
+Table 1: Query types and examples in OKGQA.
+Brown is used to highlight the placeholders (e.g., [person], [event]) in description, while Teal highlights the specific entities to replace the placeholders.
+High-level Procedure.
+First, we test $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}$ to get the original performance $g(\mathcal{G})$.
+Second, we perturb $\mathcal{G}$ to obtain $\mathcal{G}^{\prime}$.
+Third, we evaluate $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}^{\prime}$ to get the perturbed performance $g(\mathcal{G}^{\prime})$.
+Finally, we measure $g(\mathcal{G})-g(\mathcal{G}^{\prime})$ and $f(\mathcal{G},\mathcal{G}^{\prime})$ to assess how robust $\mathcal{F}_{\theta}$ is, i.e., to assess the model performance under conditions where KGs’ semantics and structure are deliberately perturbed.
+To quantify how much the perturbed KG has deviated from the original KG, i.e., $f(\mathcal{G},\mathcal{G}^{\prime})$, we leverage metrics from Raman et al.
+(2020) for capturing semantics (ATS) and structural (SC2D, SD2) similarity between KGs.
+Intuitively, ATS leverages a pre-trained LM for link prediction to measure the probability of each edge from $\mathcal{G}^{\prime}$ existing in $\mathcal{G}$, while SC2D and SD2 measure the structural similarity between two KGs based on local clustering coefficient and degree distribution.
+For each of the three metrics, higher value indicates higher similarity.
+The detailed description can be found in Appendix A.5, with visualization in Figure 5.
+For the perturbation methods, we consider four edge-based perturbation heuristics based on Raman et al. (2020) as follows:
+  • Relation Swapping (RS) randomly chooses two edges from $\mathcal{T}$ and swaps their relations.
+  • Relation Replacement (RR) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, and replaces the $e_{1}$ with another relation $e_{2}=\mathrm{argmin}_{e\in\mathcal{E}}S_{\mathcal{G}}(v_{1},e,v_{2})$, where $S_{\mathcal{G}}(\cdot)$ is a KG score function adapted from ATS.
+This yield “harder negatives” - triplets that are semantically similar but incorrect.
+  • Edge Rewiring (ER) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, then replaces $v_{2}$ with another entity $v_{3}\in\mathcal{E}\backslash\mathcal{N}_{1}(v_{1})$, where $\mathcal{N}_{1}(v_{1})$ represents the 1-hop neighborhood of $v_{1}$.
+  • Edge Deletion (ED) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$ and deletes it.
+We control perturbation level by adjusting the percentage of edges in $\mathcal{G}$ that are perturbed. Refer to Figures 5 and 6 for empirical results.
 3 Exploring KG-augmented framework for Reducing Hallucination
-To explore whether KG-augmented approaches can mitigate LLMs’ hallucination, we investigate a range of methods and discover that they can be encapsulated within a unified framework as shown in Figure 3.
-Our framework follows the paradigm of retrieval augmented generation (RAG) (Edge et al., 2024; Baek et al., 2023a), which retrieves essential information from the KGs, and then uses the retrieved knowledge to enhance the LLM’s generation.
-It has two components, i.e., Graph-guided retrieval (G-retrieval) and Graph-guided generation (G-generator), with a variety of algorithmic design choices.
-We analyze our proposed strategies within each component in §4, aiming to shed light on the best practices for leveraging KGs for reducing hallucinations in LLMs.
+To explore whether KG-augmented approaches can mitigate LLMs’ hallucination, we propose a unified framework as shown in Figure 3.
+Our framework follows the paradigm of retrieval augmented generation (RAG) Edge et al.
+(2024); Baek et al.
+(2023a), which retrieves essential information from the KGs, and then uses the retrieved knowledge to enhance the LLM’s generation (§3.1).
+It consists of two components, i.e., Graph-guided retrieval3.2) and Graph-guided generator3.3), with a variety of algorithmic design choices.
+We analyze the strategies within each component in §4, aiming to shed light on the best practices for leveraging KGs for reducing hallucinations in LLMs.
+3.1 Formalization
+We formalize the KG-augmented framework for reducing hallucination as follows.
+Given a user query $q$, a pretrained language model generates a paragraph-like answer $a$ by modeling the conditional probability $p(a|q)$.
+To explore whether KGs help reduce hallucinations of LLMs, we introduce retrieved knowledge $\mathcal{Z}$ from the KG and define:
+$p(a|q)=\sum_{\mathcal{Z}\subseteq\mathcal{G}}p_{\phi}(a|q,\mathcal{Z})p_{\theta}(\mathcal{Z}|q,\mathcal{G}),$ (1)
+where $p_{\phi}(a|q,\mathcal{Z})$ is the likelihood of generating the paragraph-like answer $a$ conditioned on $q$ and $\mathcal{Z}$ (parameterized by $\phi$), and $p_{\theta}(\mathcal{Z}|q,\mathcal{G})$ models the retrieval of knowledge subsets (parameterized by $\theta$).
+Because the number of possible subsets $\mathcal{Z}$ can be exponentially large relative to the size of $\mathcal{G}$, we approximate the sum by selecting the most probable knowledge subset: $\mathcal{Z}^{*}=\text{argmax}_{\mathcal{Z}\in\mathcal{G}}p_{\theta}(\mathcal{Z}|q,\mathcal{G})$, yielding:
+$p(a|q)\approx p_{\phi}(a|q,\mathcal{Z}^{*})p_{\theta}(\mathcal{Z}^{*}|q,\mathcal{G})$ (2)
 Figure 3: Overview of KG-augmented framework.
-Here we formalize the problem of context-relevant KG retrieval for open-ended question answering.
-Given a user query $q$, a pretrained language model (PLM) first encodes the input tokens, and models a conditional distribution $p(a|q)$ to generate an output response $a=[a_{1},\ldots,a_{T}]$.
-To explore whether KGs help reduce hallucinations of LLMs, we introduce the retrieved knowledge from the KG, denoted $\mathcal{Z}\in\mathcal{G}$, which can take various forms (e.g., triplets, paths, subgraphs, as shown in Figure 3).
-This retrieved knowledge contains information relevant to the user query, and either benefits the LLMs at its knowledge boundaries (Baek et al., 2023b) or explicitly guides its reasoning process (Luo et al., 2023a; Sun et al., 2023).
-The distribution of the retrieved knowledge $\mathcal{Z}$ is modeled as $p(\mathcal{Z}|q,\mathcal{G})$, leading to a new likelihood for generating responses, expressed as $p(a|q)$.
-$p(a|q)$ $=\sum_{\mathcal{Z}\subseteq\mathcal{G}}p_{\phi}(a|q,\mathcal{Z})p_{\theta}(\mathcal{Z}|q,\mathcal{G})$ (1)
-$\approx p_{\phi}(a|q,\mathcal{Z}^{*})p_{\theta}(\mathcal{Z}^{*}|q,\mathcal{G}),$
-where $\mathcal{Z}^{*}=\text{argmax}_{\mathcal{Z}\in\mathcal{G}}p_{\theta}(\mathcal{Z}|q,\mathcal{G})$ represents the optimal subset of retrieved knowledge from the KGs, and $p_{\phi}(a|q,\mathcal{Z}^{*})$ represents the likelihood generating the response given the query $q$ and the retrieved knowledge $\mathcal{Z}^{*}$, parameterized by $\phi$ and $\theta$.
-Due to the potential exponential growth in the number of possible retrieved knowledge with graph size, efficient approximation methods are required.
-Consequently, the first line of Eq.
-1 is approximated using the most probable retrieved knowledge $\mathcal{Z}^{*}$.
 3.2 Graph-guided retrieval (G-retrieval)
-In this section, we explore which forms of knowledge (e.g., triplets, paths, or subgraphs) should be extracted from KGs to enhance the LLMs and discuss the design of G-retrieval.
-The phase begins by encoding the user query into an embedding vector $\mathbf{q}\in\mathbb{R}^{d}$ using a language model encoder $f_{\text{LM}}(\cdot)$.
-For each node $v\in\mathcal{V}$ and edge $e\in\mathcal{E}$, their respective embeddings $\mathbf{e}_{v}$ and $\mathbf{e}_{e}$ are similarly computed using the same model, ensuring all the information is represented in a unified vector space.
-The relevance of each node or edge to the query is measured using cosine similarity as: $s_{v}=\frac{\mathbf{q}\cdot\mathbf{e}_{v}}{\|\mathbf{q}\|\|\mathbf{e}_{v}\|}$ and $s_{e}=\frac{\mathbf{q}\cdot\mathbf{e}_{e}}{\|\mathbf{q}\|\|\mathbf{e}_{e}\|}$.
-Based on these similarity scores, the most relevant nodes and edges are identified to form a set of top-$k$ nodes and edges for the query.
-To balance retrieving as many relevant nodes and edges as possible while keeping the size manageable, we adopt a prize-cost trade-off strategy: (1) Prize assignment: based on the computed similarity scores, we assign prizes to nodes and edges to quantify their relevance to the query.
+Our goal in G-retrieval is to extract a compact yet informative subset $\mathcal{Z}^{*}$ from the KG that best supports answering the user query $q$.
+We first encode the query and all KG elements (nodes/edges) into a unified embedding space using a language model.
+We then measure the relevance of each element to $q$ (e.g., via cosine similarity) and identify a set of top-$k$ nodes and edges for the query.
+To balance retrieving as many relevant nodes and edges as possible while keeping the $\mathcal{Z}^{*}$ size manageable, we adopt a prize-cost trade-off strategy Balas (1989) to guide the retrieval process: (1) Prize assignment: based on the computed similarity scores, we assign prizes to nodes and edges to quantify their relevance to the query.
 Specifically, we assign the top-$k$ nodes/edges with descending prize values from $k$ to 1, while nodes and edges outside the top-$k$ receive a prize of 0.
@@
 (2) Cost allocation: to manage the retrieved knowledge size, we assign penalties as cost $C_{e}$ during the expansion of the retrieved paths or subgraphs.
-These prizes and costs will be used in the retrieval process, by aiming to maximize the prizes while minimizing costs.
-Triplet-retrieval: the triplet retrieval approach retrieves a fixed number $k$ of triplets with the highest total prize assigned to their respective head entity, relation, and tail entities.
-Path-retrieval: based on the defined prize assignment and cost allocation, the path-retrieval approach starts from a fixed number of $k$ of high-prize nodes to construct sequences $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$ by aiming to greedily maximize their scores: $S(\mathcal{P})=\sum_{i=1}^{n}p_{v_{i}}+\sum_{i=1}^{n-1}p_{e_{i}}-\sum_{i=1}^{n-1}c_{e}$.
-We use a priority queue to iteratively extend paths by selecting the highest-scoring extensions while avoiding cycles and adhering to maximum path length and quantity constraints.
-Finally, the top-scoring paths are sorted and returned.
+The final retrieval process aims to maximize the total prize (i.e., relevance) while minimizing associated costs.
+We explore three retrieval variants for G-retrieval design (e.g., triplets, paths and subgraphs) as demonstrated in Figure 3.
+  • Triplet-retrieval: retrieves a fixed number of triplets with the highest total prize assigned to their respective triplets.
+  • Path-retrieval: starting from a fixed number of $k$ of high-prize nodes, we greedily expand paths $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$ to maximize score: $S(\mathcal{P})=\sum_{i=1}^{n}p_{v_{i}}+\sum_{i=1}^{n-1}p_{e_{i}}-\sum_{i=1}^{n-1}c_{e}$.
+We use a priority queue to iteratively return paths with top-scores and subject to maximum lengths and cycles.
 The details of path-retrieval can be found in Appendix A.3.
-Sub-graph retrieval: the sub-graph retrieval approach also uses the prize assignment and cost allocation, following He et al.
-(2024a) to optimize the process using Prize-Collecting Steiner Tree (PCST) algorithm.
-The PCST problem seeks to build a connected subgraph $\mathcal{S}$ that maximizes the total prizes of included nodes and edges while minimizing the total cost of the edges.
-The score of the subgraph is defined as $S(\mathcal{S})=\sum_{n\in V_{\mathcal{S}}}p_{v_{i}}+\sum_{e\in E_{\mathcal{S}}}p_{e_{i}}-\sum_{einE_{\mathcal{S}}}c_{e}$.
+  • Sub-graph retrieval: building on He et al.
+(2024a), we use the Prize-Collecting Steiner Tree (PCST) algorithm to find a connected subgraph $\mathcal{S}$ that maximizes $S(\mathcal{S})=\sum_{n\in V_{\mathcal{S}}}p_{v_{i}}+\sum_{e\in E_{\mathcal{S}}}p_{e_{i}}-\sum_{einE_{\mathcal{S}}}c_{e}$.
 Unlike in path-retrieval, we only yield one subgraph that maximizes the total score.
 3.3 Graph-guided Generation (G-Generator)
-After retrieving $\mathcal{Z}^{*}$, the graph-guided generation phase learns to use this knowledge to generate the corresponding output to answer the user query.
-The generation is modeled as a sequential decision-making process, where at each time step $t$, the next token $a_{t}$ is generated by G-Generator, conditioned on the query $q$, the retrieved knowledge $\mathcal{Z}^{*}$, and all previously generated tokens $a_{0:t-1}$.
-$p(a|q,\mathcal{Z}^{*})=\prod_{t=1}^{T}p_{\theta}\left(a_{t}|q,\mathcal{Z}^{*},a_{0:t-1}\right)\vspace{-2mm}$ (2)
-The model learns the parameters $\theta$ to maximize the likelihood of generating the correct output sequence by computing the probability distribution over possible next tokens at each time step.
-This ensures that the generated output not only considers the previously generated tokens but also the retrieved knowledge.
-The generation continues until a termination condition is met, such as generating an end-of-sequence token or reaching the maximum sequence length $T$.
+After retrieving $\mathcal{Z}^{*}$, the G-Generator use this knowledge to generate the paragraph-like response the user query.
+The generation is modeled as a sequential decision-making process: at each time step $t$, token $a_{t}$ is generated conditioned on $q$, $\mathcal{Z}^{*}$, and the previously generated tokens $a_{0:t-1}$:
+$p(a|q,\mathcal{Z}^{*})=\prod_{t=1}^{T}p_{\theta}(a_{t}|q,\mathcal{Z}^{*},a_{0:t-1}),$ (3)
+where $\theta$ denotes the parameters of a neural text generation model. The generation stops when an end-of-sequence token is produced or when the maximum sequence length $T$ is reached.
 4 Experiments
-4.1 Evaluation Metrics
-To quantify the hallucinations of LLMs, we leverage two public metrics, FActScore (Min et al., 2023) (factual precision in atomicity score) and SAFE (Wei et al., 2024) (search-augmented factuality evaluator).
-FActScore is designed to measure factual precision in text by decomposing a long form generation into multiple atomic facts and validates each separately against a reliable knowledge base, such as Wikipedia.
-We measure the proportion of facts that are supported by the knowledge source out of the total atomic facts.
-SAFE, on the other hand, adopts a more dynamic approach to fact-checking.
-It employs a language model as an investigative agent, which iteratively uses Google Search queries and reasons about whether the search results support or do not support the fact.
-Figure 4: Comparison results of different forms of information over different queries.
-Table 2: Comparison results of various forms of information extracted from the KGs.
+In this section, we first introduce the evaluation metrics, and then focus on two main research questions: RQ1: Can KGs reduce hallucination in LLMs?
+and RQ2: How are KG-Aware methods affected by noise/perturbations in KGs?
+4.1 Evaluation Metrics & Setup
+We quantify LLM hallucinations using two public metrics: FActScore (Min et al., 2023) and SAFE (Wei et al., 2024).
+FActScore measures factual precision by decomposing a long-form text into atomic facts and validating each against a reliable knowledge base like Wikipedia.
+In contrast, SAFE employs a language model as an investigative agent that iteratively employs Google Search queries to assess whether search results support the fact.
+For both metrics, we report the proportion of supported atomic facts out of the total atomic facts extracted from LLM responses.
+In addition to the hallucination metrics, we propose four metrics using LLM-as-evaluator (Li et al., 2024) to quantify the quality of generated responses from LLM (Edge et al., 2024; Wang et al., 2023).
+In specific, we use G-Eval (Liu et al., 2023) framework for the evaluation and provide relevant Wikipedia pages of each query as context to enhance G-Eval’s robustness and stability.
+The four metrics are defined as follows: Context Relevance: measures how well the generated response aligns with the provided context.
+Comprehensiveness: assesses how thoroughly the answer addresses all aspects and details of the question.
+Correctness: measures the clarity and specificity of the generated answer to the question.
+Empowerment: evaluates how well the generated answer helps the reader understand the topic and make informed decisions.
+The detailed prompt can be found in Appendix D.
+We use gpt-4o-mini (from November 2024 to January 2025) as LLM backbone for all the evaluation metrics.
+As using LLM-as-evaluator frameworks may raise concern regarding potential self-enhancement or bias from the selection of the models Gu et al.
+(2024); Li et al.
+(2024), we conduct additional validation in Appendix A.4 (including human evaluation alignment and cross-validation across different LLM backbones), and find that the choice of LLM in the LLM-as-evaluator framework has little impact on the overall evaluation and demonstrate high correlation with the human evaluation, supporting the reliability of our testing.
+For other testing LLM backbones mentioned in this section, we consider a range of widely used LLMs of different scales, including GPT-4o, GPT-4o-mini (from November 2024 to January 2025), Llama-3.1-8B-instruct (Dubey et al., 2024), Mistral-7B-instruct-v0.3 (Jiang et al., 2023a), and Gemma-2-9B-it (Team et al., 2024).
+Considering the trade-off between cost and performance, we use text-embedding-3-small model from OpenAI (from November 2024 to January 2025) as embedding model for G-retrieval process.
 Models G-Eval Hallucination
@@
 gemma-2-9b-it $55.24\%\pm 1.49\%$ $52.27\%\pm 1.21\%$ $47.14\%\pm 0.36\%$ $50.36\%\pm 0.51\%$ $78.00\%\pm 0.77\%$ $44.32\%\pm 1.58\%$
+Baseline: With Wikipedia documents
+GPT-4o - IRCoT $73.12\%\pm 0.32\%$ $69.23\%\pm 0.42\%$ $66.33\%\pm 0.34\%$ $65.51\%\pm 0.11\%$ $87.39\%\pm 0.68\%$ $69.45\%\pm 0.34\%$
+GPT-4o-mini - IRCoT $70.31\%\pm 0.32\%$ $64.42\%\pm 1.31\%$ $61.37\%\pm 0.48\%$ $63.89\%\pm 0.72\%$ $84.72\%\pm 0.48\%$ $65.72\%\pm 1.03\%$
 Var-1: With CoT Prompting
@@
 GPT-4o - CoT+SC $75.81\%\pm 0.65\%$ $71.62\%\pm 0.74\%$ $66.55\%\pm 0.75\%$ $68.74\%\pm 0.15\%$ $79.03\%\pm 0.48\%$ $53.23\%\pm 0.78\%$
-llama-3.1-8b-instruct - CoT $61.54\%\pm 0.95\%$ $58.35\%\pm 1.05\%$ $53.31\%\pm 0.71\%$ $56.42\%\pm 0.83\%$ $77.07\%\pm 0.85\%$ $46.15\%\pm 0.54\%$
 llama-3.1-8b-instruct - CoT+SC $63.69\%\pm 0.32\%$ $60.44\%\pm 0.59\%$ $55.46\%\pm 0.52\%$ $58.53\%\pm 1.11\%$ $76.00\%\pm 0.63\%$ $45.05\%\pm 0.97\%$
-mistral-7B-Instruct-v0.3 - CoT $59.58\%\pm 0.43\%$ $56.23\%\pm 2.31\%$ $51.28\%\pm 1.31\%$ $54.33\%\pm 0.72\%$ $75.04\%\pm 0.95\%$ $43.03\%\pm 1.03\%$
 mistral-7B-Instruct-v0.3 - CoT+SC $61.35\%\pm 0.93\%$ $58.33\%\pm 1.02\%$ $53.42\%\pm 0.79\%$ $56.47\%\pm 0.85\%$ $74.30\%\pm 0.21\%$ $42.00\%\pm 0.29\%$
-gemma-2-9b-it - CoT $57.34\%\pm 1.05\%$ $54.12\%\pm 0.32\%$ $49.27\%\pm 0.85\%$ $52.12\%\pm 0.95\%$ $72.07\%\pm 1.05\%$ $40.13\%\pm 0.49\%$
 gemma-2-9b-it - CoT+SC $59.42\%\pm 0.27\%$ $56.27\%\pm 0.84\%$ $51.34\%\pm 1.42\%$ $54.34\%\pm 1.31\%$ $71.09\%\pm 0.43\%$ $39.85\%\pm 1.03\%$
@@
 […]
-In addition to the hallucination metrics, we also propose four metrics to quantify the desirable qualities for generating sensible responses (Edge et al., 2024; Wang et al., 2023).
-Given the complexities inherent in an open-ended setting, collecting and annotating ground-truth answers poses significant challenges.
-Therefore, we leverage LLMs as automatic reference-free evaluators to assess these metrics based on G-Eval (Liu et al., 2023), a framework of using LLMs with chain-of-thought (CoT) and a form-filling paradigm.
-To enhance G-Eval’s robustness and stability, we provide the relevant Wikipedia pages as context in the prompt when calculating the metrics.
-The metrics are defined as follows: Context Relevance: measures how well the generated response aligns with the provided context.
-Comprehensiveness: assesses how thoroughly the answer addresses all aspects and details of the question.
-Correctness: measures the clarity and specificity of the generated answer to the question.
-Empowerment: evaluates how well the generated answer helps the reader understand the topic and make informed decisions.
-The detailed prompt can be found in Appendix C.
-Models.
-We incorporate a range of widely used LLMs of different scales, including GPT-4o, GPT-4o-mini, Llama-3.1-8B-instruct (Dubey et al., 2024), Mistral-7B-instruct-v0.3 (Jiang et al., 2023a), and Gemma-2-9B-it (Team et al., 2024).
-We use the embedding model (text-embedding-3-small) from OpenAI for the G-retrieval process, considering cost and performance.
+Table 2: Comparison results of various forms of information extracted from the KGs.
 4.2 RQ1: Main Results - Can KGs Reduce Hallucination in LLMs?
@@
 We use zero-shot and few-shot prompting as baselines without injecting external knowledge.
-In addition, compared to adding external knowledge from KGs, we also consider leveraging LLMs’ internal knowledge to do Chain-of-thought (Kim et al., 2023), or self-consistency (Wang et al., 2022) as baselines.
-For LLMs augmented with KGs, we consider three KG retrieval forms: triplets, paths, and subgraphs (as shown in Figure 3) to study the impact of G-retrieval for reducing LLMs’ hallucinations.
+In addition, we consider leveraging LLMs’ internal knowledge to do Chain-of-thought (Kim et al., 2023), or self-consistency (Wang et al., 2022), and more general RAG systems like IRCoT (Trivedi et al., 2022a) which retrieves paragraphs from Wikipedia to augment CoT generation.
+For LLMs augmented with KGs, we consider three KG retrieval variants: triplets, paths, and subgraphs to study the impact of G-retrieval for reducing LLMs’ hallucinations.
 The results are shown in Table 2 and Figure 4.
 We obtain some intriguing findings:
+Figure 4: Comparison results of different forms of information over different queries.
 Retrieving KG information can indeed mitigate factual errors in the responses.
@@
 The findings also apply to open-source models like mistral-7B-Instruct-v0.3 and Llama-3.1-8B-instruct, illustrating the consistency of the finding.
+In addition, compared with RAG method IRCoT Trivedi et al.
+(2022b), leveraging Wikipedia documents, improves performance over zero-shot and 4-shot prompting by providing broad contextual support, it struggles with correctness and hallucination control due to the potential introduction of irrelevant or conflicting information.
+Our KG-based methods consistently outperform IRCoT, particularly in correctness, SAFE, and FActScore.
 Directly performing reasoning in the LLM itself does not mitigate hallucinations.
-We benchmark the hallucination ratio of LLMs using internal reasoning strategies like Chain-of-thought and Self-consistency, as shown in Var-1 in Table 2.
-It shows that these methods can improve response quality (i.e., correctness, comprehensiveness, and empowerment) compared to baselines, but do not consistently improve factuality, and sometimes even diminish factuality.
-This shows that relying solely on internal reasoning does not mitigate hallucination, emphasizing the importance of external knowledge to counter the hallucination issue.
+We benchmark the hallucination ratio of LLMs using internal reasoning strategies like CoT and Self-consistency, as shown in Var-1 in Table 2.
+It shows that these methods can improve response quality (i.e., G-Eval) compared to baselines, but do not consistently improve factuality, and sometimes even diminish.
+This shows that relying solely on internal reasoning is inadequate for mitigating hallucinations, highlighting the necessity for external knowledge to address this issue effectively.
 (a) Edge Deletion
@@
 We demonstrate the performance of different retrieval methods across different query types in Figure 4, showing that subgraphs achieve the best performance.
-Especially for simpler queries (“Character Description” and “Event Description” which do not require intensive reasoning), subgraph retrieval methods achieve better performance compared to paths and triplets.
+Especially for simpler queries (“Character Description” and “Event Description” which do not require intensive reasoning).
 Even for queries like “Relationship Explanation” and “Cause Explanation” which require stepwise reasoning, subgraph methods still demonstrate promising performance.
@@
 For the edge deletion methods, only if the perturbation level reaches 1.0, the ATS goes to 0, otherwise, the ATS remains higher compared to other settings.
-Relation replacement has a relatively stronger impact on ATS, i.e., on the semantics of the perturbed KGs.
-For SC2D and SD2 (which intuitively measure structural similarity), the four methods show similar levels of decrease in most cases.
 Figure 6 illustrates the hallucination ratio using different methods on the perturbed KGs.
@@
 However, it still outperforms the baseline when the perturbation level is smaller than 40%.
-Our findings on OKGQA-P illustrate that the effectiveness of KG-derived information diminishes with a perturbation level at 50%, surpassing this level leads to a further decrease in performance.
+In summary, we find that effectiveness of KG-derived information diminishes with a perturbation level at 50%, surpassing this level leads to a further decrease in performance.
 We think that before this perturbation level at 50%, incorporating external knowledge from KGs can mitigate hallucinations in LLMs compared to baseline using CoT.
-In practical scenarios, platforms like Wikidata are less likely to be severely perturbed than 50% due to ongoing updates and community-based quality control, ensuring the relevance of our findings in real-world applications.
+Considering practical scenarios, platforms like Wikidata are less likely to have perturbations as severe as 50% due to their ongoing updates and community-based quality control.
+This ensures the relevance and applicability of our findings in real-world settings.
 5 Conclusion
@@
 Our results underscore the significance of integrating KGs with LLMs to help reduce hallucination of LLMs, even in circumstances where the KGs are contaminated.
-Limitations
+6 Limitations
 Our proposed benchmark primarily use DBpedia as the knowledge source, which may not generalize well to testing scenarios requiring highly specialized or domain-specific knowledge.
@@
 In dynamic environments where knowledge is continuously updated, maintaining and integrating real-time changes remains a challenge and may requires further design.
-Appendix A Appendix
-A.1 Intuition of Open-ended Questions?
-In previous close-ended KGQA research Mihaylov et al.
-(2018); Yih et al.
-(2016); Talmor and Berant (2018), “hallucination” has not been considered at all.
-The primary reason is that close-ended KGQA models typically generate concise entity names as answers, which are directly compared against ground truth labels using metrics such as precision or accuracy.
-Any incorrect entity name is immediately reflected in the evaluation results, leaving little room for hallucination to manifest as a separate problem.
-While effective for simple tasks, we believe this closed-ended setting is too naive and fails to generalize to real-world, complex QA situations, where models must provide more detailed, nuanced responses Kantharaj et al.
-(2022).
-In contrast, our work focuses on open-ended KGQA, where LLMs are prompted to generate more elaborate answers that include reasoning paths and supporting facts derived from the background KG.
-This broader output space not only provides a larger field to assess hallucination but also increases the likelihood of factual inaccuracies, as LLMs tend to produce more errors when generating longer and more complex responses.
-While hallucination is minimal in close-ended KGQA due to the short and factual nature of entity names, open-ended KGQA highlights the susceptibility of LLMs to hallucinate when generating rich, explanatory answers.
-By introducing this open-ended setting, OKGQA, we (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on method design.
-We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs to reduce hallucination.
-A.2 Query Construction
+Appendix A Implementation Details
+A.1 Query Construction
 In this section, we discuss the details of the query construction of OKGQA.
-It follows a template-based method with LLMs to generate a diverse range of queries.
-To ensure that the generated queries accurately represent real-world scenarios and complexities, we integrate the corresponding Wikipedia pages of each entity as contextual information for the LLMs’ generation.
-Furthermore, we follow a human-in-the-loop process to optimize the instruction used for generation, as shown in Figure 7.
-Intuitively, starting with an initial instruction, we generate a large number of query candidates, then use an LLM as an automatic evaluator to evaluate the quality of the query, denoted as a set of scores from different metrics $s_{\text{auto}}$ ranging from 1 to 10, with higher scores indicating better performance.
-Then, these queries are manually evaluated and assigned a set of human-label scores $s_{\text{human}}$ corresponding to $s_{\text{auto}}$, normalized to the same range as $s_{\text{auto}}$.
-We then optimize the input instructions by iteratively generating instructions to minimize the gap between $s_{\text{human}}$ and $s_{\text{auto}}$.
+We first introduce the human-in-the-loop process to optimize the instruction for generating the queries, as shown in §A.1.1.
+We then present the metrics for quantify the generated queries in §A.1.2.
+Subsequently, we provide experiments results of human-in-the-loop process and demonstrate the Pearson correlation coefficients between human and LLM scores across rounds of optimization, and verify the inter-rater reliability across our evaluators in §A.1.3.
+A.1.1 Human-in-the-loop for instruction optimization
+To ensure that the generated queries accurately represent real-world scenarios and complexities, we propose a human-in-the-loop process to optimize the instruction used for generation, as shown in Figure 7.
+To ensure clarity, we summarize this optimization process here:
+  • Step 1: Generate a set of queries from an initial instruction.
+  • Step 2: Collect automatic evaluation scores $s_{\mathbf{auto}}$ by LLMs and human-label scores $s_{\mathbf{human}}$ by human annotators for these queries (normalized to the same range).
+  • Step 3: Identify patterns of discrepancies between these scores.
+  • Step 4: Let the LLM analysis the identified patterns to generate new instructions,
+The step 3 and 4 are conducted by prompting LLM with prompt specified in §D.3, and steps 1 to 4 are running iteratively to reducing $s_{\mathbf{auto}}$ and $s_{\mathbf{human}}$ discrepancies.
 This process quite mimics the way of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and inherits the benefit that labeling the reward of the LLMs’ output is much easier than labeling the output directly.
 Figure 7: Human-in-the-loop of query construction.
-Specifically, we consider five angles to measure the quality of the generated queries: (1) Naturalness: assessing how fluid and human-like the query sounds; (2) Relevance: measuring whether the query pertains directly to the entity and the context provided; (3) Specificity: determining the level of detail and granularity included in the query, ensuring it is not too broad or vague; (4) Novelty: evaluating the uniqueness of the query, ensuring it is not just a repetitive or common question; (5) Actionability: gauging whether the query prompts clear, definite answers or actions that are feasible within the given context.
+A.1.2 Metrics for generated queries
+We consider five metrics to measure the quality of the generated queries: (1) Naturalness: assessing how fluid and human-like the query sounds; (2) Relevance: measuring whether the query pertains directly to the entity and the context provided; (3) Specificity: determining the level of detail and granularity included in the query, ensuring it is not too broad or vague; (4) Novelty: evaluating the uniqueness of the query, ensuring it is not just a repetitive or common question; (5) Actionability: gauging whether the query prompts clear, definite answers or actions that are feasible within the given context.
 Each of these angles contributes to a holistic evaluation of the query’s effectiveness and relevance in real-world applications.
-We have three evaluators participating in the manual assessment of query quality.
+A.1.3 Verifying human-in-the-loop
+For the human-label scores $s_{\mathbf{human}}$ collection, we have three evaluators participating in the manual assessment of query quality.
 All of the evaluators are computer science majors with fluent English skills.
@@
 It shows that as the rounds progress, agreement between humans and LLMs increases, suggesting that iterative feedback improves alignment between human annotation and LLM responses.
-Table 3: Pearson correlation coefficients between human and LLM scores across rounds.
 Metric Round 1 Round 2 Round 3 Round 4
@@
 Actionability 0.33 0.41 0.48 0.53
+Table 3: Pearson correlation coefficients between human and LLM scores across rounds.
 In addition, we also consider verifying the inter-rater reliability across three evaluators as shown in Table 4.
@@
 Meanwhile, Actionability performs slightly better, nearing the “Substantial” range, indicating moderately consistent criteria.
-Table 4: Cohen’s Kappa coefficient for various metrics.
 Metric Evaluator 1 & 2 Evaluator 1 & 3 Evaluator 2 & 3
@@
 Actionability 0.67 0.65 0.68
-A.3 Personalized PageRank (PPR)
+Table 4: Cohen’s Kappa coefficient for various metrics.
+A.2 Personalized PageRank (PPR)
 In this section, we discuss the details of the PPR algorithm used in §2.1 to prune the graph from DBPedia and concentrate on nodes most pertinent to the central nodes of interest.
@@
 […]
-A.4 Prize-Cost-based Path Retrieval
-The path-retrieval method is designed to construct and evaluate paths in a graph based on predefined prize assignments and cost allocations.
+Setting Context Relevance Comprehensiveness Correctness Empowerment SAFE FActScore
+OKGQA (subgraphs) $75.70\%\pm 0.44\%$ $71.51\%\pm 0.83\%$ $66.43\%\pm 0.76\%$ $69.60\%\pm 0.65\%$ $88.71\%\pm 0.72\%$ $70.12\%\pm 0.87\%$
++ Multi-lingual context $75.14\%\pm 0.33\%$ $72.32\%\pm 0.19\%$ $66.72\%\pm 0.74\%$ $70.32\%\pm 0.57\%$ $90.32\%\pm 0.48\%$ $72.83\%\pm 0.93\%$
+Table 5: Comparison of GPT-4o-mini Performance Using Monolingual and Multilingual Subgraphs
+A.3 Prize-Cost-based Path Retrieval
+In this section, we detail the path-retrieval method used in §3.2.
+It is designed to construct and evaluate paths in a graph based on predefined prize assignments and cost allocations.
 The objective is to form sequences of nodes and edges, represented as $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$, that maximize the overall score and minimize the costs.
@@
 When no more paths can be added or the priority queue is empty, the algorithm sorts the paths by their scores from highest to lowest.
+A.4 LLM Evaluation Clarity
+To address the concern regarding potential self-enhancement bias in LLM-as-evaluator frameworks, we have conducted extensive validation of our evaluation approach.
+In specific, we randomly sample 100 questions and evaluated them using three different LLMs (gpt-4o-mini, llama-3.1-8b-instruct, and gemma-2-9b-it).
+We measured inter-model agreement using Cohen’s Kappa as shown in Table 6, which showed substantial to almost perfect consistency.
+This indicates that the evaluation results are consistent across different LLMs, even when the model generating the responses is not the same as the one evaluating them (e.g., using gpt-4o-mini for generation and llama-3.1-8b-instruct for evaluation).
+These findings confirm that the evaluation is robust and independent of the specific LLM used as the evaluator.
+Metric LLM 1 & 2 LLM 1 & 3 LLM 2 & 3
+G-Eval 0.84 0.81 0.82
+FactScore 0.78 0.74 0.78
+SAFE 0.74 0.70 0.72
+Table 6: Cohen’s Kappa coefficient for different LLM pair comparisons.
+For the G-Eval, we use the average score of four sub-metrics for better readability.
+LLM 1: gpt-4o-mini; LLM 2: llama-3.1-8b-instruct; LLM 3: gemma-2-9b-it)
+In addition, we also collect human evaluations for these 100 samples.
+Three experts annotators rate each anonymized response on context relevance, comprehensiveness, correctness and empowerment using a 1-5 Likert scale.
+The average human ratings are computed and compared with automated scores using G-Eval.
+The Pearson’s correlation yields a score of 0.78, indicating strong alignment between human judgment and LLM-based evaluation.
+Combined with the inter-model agreement shown in Table 6, these results confirm that our evaluation is robust, consistent, and largely independent of the specific LLM used as the evaluator.
 A.5 KG Similarity Metrics
-To measure how much the perturbed KGs are deviated from the original KG used in §2. We re-present the according metrics from (Raman et al., 2020) below:
-Aggregated Triple Score (ATS)  ATS measures semantic similarity between two KGs.
+In this section, we introduce the metrics used in §2 to measure the deviation of the perturbed KGs from the original KG.
+These metrics are adapted from  Raman et al.
+(2020) as presented below.
+ATS is mainly used to measure the semantic similarity between two KGs, while SC2D and SD2 are used to measure the structural similarity.
+Aggregated Triple Score (ATS):
+ATS measures semantic similarity between two KGs.
 Let $s_{\mathcal{G}}$ be an edge (triple) scoring function, such that $s_{\mathcal{G}}(e_{1},r,e_{2})$ measures how likely edge $(e_{1},r,e_{2})$ is to exist in $\mathcal{G}$.
@@
 Note that $f_{\text{ATS}}(\mathcal{G},\mathcal{G})$ may not equal 1, since $s_{\mathcal{G}}$ may not perfectly generalize to KGs beyond those it was trained on.
-Similarity in Clustering Coefficient Distribution (SC2D)  SC2D measures structural similarity between two KGs and is derived from the local clustering coefficient (Saramäki et al., 2007; Onnela et al., 2005; Fagiolo, 2007).
+Similarity in Clustering Coefficient Distribution (SC2D):
+SC2D measures structural similarity between two KGs and is derived from the local clustering coefficient (Saramäki et al., 2007; Onnela et al., 2005; Fagiolo, 2007).
 For a given entity in $\mathcal{G}$ (treated here as undirected), the local clustering coefficient is the fraction of possible triangles through the entity that exist (i.e., how tightly the entity’s neighbors cluster around it).
@@
 SC2D is defined as $f_{\text{SC2D}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=1-\frac{\lVert\mathbf{c}_{o}-\mathbf{c}_{p}\rVert_{2}}{\lVert\mathbf{c}_{o}-\mathbf{c}_{p}\rVert_{2}+1}\in[0,1]$, with higher value indicating higher similarity.
-Similarity in Degree Distribution (SD2)  SD2 also measures structural similarity between two KGs, while addressing SC2D’s ineffectiveness when the KGs’ entities have tiny local clustering coefficients (e.g., the item KG used by recommender systems is roughly bipartite).
+Similarity in Degree Distribution (SD2):
+SD2 also measures structural similarity between two KGs, while addressing SC2D’s ineffectiveness when the KGs’ entities have tiny local clustering coefficients (e.g., the item KG used by recommender systems is roughly bipartite).
 In such cases, SC2D is always close to one regardless of the perturbation method, thus rendering SC2D useless.
@@
 SD2 is defined as $f_{\text{SD2}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=1-\frac{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}}{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}+1}\in[0,1]$, with higher value indicating higher similarity.
-Appendix B Related Work
+Appendix B Extension of OKGQA
+In this section, we extend our benchmark by incorporating multilingual context and validating our query generation against DBpedia’s structure.
+We first introduce the multilingual setup of our dataset anc compare the performance of multilingual subgraphs with the monolingual subgraphs (§B.1).
+We then analyze the relationship between generated queries and DBpedia by examining query generation, entity/relation coverage, and subgraph alignment (§B.2).
+We also compare OKGQA with the existing widely used benchmarks in Table 7.
+B.1 Multilingual Setup of OKGQA
+KGs typically include entities and relations in multiple languages, providing a richer context that can benefit our OKGQA setting.
+In this experiment, we investigate whether incorporating multilingual context improves overall performance.
+Specifically, we randomly sample 300 queries from our dataset and generate subgraphs that include multilingual entities and relations from DBpedia.
+We then apply PPR consistent with our original method in §2.1 to reduce the KG size.
+For this multilingual setting, we consider five languages—Greek, Polish, Portuguese, Spanish, and English—which cover the majority of entities in DBpedia.
+We compare the performance of GPT-4o-mini using the new multilingual subgraphs against the original monolingual subgraphs, as shown in Table 5.
+Our findings indicate that including multilingual context generally leads to better performance across multiple metrics.
+Intuitively, this additional multi-lingual context may provides more knowledge from different perspectives (which could provide more context, but also may requires more techniques for handle challenges like duplicates across languages) and also provide another way to validate the factuality of the resources stored in the KGs (which can provide more authenticity through cross validation from different languages).
+B.2 Generated Query-DBpedia Alignment
+We analyze the alignment between our generated queries and DBpedia along three dimensions: query generation, entity/relation coverage, and subgraph alignment as follows:
+Query Generation:
+Each query is directly generated from DBpedia entities and their relationships.
+For example, when asking about Microsoft’s founder, we first confirm that both “Microsoft” and “Bill Gates” exist in DBpedia and are connected by the founded_by relation, ensuring that our queries are firmly grounded in the knowledge graph.
+Entity and Relation Coverage:
+Our analysis indicates that:
+  • 92% entities mentioned in the queries can be detected from DBpedia entities.
+  • 87% queries have complete relation paths connecting the relevant entities from DBPedia.
+  • Entities/relations mentioned in queries cover 72% of DBpedia’s most common entities/predicates and span diverse entity types (e.g., Person, Organization, and Event).
+Subgraph Alignment:
+We evaluate the structure of the sampled subgraphs for each query and find that:
+  • 75% of the queries retrieve subgraphs within 3–4 hops, which aligns with the typical depth for DBpedia reasoning tasks.
+  • On average, each subgraph contains 48 nodes and 152 edges, with an average node degree of 3.17 and a clustering coefficient of 0.69, which also aligns with the property of DBPedia.
+These statistics support that our dataset accurately reflects DBpedia’s structure, ensuring both authenticity and complexity in the generated queries.
+Dataset # Questions Question Type Focus Areas Source of Questions Knowledge Base Hallucination Detection Unreliable KG
+OKGQA 850 / 2,050 Open-ended Evaluating hallucination and reasoning capabilities in LLMs when augmented with Knowledge Graphs; diverse queries requiring complex reasoning Curated DBPedia
+[…]
+WebQuestions 5,810 Factoid Questions derived from Google Suggest queries, focusing on simple factual information User queries Freebase
+[…]
+ComplexWebQuestions 34,689 Multi-hop Factoid Extends WebQuestions with more complex, multi-hop questions requiring compositional reasoning User queries Freebase
+[…]
+GrailQA 64,331 Varied Factoid Evaluates generalization in KBQA with questions requiring different levels of reasoning Crowdsourced Freebase
+Table 7: Comparison of OKGQA with existing benchmarks along with their question types, focus areas, and additional properties.
+Appendix C Related Work
 Due to the stochastic decoding process of Large Language Models (LLMs), i.e., sampling the next token in the sequence, LLMs exhibit probabilistic behaviors: (1) potentially yielding varied outputs of the same input across different instances (Agrawal et al., 2023); (2) cannot accurately interpret phrases or terms when the context is vague and resides in a knowledge gap of the model.
@@
 (2022) propose “rethinking with retrieval” to use decomposed reasoning steps from CoT prompting to retrieve external knowledge, leading to more accurate and faithful explanations.
-IR-CoT (Trivedi et al., 2022) interleaves the generation of CoT with knowledge retrieval from corresponding KGs, iteratively guiding both retrieval and reasoning for multi-step questions.
+IR-CoT (Trivedi et al., 2022b) interleaves the generation of CoT with knowledge retrieval from corresponding KGs, iteratively guiding both retrieval and reasoning for multi-step questions.
 MindMap (Wen et al., 2023) introduce a plug-and-play approach to evoke graph-of-thoughts reasoning in LLMs.
@@
 The benchmark extends the assessment of closed-ended question answering to an open-ended setting, which can further support the assessment of hallucination of LLMs.
-Appendix C Prompt List
+Appendix D Prompt List
 In this section, we present all the prompts required for the main experiments.
 To enhance clarity, we provide only one example in the prompt labeled as Example 1; the other few-shot examples utilized are labeled as Other In-Context Few-shots within the prompt.
-C.1 Knowledge-Augmented Generation
+D.1 Knowledge-Augmented Generation
 System Instruction: “You are a helpful assistant designed to answer the users’ open-ended questions.
@@
 Answer:
-C.2 OKGQA Query Generation Prompting
+D.2 OKGQA Query Generation Prompting
 System Instruction: “Generate open-ended questions about different types: character description, event description, cause explanation, relationship explanation, trend prediction, outcome prediction, contrast analysis, historical comparison, evaluation and reflection, and application and practice.
@@
 Generation:
-C.3 Prompts for Instruction Tuner
+D.3 Prompts for Instruction Tuner
 Act as an “Instruction Tuner” for the LLM, you will be given the inputs: (1) the {Current Instruction} used to guide the LLMs’s evaluation, including specific examples with ground truth labels; (2) {Current Errors} that emerged with this instruction are applied to the dataset.
@@
 Include 2-3 examples at the end of your response to demonstrate how the new instruction would be applied.
-C.4 Metrics Prompt for G-eval
+D.4 Metrics Prompt for G-eval
 System Instruction: “You are a helpful assistant designed to evaluate the quality of the response to a query. Your task is to rate the response on one metric defined as below:”

### v4

@@
 Abstract
-Recent works integrating Knowledge Graphs (KGs) have led to promising improvements in enhancing the reasoning accuracy of Large Language Models (LLMs).
-However, current benchmarks focus mainly on closed-ended tasks, leaving a gap in the assessment of more complex real-world scenarios.
-This gap has also obscured the evaluation of KGs’ potential to mitigate the problem of hallucination in LLMs.
-To fill the gap, we introduce OKGQA, a new benchmark specifically designed to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios.
-OKGQA is designed to closely reflect the complexities of practical applications using questions from different types, and incorporates specific metrics to measure both hallucination ratio and the enhancement in reasoning capabilities.
-To consider the scenario in which KGs may have varying levels of mistakes, we propose another benchmark variant OKGQA-P to assess model performance when the semantics and structure of KGs are deliberately perturbed and contaminated.
-OKGQA aims to (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on method design.
-We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs to reduce hallucination.
+Recent works integrating Knowledge Graphs (KGs) have shown promising improvements in enhancing the reasoning capabilities of Large Language Models (LLMs).
+However, existing benchmarks primarily focus on closed-ended tasks, leaving a gap in evaluating performance on more complex, real-world scenarios.
+This limitation also hinders a thorough assessment of KGs’ potential to reduce hallucinations in LLMs.
+To address this, we introduce OKGQA [^1] Code and data are released at https://github.com/Y-Sui/OKGQA, a new benchmark specifically designed to evaluate LLMs augmented with KGs in open-ended, real-world question answering settings.
+OKGQA reflects practical complexities through diverse question types and incorporates metrics to quantify both hallucination rates and reasoning improvements in LLM+KG models.
+To consider the scenarios in which KGs may contain varying levels of errors, we propose a benchmark variant, OKGQA-P, to assess model performance when the semantics and structure of KGs are deliberately perturbed and contaminated.
+In this paper, we aims to (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on method design.
+We believe this study can facilitate a more complete performance comparison and encourages continuous improvement in integrating KGs with LLMs to mitigate hallucination, and make LLMs more trustworthy.
 Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study Over Open-ended Question Answering
@@
 1 Introduction
-Contemporary LLMs are prone to producing hallucinations due to gaps in their knowledge Gekhman et al.
+Contemporary large language models (LLMs) are prone to producing hallucinations—plausible-sounding but incorrect or irrelevant outputs—due to gaps and inconsistencies in their training data Gekhman et al.
 (2024); Lee et al.
 (2022).
-These inaccuracies commonly stem from misinformation, biases, or errors in the training data, and lead to responses that seem plausible but may be irrelevant or incorrect Weng (2024).
-This issue is particularly concerning in high-stakes contexts such as healthcare He et al.
-(2023) and science Taylor et al.
-(2022) [^1] Code and data are released at https://anonymous.4open.science/r/OKGQA-CBB0.
-Figure 1: Comparison between Open-ended Question and Close-ended Question over Knowledge Graph.
-To address this limitation, researchers have turned to leveraging external knowledge graphs (KGs) as a complementary Pan et al.
+These inaccuracies often arise from misinformation, biases, or errors embedded in the data Weng (2024); Chen et al.
+(2025), posing significant risks in high-stakes domains such as healthcare He et al.
+(2023b) and scientific research Taylor et al.
+(2022).
+Figure 1: Comparison between open-ended and closed-ended questions over knowledge graphs.
+To mitigate these limitations, recent research has explored augmenting LLMs with external knowledge graphs (KGs) Pan et al.
 (2024); Luo et al.
@@
 (2024).
-KGs offer structured and explicit factual information—often domain-specific—and allow each piece of data to be traced back to its source Zheng et al.
+KGs provide structured, explicit, and often domain-specific factual information, with each fact traceable to its original source Zheng et al.
 (2023); Agrawal et al.
 (2023).
-This traceability not only enables verification of the model’s reasoning but also brings transparency to the decision-making process, making KGs a promising method for enhancing the reliability of LLM reasoning.
-Find more details of related works in §C.
-However, current benchmarks for testing these LLM+KG models are predominantly closed-ended Jin et al.
+These properties not only enable verification of the model’s reasoning but also bring transparency to the decision-making process, making KGs a promising avenue for improving the reliability and trustworthiness of LLM outputs.
+A detailed review of related work is provided in Appendix §C.
+However, existing benchmarks for evaluating these LLM+KG models predominantly focus on closed-ended tasks Jin et al.
 (2020); Puerto et al.
-(2023), restricting the model’s output to a limited set of entities, relations Talmor et al.
+(2021), where model outputs are restricted to a fixed set of entities, relations Talmor et al.
 (2019); Mihaylov et al.
-(2018) or logical forms Yih et al.
+(2018), or logical forms Yih et al.
 (2016); Talmor and Berant (2018).
-While these benchmarks are useful to measure retrieval and basic reasoning, they do not adequately capture whether a model is hallucinating.
-In closed-ended settings, errors can stem from incorrect retrieval or from fabricating (hallucinating) answers, yet conventional metrics (e.g., accuracy or precision) cannot distinguish between these two issues.
-This becomes problematic for more complex, real-world applications that demand nuanced answers Kantharaj et al.
+While these benchmarks are useful to measure retrieval accuracy and basic reasoning, such benchmarks fall short in detecting whether a model is hallucinating.
+In closed-ended settings, errors can stem from incorrect retrieval or fabricated (hallucinated) answers, but conventional metrics like accuracy or precision cannot distinguish between these two failure modes.
+This limitation becomes problematic for more complex, real-world applications that demand nuanced answers Kantharaj et al.
 (2022).
-In contrast, our work focuses on open-ended KGQA, where LLMs are prompted to generate more elaborate answers, include reasoning paths and supporting facts derived from the KG (as shown in Figure 1).
-This broader output space offers two key advantages: First, it enables direct assessments of hallucination with metrics like FActScore Min et al.
-(2023) or SAFE Wei et al.
-(2024)), which decompose longer responses into atomic statements for factual consistency checks with external knowledge sources like Wikipedia.
-Second, it increases the likelihood of exposing factual errors which helps to assess the phenomenon indicated in Qiu et al.
-(2024): where longer, more complex responses provide more opportunities for errors to occur.
-By adopting this open-ended approach, we aim to (1) explore whether KGs can make LLMs more trustworthy in the open-ended setting, and (2) conduct a comparative analysis to shed light on methods design and direction for leveraging KGs to reduce LLMs’ hallucination.
-To achieve this, we introduce a new benchmark, Open-ended Knowledge-Graphs Question Answering (OKGQA), specifically designed to assess LLMs enhanced with KGs in an open-ended QA setting.
-OKGQA is designed to closely reflect the complexities of practical applications with diverse questions types as mentioned in Table 1, ensuring that all the queries cannot be answered simply by retrieving isolated KG facts.
-To consider the scenarios for potentially contaminated or imperfect KGs (i.e., attributes may be mislabeled, relations may not exist, etc.), we also propose a variant OKGQA-P (§2.2) to assess model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated.
-In both settings, we measure the degree of hallucination and the overall quality of the responses (see §4.1 for details).
-Based on our experiments, we find that (1) integrating KG information generally mitigates factual errors, especially for queries requiring deep reasoning; (2) directly performing reasoning in the LLM itself (e.g., internal reasoning strategies like Chain-of-thought Kim et al.
+In contrast, our work focuses on open-ended knowledge graph question answering (KGQA), where LLMs are aimed to generate detailed answers that include explicit reasoning paths and supporting facts derived from the KG, as illustrated in Figure 1.
+This expanded output space offers two key advantages.
+First, it enables direct measurement of hallucination using metrics such as FActScore Min et al.
+(2023) and SAFE Wei et al.
+(2024), which decompose complex responses into atomic statements and verify their factual consistency against external knowledge sources like Wikipedia.
+Second, longer and more complex answers increase the likelihood of exposing factual errors, aligning with observations from Qiu et al.
+(2024) that hallucinations tend to accumulate in extended responses.
+By adopting this open-ended paradigm, we aim to (1) explore whether KGs can enhance the trustworthiness of LLMs in realistic, open-ended scenarios, and (2) provide a comparative analysis to guide the design of methods that leverage KGs to reduce hallucination.
+To this end, we introduce Open-ended Knowledge-Graphs Question Answering (OKGQA), a new benchmark tailored to evaluate LLMs augmented with KGs in open-ended QA settings.
+OKGQA reflects the complexities of practical applications by incorporating diverse question types (see Table 1), ensuring that queries cannot be answered by simply retrieving isolated KG facts.
+To simulate real-world conditions where KGs may be imperfect or contaminated—for example, with mislabeled attributes or spurious relations—we propose a benchmark variant, OKGQA-P (§2.2), which evaluates model robustness when KG semantics and structure are deliberately perturbed.
+In both settings, we assess hallucination rates alongside overall response quality (details in §4.1).
+Based on our experiments, we find that (1) integrating KG information generally reduces factual errors, especially for queries requiring deeper reasoning; (2) relying solely on internal LLM reasoning strategies (e.g., Chain-of-Thought Kim et al.
 (2023) and Self-Consistency Wang et al.
-(2022)) may cause bias and hallucination; (3) subgraph-based methods often achieve the best performance for simpler query types; and (4) incorporating KGs effectively reduces hallucinations in LLMs even when the KG is partially contaminated.
+(2022)) can introduce biases and hallucinations; (3) subgraph-based methods often achieve the best performance for simpler query types; and (4) incorporating KGs effectively reduces hallucinations in LLMs even when the KG is partially contaminated.
 2 OKGQA: An Open-ended Knowledge Graph Question-Answering Benchmark
@@
 We then introduce OKGQA-P, a benchmark variant designed to evaluate model robustness under KG perturbations, detailing our perturbation methods and the metrics used to assess semantic and structural deviations.
-Due to the page limitation, we also include some extension of our benchmark in Appendix §B, including multilingual setup, and more analysis.
+Due to page limitations, we provide additional extensions of our benchmark—including a multilingual setup and further analyses—in Appendix §B.
 2.1 Dataset Construction
@@
 Queries.
-We utilize a template-based approach to generate a diverse range of queries using LLMs, including categories such as descriptive, explanatory, predictive, comparative, and critical queries.
-Details regarding specific templates and example queries can be found in Table 1, while the corresponding prompts are provided in the Appendix D.
-To ensure that the generated queries represent real-world scenarios and complexities, we employ an iterative optimization approach that utilizes both automated and human evaluation to refine the query generation process (the details are given in Appendix A.1).
+We utilize a template-based approach to generate a diverse set of queries using LLMs, including categories such as descriptive, explanatory, predictive, comparative, and critical queries.
+Specific templates and example queries are detailed in Table 1, with the corresponding prompts provided in Appendix D.
+To ensure that the generated queries reflect real-world complexity and relevance, we adopt an iterative optimization process that combines automated and human evaluations to refine the query generation process (see Appendix A.1 for details).
 Initially, we generate a diverse set of query candidates from a seed instruction.
-These candidates undergo automated evaluation using an LLM-based evaluator, which assigns quality scores $s_{\text{auto}}$ on a scale of 1-10, with higher scores indicating better performance across multiple metrics.
-Subsequently, human evaluators assess the same queries, producing corresponding normalized scores $s_{\text{human}}$ within the same range.
-To refine the dataset, we iteratively optimize the input instructions by minimizing the discrepancy between $s_{\text{human}}$ and $s_{\text{auto}}$.
-This optimization process ensures alignment between automated and human quality assessments.
-The queries are also categorized by complexity with detailed statistics in Figure 2.
+These candidates are then scored automatically by an LLM-based evaluator, which assigns quality scores $s_{\text{auto}}$ on a scale of 1-10, where higher scores indicate better performance across multiple criteria.
+Subsequently, human evaluators assess the same queries, producing normalized scores $s_{\text{human}}$ within the same range as $s_{\text{auto}}$.
+To refine the query generation, we iteratively optimize the input instructions by minimizing the discrepancy between $s_{\text{human}}$ and $s_{\text{auto}}$, thereby aligning automated assessments with human judgment.
+Additionally, the generated queries are categorized by complexity, with detailed statistics shown in Figure 2.
 KG Sub-graphs.
-To reduce the size of KGs while retaining relevant information, we follow previous work Yih et al.
+To manage the size of KGs while covering relevant context for the queries, we follow previous work Yih et al.
 (2016); Talmor and Berant (2018) by sampling subgraphs from DBpedia (Noted that all queries in OKGQA can be answered using DBpedia).
-We extract all triples contained within the $K$-hop neighbors from the entities mentioned in the query.
-We set $K=2$ to balance graph size and computational feasibility.
+Specifically, we collect all triples within the $K$-hop neighborhood of entities mentioned in each query.
+We set $K=2$ to balance graph coverage and computational feasibility.
 As increasing beyond $2$-hop subgraphs generally leads to exponential growth in edges and nodes Jin et al.
-(2020), which increase excessive noise and complicating information retrieval [^2] This choice is also informed by common practices in other benchmarks, such as WebQSP Yih et al.
+(2020), which increases excessive noise and complicates graph retrieval [^2] This choice aligns with common practices in benchmarks such as WebQSP Yih et al.
 (2016) and CWQ Talmor and Berant (2018), where 2-hop subgraphs are widely used for similar KGQA tasks..
-To further reduce the size of the 2-hop subgraphs, we leverage Personalized Page-Rank (PPR) Bahmani et al.
+To further reduce the size of subgraphs, we leverage Personalized Page-Rank (PPR) Bahmani et al.
 (2010) to prune the nodes/edges that are not relevant to the query (the details of the PPR algorithm are discussed in Appendix A.2).
@@
 To quantify how much the perturbed KG has deviated from the original KG, i.e., $f(\mathcal{G},\mathcal{G}^{\prime})$, we leverage metrics from Raman et al.
-(2020) for capturing semantics (ATS) and structural (SC2D, SD2) similarity between KGs.
+(2020) to evaluate semantics (ATS) and structural (SC2D, SD2) similarity between perturbed KG and original KG.
 Intuitively, ATS leverages a pre-trained LM for link prediction to measure the probability of each edge from $\mathcal{G}^{\prime}$ existing in $\mathcal{G}$, while SC2D and SD2 measure the structural similarity between two KGs based on local clustering coefficient and degree distribution.
-For each of the three metrics, higher value indicates higher similarity.
-The detailed description can be found in Appendix A.5, with visualization in Figure 5.
+For each metric, higher value indicates higher similarity.
+The detailed description of these metrics can be found in Appendix A.5, with corresponding results shown in Figure 5.
 For the perturbation methods, we consider four edge-based perturbation heuristics based on Raman et al. (2020) as follows:
@@
 3.1 Formalization
-We formalize the KG-augmented framework for reducing hallucination as follows.
+We formalize the KG-augmented framework as follows.
 Given a user query $q$, a pretrained language model generates a paragraph-like answer $a$ by modeling the conditional probability $p(a|q)$.
-To explore whether KGs help reduce hallucinations of LLMs, we introduce retrieved knowledge $\mathcal{Z}$ from the KG and define:
+To explore whether KGs help reduce hallucinations of LLMs, we introduce the retrieved knowledge $\mathcal{Z}$ from the KG and define:
 $p(a|q)=\sum_{\mathcal{Z}\subseteq\mathcal{G}}p_{\phi}(a|q,\mathcal{Z})p_{\theta}(\mathcal{Z}|q,\mathcal{G}),$ (1)
@@
   • Sub-graph retrieval: building on He et al.
-(2024a), we use the Prize-Collecting Steiner Tree (PCST) algorithm to find a connected subgraph $\mathcal{S}$ that maximizes $S(\mathcal{S})=\sum_{n\in V_{\mathcal{S}}}p_{v_{i}}+\sum_{e\in E_{\mathcal{S}}}p_{e_{i}}-\sum_{einE_{\mathcal{S}}}c_{e}$.
+(2024), we use the Prize-Collecting Steiner Tree (PCST) algorithm to find a connected subgraph $\mathcal{S}$ that maximizes $S(\mathcal{S})=\sum_{n\in V_{\mathcal{S}}}p_{v_{i}}+\sum_{e\in E_{\mathcal{S}}}p_{e_{i}}-\sum_{einE_{\mathcal{S}}}c_{e}$.
 Unlike in path-retrieval, we only yield one subgraph that maximizes the total score.
@@
 4 Experiments
-In this section, we first introduce the evaluation metrics, and then focus on two main research questions: RQ1: Can KGs reduce hallucination in LLMs?
+In this section, we first introduce the evaluation metrics and experiment setup, and then focus on two main research questions and provide relevant analysis: RQ1: Can KGs reduce hallucination in LLMs?
 and RQ2: How are KG-Aware methods affected by noise/perturbations in KGs?
@@
 For both metrics, we report the proportion of supported atomic facts out of the total atomic facts extracted from LLM responses.
-In addition to the hallucination metrics, we propose four metrics using LLM-as-evaluator (Li et al., 2024) to quantify the quality of generated responses from LLM (Edge et al., 2024; Wang et al., 2023).
+In addition to the hallucination metrics, we propose four metrics using LLM-as-evaluator (Li et al., 20 2) to quantify the quality of generated responses from LLM (Edge et al., 2024; Wang et al., 2023).
 In specific, we use G-Eval (Liu et al., 2023) framework for the evaluation and provide relevant Wikipedia pages of each query as context to enhance G-Eval’s robustness and stability.
-The four metrics are defined as follows: Context Relevance: measures how well the generated response aligns with the provided context.
-Comprehensiveness: assesses how thoroughly the answer addresses all aspects and details of the question.
-Correctness: measures the clarity and specificity of the generated answer to the question.
-Empowerment: evaluates how well the generated answer helps the reader understand the topic and make informed decisions.
-The detailed prompt can be found in Appendix D.
+The four metrics are defined as follows: (1) Context Relevance: measures how well the generated response aligns with the provided context.
+(2) Comprehensiveness: assesses how thoroughly the answer addresses all aspects and details of the question.
+(3) Correctness: measures the clarity and specificity of the generated answer to the question.
+(4) Empowerment: evaluates how well the generated answer helps the reader understand the topic and make informed decisions.
+The corresponding prompts are detailed in Appendix D.
 We use gpt-4o-mini (from November 2024 to January 2025) as LLM backbone for all the evaluation metrics.
-As using LLM-as-evaluator frameworks may raise concern regarding potential self-enhancement or bias from the selection of the models Gu et al.
-(2024); Li et al.
-(2024), we conduct additional validation in Appendix A.4 (including human evaluation alignment and cross-validation across different LLM backbones), and find that the choice of LLM in the LLM-as-evaluator framework has little impact on the overall evaluation and demonstrate high correlation with the human evaluation, supporting the reliability of our testing.
-For other testing LLM backbones mentioned in this section, we consider a range of widely used LLMs of different scales, including GPT-4o, GPT-4o-mini (from November 2024 to January 2025), Llama-3.1-8B-instruct (Dubey et al., 2024), Mistral-7B-instruct-v0.3 (Jiang et al., 2023a), and Gemma-2-9B-it (Team et al., 2024).
+As using LLM-as-evaluator frameworks may raise concern regarding potential self-enhancement or bias from the selection of the backbone models Gu et al.
+(20 2); Li et al.
+(20 2), we conduct additional analysis in Appendix A.4 (including human evaluation alignment and cross-validation across different LLM backbones), and find that the choice of LLMs in the LLM-as-evaluator framework has little impact on the overall evaluation and the results demonstrate high correlation with the human evaluation, supporting the reliability of our testing.
+For our experiments, we consider a range of widely used LLMs of different scales for testing, including GPT-4o, GPT-4o-mini (from November 2024 to January 2025), Llama-3.1-8B-instruct (Dubey et al., 2024), Mistral-7B-instruct-v0.3 (Jiang et al., 2023a), and Gemma-2-9B-it (Team et al., 2024).
 Considering the trade-off between cost and performance, we use text-embedding-3-small model from OpenAI (from November 2024 to January 2025) as embedding model for G-retrieval process.
+To ensure the reproducibility of the experiments, we set $\mathrm{temperature}=0.7$ and $\mathrm{top\_p}=1.0$ for all models.
+We use the API service from OpenAI [^3] https://openai.com/ and OpenRouter [^4] https://openrouter.ai/ for our experiments which host detailed snapshots of the utilized model versions.
 Models G-Eval Hallucination
@@
 Directly performing reasoning in the LLM itself does not mitigate hallucinations.
-We benchmark the hallucination ratio of LLMs using internal reasoning strategies like CoT and Self-consistency, as shown in Var-1 in Table 2.
-It shows that these methods can improve response quality (i.e., G-Eval) compared to baselines, but do not consistently improve factuality, and sometimes even diminish.
+We benchmark the hallucination ratio of LLMs using internal reasoning strategies like CoT and Self-consistency.
+As shown in Var-1 in Table 2, these methods can improve response quality (i.e., G-Eval) compared to baselines, but do not consistently improve factuality, and sometimes even diminish.
 This shows that relying solely on internal reasoning is inadequate for mitigating hallucinations, highlighting the necessity for external knowledge to address this issue effectively.
@@
 4.3 RQ2: How Are KG-Aware Methods Affected by Noise / Perturbations in KGs?
-We benchmark different KG-augmented LLMs on our OKGQA-P setting, where we deliberately perturb and contaminate the semantics and structure of KGs to simulate the real-world situation where KGs may not have high quality.
+We test different KG-augmented LLMs on our OKGQA-P setting, where we deliberately perturb and contaminate the semantics and structure of KGs to simulate the real-world situation where KGs may not have high quality.
 Specifically, we consider different perturbation methods discussed in §2.2 and control the perturbation level based on the percentage of KG edges being perturbed.
@@
 However, it still outperforms the baseline when the perturbation level is smaller than 40%.
-In summary, we find that effectiveness of KG-derived information diminishes with a perturbation level at 50%, surpassing this level leads to a further decrease in performance.
+In summary, we find that the effectiveness of KG-derived information diminishes with a perturbation level at 50%, surpassing this level leads to a further decrease in performance.
 We think that before this perturbation level at 50%, incorporating external knowledge from KGs can mitigate hallucinations in LLMs compared to baseline using CoT.
@@
 In this section, we discuss the details of the query construction of OKGQA.
-We first introduce the human-in-the-loop process to optimize the instruction for generating the queries, as shown in §A.1.1.
-We then present the metrics for quantify the generated queries in §A.1.2.
-Subsequently, we provide experiments results of human-in-the-loop process and demonstrate the Pearson correlation coefficients between human and LLM scores across rounds of optimization, and verify the inter-rater reliability across our evaluators in §A.1.3.
+We first introduce the human-in-the-loop process to optimize the seed instruction for generating the queries, as shown in §A.1.1.
+We then present the metrics to quantify the generated queries in §A.1.2.
+Subsequently, we provide experiments results of human-in-the-loop process and demonstrate the Pearson correlation coefficients between human and LLM scores across different rounds of optimization, and verify the inter-rater reliability across different LLM evaluators in §A.1.3.
 A.1.1 Human-in-the-loop for instruction optimization
-To ensure that the generated queries accurately represent real-world scenarios and complexities, we propose a human-in-the-loop process to optimize the instruction used for generation, as shown in Figure 7.
+To ensure that the generated queries represent real-world scenarios and complexities, we propose a human-in-the-loop process to optimize the seed instruction used for generation, as shown in Figure 7.
 To ensure clarity, we summarize this optimization process here:
@@
   • Step 4: Let the LLM analysis the identified patterns to generate new instructions,
-The step 3 and 4 are conducted by prompting LLM with prompt specified in §D.3, and steps 1 to 4 are running iteratively to reducing $s_{\mathbf{auto}}$ and $s_{\mathbf{human}}$ discrepancies.
-This process quite mimics the way of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and inherits the benefit that labeling the reward of the LLMs’ output is much easier than labeling the output directly.
+Steps 3 and 4 are performed by prompting the LLM with the instruction specified in §D.3, and the entire process from steps 1 to 4 is iterated to minimize the discrepancy between $s_{\mathbf{auto}}$ and $s_{\mathbf{human}}$.
+This procedure closely resembles the way of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and inherits the benefit that labeling the reward of the LLMs’ output is much easier than directly labeling the outputs.
 Figure 7: Human-in-the-loop of query construction.
@@
 In addition, we also consider verifying the inter-rater reliability across three evaluators as shown in Table 4.
-We report the Cohen’s Kappa coefficient for each pair of evaluators as follows.
-Using the Landis and Koch (1977) interpretation guidelines, the Cohen’s Kappa coefficients for Naturalness and Relevance (ranging from 0.79 to 0.85) fall within the “Substantial” to “Almost Perfect” categories, indicating strong inter-rater reliability for these metrics.
-This reflects a shared understanding of the evaluation criteria, resulting in consistent ratings among evaluators.
+We report the Cohen’s Kappa coefficient for each pair of evaluators.
+According to the interpretation guidelines from Landis and Koch (1977), the Cohen’s Kappa coefficients for Naturalness and Relevance (ranging from 0.79 to 0.85) fall within the “Substantial” to “Almost Perfect” categories, indicating strong inter-rater reliability for these metrics.
+This indicates that the evaluators share a common understanding of the evaluation criteria, leading to consistent ratings across evaluators.
 For Specificity, Novelty, and Actionability, the coefficients range from 0.58 to 0.68, placing them primarily in the “Moderate” to “Substantial” categories.
@@
 To efficiently manage the exploration of potential paths, we utilize a priority queue, a data structure that allows paths to be organized based on their scores, ensuring that the highest-scoring paths are processed first.
-The method starts by picking a number of starting nodes with high prizes.
-We then expand these starting points by exploring neighboring nodes.
-For each neighbor, the method calculates a new score.
-This score is the sum of the neighbor’s prize and the edge’s prize minus the edge’s cost.
-If this neighbor hasn’t been visited before, which helps avoid looping, the algorithm adds this neighbor to the path.
-This new path is then added to the priority queue.
-This expansion keeps going until the path reaches a maximum length or can’t be extended further.
-The algorithm keeps track of paths already explored to avoid repetition and ensure paths don’t loop back on themselves.
-When no more paths can be added or the priority queue is empty, the algorithm sorts the paths by their scores from highest to lowest.
+The method starts by picking a number of starting nodes with high prizes.From each starting node, it expands by exploring neighboring nodes.
+For each neighbor, a new score is computed as the sum of the neighbor’s prize and the edge’s prize minus the edge’s cost.
+If this neighbor has not been visited before, the algorithm appends it to the current path and adds this extended path to the priority queue.
+This expansion process continues until paths reach a maximum length or can no longer be extended.
+The algorithm maintains a record of explored paths to avoid repetition and cycles.
+Once no further expansions are possible or the priority queue is empty, the algorithm sorts the collected paths in descending order of their scores.
 A.4 LLM Evaluation Clarity
-To address the concern regarding potential self-enhancement bias in LLM-as-evaluator frameworks, we have conducted extensive validation of our evaluation approach.
-In specific, we randomly sample 100 questions and evaluated them using three different LLMs (gpt-4o-mini, llama-3.1-8b-instruct, and gemma-2-9b-it).
-We measured inter-model agreement using Cohen’s Kappa as shown in Table 6, which showed substantial to almost perfect consistency.
-This indicates that the evaluation results are consistent across different LLMs, even when the model generating the responses is not the same as the one evaluating them (e.g., using gpt-4o-mini for generation and llama-3.1-8b-instruct for evaluation).
+To address the concern regarding potential self-enhancement bias in LLM-as-evaluator frameworks, we provide extensive validation of our evaluation approach.
+In specific, we randomly sample 100 questions and evaluated them using three different LLMs (i.e., gpt-4o-mini, llama-3.1-8b-instruct, and gemma-2-9b-it).
+We measure the inter-model agreement using Cohen’s Kappa as shown in Table 6.
+It indicates that the evaluation results are consistent across different LLMs, even when the model generating the responses is not the same as the one evaluating them (e.g., using gpt-4o-mini for generation and llama-3.1-8b-instruct for evaluation).
 These findings confirm that the evaluation is robust and independent of the specific LLM used as the evaluator.
+In addition, we also collect human evaluations for these 100 samples.
+Three expert annotators rate each anonymized response on context relevance, comprehensiveness, correctness, and empowerment using a 1–5 Likert scale.
+The average human ratings are then computed and compared with automated scores obtained via G-Eval.
+The Pearson correlation coefficient of 0.78 indicates strong alignment between human judgments and LLM-based evaluations.
+Together with the inter-model agreement reported in Table 6, these results demonstrate that our evaluation is robust, consistent, and largely independent of the specific LLM used as the evaluator.
 Metric LLM 1 & 2 LLM 1 & 3 LLM 2 & 3
@@
 LLM 1: gpt-4o-mini; LLM 2: llama-3.1-8b-instruct; LLM 3: gemma-2-9b-it)
-In addition, we also collect human evaluations for these 100 samples.
-Three experts annotators rate each anonymized response on context relevance, comprehensiveness, correctness and empowerment using a 1-5 Likert scale.
-The average human ratings are computed and compared with automated scores using G-Eval.
-The Pearson’s correlation yields a score of 0.78, indicating strong alignment between human judgment and LLM-based evaluation.
-Combined with the inter-model agreement shown in Table 6, these results confirm that our evaluation is robust, consistent, and largely independent of the specific LLM used as the evaluator.
 A.5 KG Similarity Metrics
@@
 One emerging research trend is enhancing LLMs through integrating external knowledge graphs (Agrawal et al., 2023).
-KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023; Sui et al., 2022).
+KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning (Sui, 2021; Zheng et al., 2023; Agrawal et al., 2023).
 Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance.
-This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent.
-Researchers employ diverse strategies to augment the LLMs by integrating external KGs (Sui et al., 2024; He et al., 2024b).
+This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning.
+Researchers employ diverse strategies to augment the LLMs by integrating external KGs (Sui et al., 2024; He et al., 20 2, 2025; Sui et al., 2022).
 For example, KAPING (Baek et al., 2023b) matches entities in questions to retrieve related triples from knowledge graphs for zero-shot question answering.
@@
 Following the succuess of internal reasoning-enhancement methods like Chain-of-thoughts (CoT) (Wei et al., 2022), Reflexion (Shinn et al., 2024), and Tree-of-thoughts (ToT), He et al.
-(2022) propose “rethinking with retrieval” to use decomposed reasoning steps from CoT prompting to retrieve external knowledge, leading to more accurate and faithful explanations.
+(2023a) propose “rethinking with retrieval” to use decomposed reasoning steps from CoT prompting to retrieve external knowledge, leading to more accurate and faithful explanations.
 IR-CoT (Trivedi et al., 2022b) interleaves the generation of CoT with knowledge retrieval from corresponding KGs, iteratively guiding both retrieval and reasoning for multi-step questions.
 MindMap (Wen et al., 2023) introduce a plug-and-play approach to evoke graph-of-thoughts reasoning in LLMs.
-Similarly, RoG (Luo et al., 2023b) use KGs to create faithful reasoning paths based on various relations, enabling interpretable reasoning in LLMs.
+Similarly, RoG (Luo et al., 2023b) use KGs to create faithful reasoning paths based on various relations, enabling interpretable reasoning.
 However, current benchmarks for testing the capabilities of these LLM+KG models are predominantly closed-ended, restricting responses to a limited set of entities/relations or a set of logical forms derived from specific facts of KG.